Jefferson Lab HALL A Logbook for November 08 - All EXCEPT Automatic Entries
Main INDEX
Sunday
251549 11/30/08 23:50 bradshaw Polarization - Swing Shift
251548 11/30/08 23:45 Chiranjib LHRS VDC efficiency
251547 11/30/08 23:41 xqian BigBite Wire Chamber Gas Start Bubbling
251544 11/30/08 23:04 xqian another Hitting efficiency plot
251543 11/30/08 22:39 deng For run 4075, the Happex run is 30488
251540 11/30/08 22:21 deng check list for lumis
251539 11/30/08 21:33 deng happex restart at 21:30pm
251536 11/30/08 21:09 xqian Hitting efficiency for RUN 4072
251535 11/30/08 21:06 xqian Data Check for the BigBite Wire Chamber Gas Problem
251534 11/30/08 20:59 deng check list for lumis, run= 30487
251533 11/30/08 20:43 M. Iodice Run 4073 RICH DAQ TEST hits threshold 180
251532 11/30/08 20:42 Kalyan Dead time is 12% for run 4073
251529 11/30/08 20:27 adaq Target Field and Oven Temps snapshot
251528 11/30/08 20:27 adaq Target Cell Temps snapshot
251527 11/30/08 20:26 adaq Radiation and Ozone snapshot
251526 11/30/08 20:26 adaq Hall A Tools snapshot
251525 11/30/08 20:26 adaq BigBox Magnet snapshot
251524 11/30/08 20:26 adaq Beam Position and Current snapshot
251523 11/30/08 20:26 adaq Beam Energy in A and C snapshot
251522 11/30/08 20:05 adaq Scaler rates (BigBite) snapshot
251521 11/30/08 19:52 Kalyan Replaced argon/ethane gas bottles for BB wirechambers
251518 11/30/08 18:53 adaq Hall A general tools screen shots
251517 11/30/08 18:43 H. Gao Spot++ run 4072
251514 11/30/08 18:04 bradshaw Absorption Spectrum
251513 11/30/08 18:02 deng check list for Lumis
251512 11/30/08 17:17 deng Happex restart at 17:20 pm
251509 11/30/08 16:40 adev BigBite WireChamber Chamber Turn on/off!
251508 11/30/08 16:30 adev BigBite WireChamber Chamber Turn on/off!
251507 11/30/08 16:18 M. Iodice DAQ rich num hits threshold scan
251506 11/30/08 16:16 Kalyan BB Shower HV tripped during run 4069
251503 11/30/08 16:05 gaskell Day Shift Summary
251502 11/30/08 16:05 adaq Scaler rates (BigBite) snapshot
251496 11/30/08 15:44 M. Iodice RICH HV 2150 for runs 4064 and 4065
251491 11/30/08 14:57 xqian BigBite Wire Chamber hitting efficiency
251490 11/30/08 14:49 gaskell spot++
251489 11/30/08 14:33 gaskell restored links to adaql10
251488 11/30/08 14:24 xqian BigBite Wire Chamber Standard Check
251485 11/30/08 13:59 Ole Hansen Replaced adaql10 CPU fan
251484 11/30/08 13:20 M. Iodice Run 4064 - Happex is now back in DAQ
251483 11/30/08 13:19 yawei ROC 23 works, happex run#30486 starts
251480 11/30/08 13:07 gaskell rebooted ROC 23 (parity DAQ)
251479 11/30/08 12:57 M. Iodice Rich HV change from run 4063
251478 11/30/08 12:55 M. Iodice Happex DAQ hanging
251477 11/30/08 12:53 gaskell re-directed online analysis to adaqfs/scratch
251476 11/30/08 12:41 gaskell ROC 23 down - no HAPPEX run
251475 11/30/08 12:26 adaq Target Field and Oven Temps snapshot
251474 11/30/08 12:26 adaq Target Cell Temps snapshot
251473 11/30/08 12:26 adaq Radiation and Ozone snapshot
251472 11/30/08 12:26 adaq Hall A Tools snapshot
251471 11/30/08 12:26 adaq BigBox Magnet snapshot
251470 11/30/08 12:26 adaq Beam Position and Current snapshot
251469 11/30/08 12:26 adaq Beam Energy in A and C snapshot
251467 11/30/08 12:05 adaq Scaler rates (BigBite) snapshot
251465 11/30/08 11:58 gaskell disk space
251463 11/30/08 11:55 Ole Hansen adaql10 - help is on the way
251461 11/30/08 11:43 gaskell adaql10 stopped again
251460 11/30/08 11:04 M. Iodice new happex run for runs 4059 and 4060
251456 11/30/08 10:16 Yi Qiang Updated Summary for Nov 29, Owl - Swing
251455 11/30/08 10:06 yawei left arm VDC efficiency, run#4059
251454 11/30/08 09:56 yawei preshower and shower energy check, run#4058
251451 11/30/08 09:33 Yi Qiang Experiment Summary for Nov 29, Owl - Swing
251450 11/30/08 09:21 yawei Laser Absorption Spectrum
251449 11/30/08 09:04 yawei spot++ for run #4059
251446 11/30/08 08:04 adaq Scaler rates (BigBite) snapshot
251444 11/30/08 07:58 R. Michaels owl summary
251441 11/30/08 07:07 R. Michaels tools
251440 11/30/08 06:48 Yi Zhang LHRS Chamber Efficiency
251437 11/30/08 06:38 R. Michaels auto BANU problem/repair
251436 11/30/08 05:47 A. Camsonne Threshold change on number of hits for RICH
251433 11/30/08 05:11 R. Michaels/ Absorption Spectrum
251432 11/30/08 04:49 R. Michaels new HAPPEX run 30484
251429 11/30/08 04:41 R. Michaels DT problem back to 8% (mystery solved !!)
251428 11/30/08 04:30 adev BigBite WireChamber Measured Voltage change!
251427 11/30/08 04:26 adaq Target Field and Oven Temps snapshot
251426 11/30/08 04:26 adaq Target Cell Temps snapshot
251425 11/30/08 04:26 adaq Radiation and Ozone snapshot
251424 11/30/08 04:26 adaq Hall A Tools snapshot
251423 11/30/08 04:25 adaq BigBox Magnet snapshot
251422 11/30/08 04:25 adaq Beam Position and Current snapshot
251421 11/30/08 04:25 adaq Beam Energy in A and C snapshot
251420 11/30/08 04:25 R. Michaels restarted hvgui on adaql3
251419 11/30/08 04:20 adev BigBite WireChamber Measured Voltage change!
251418 11/30/08 04:11 Annand 3He Target
251417 11/30/08 04:09 adev BigBite WireChamber Measured Voltage change!
251416 11/30/08 04:06 adaq target: Spin Flip Report: 3He Target status
251415 11/30/08 04:04 adaq Scaler rates (BigBite) snapshot
251414 11/30/08 04:01 R. Michaels fix start/end GUI pop-up
251412 11/30/08 03:56 R. Michaels DT = 25%
251409 11/30/08 03:50 R. Michaels spot++
251407 11/30/08 03:43 R. Michaels junk runs, power-cycled TS11
251403 11/30/08 03:24 adev SCALER = -1
251402 11/30/08 03:18 R. Michaels high deadtime, scaler=-1
251401 11/30/08 03:17 adev SCALER = -1
251397 11/30/08 03:10 adev SCALER = -1
251396 11/30/08 03:03 adev SCALER = -1
251391 11/30/08 02:44 R. Michaels BB HV trip
251387 11/30/08 02:25 R. Michaels adaql10 died again. move to another (new) UPS.
251384 11/30/08 01:26 ycwang L-arm VDC efficiency at Run #4042
251383 11/30/08 01:23 ycwang RICH Detector number of photons
251382 11/30/08 01:00 Bob/Youcai hit rejection on RICH
251379 11/30/08 00:35 R. Michaels adaql10 shut itself off at 23:30
251378 11/30/08 00:04 adaq Scaler rates (BigBite) snapshot
251377 11/30/08 00:03 Gao Swing shift summary!
Saturday
251371 11/29/08 21:52 ycwang L-arm VDC efficiency
251368 11/29/08 20:49 Jin Golden RootFile -> 4029; Reduce replay number
251365 11/29/08 20:25 adaq Target Field and Oven Temps snapshot
251364 11/29/08 20:25 adaq Target Cell Temps snapshot
251363 11/29/08 20:25 adaq Radiation and Ozone snapshot
251362 11/29/08 20:25 adaq Hall A Tools snapshot
251361 11/29/08 20:25 adaq BigBox Magnet snapshot
251360 11/29/08 20:25 adaq Beam Position and Current snapshot
251359 11/29/08 20:25 adaq Beam Energy in A and C snapshot
251358 11/29/08 20:04 adaq Scaler rates (BigBite) snapshot
251355 11/29/08 19:34 Jin Web Page for Target & Beam Real Time Status
251354 11/29/08 19:13 M.Shabestari Target ---> TO report: Swing Shift
251353 11/29/08 19:08 adaq target: Spin Flip Report: Target ---> Spin Flip Report: TO report - Swing shift
251352 11/29/08 19:01 M.Shabestari Laser absorption spectrum
251351 11/29/08 19:00 H. Gao Hall Tool screen
251348 11/29/08 18:47 H. Gao Spot++ on Run 4036
251347 11/29/08 17:51 M.Shabestari Target Stripcharts
251344 11/29/08 17:39 adaq BigBite Wire Chamber Hitting efficiency
251343 11/29/08 17:33 Jin VDC Efficiency, run # 4033
251342 11/29/08 17:24 S. Sirca Replay crashes with X11 connection error
251340 11/29/08 16:48 M.Shabestari BigBite HV tripped
251338 11/29/08 16:38 A. Camsonne Start and End of run windows
251336 11/29/08 16:08 Richard Lind Day Shift Summary
251335 11/29/08 16:04 adaq Scaler rates (BigBite) snapshot
251334 11/29/08 15:52 M. Iodice coda start and end run pop up window
251332 11/29/08 15:47 xqian BigBite Wire Chamber Standard Check
251330 11/29/08 15:41 yawei L-arm VDC efficiency, run #4029
251329 11/29/08 15:37 M. Iodice runs 4030 and 4031 junk - coda problems
251324 11/29/08 14:02 M. Iodice Hall A General Tool Screen
251323 11/29/08 13:57 M. Iodice spot++ on run 4029
251322 11/29/08 13:46 Lindgren Beam unstable
251321 11/29/08 13:43 Katich Wiki - Instructions for Lumi HV Change (for experts)
251318 11/29/08 12:39 Moffit Re: HAPPEX DAQ bcm4... missing asymmetry plots.
251317 11/29/08 12:34 Katich Lumi Shift worker checklist updated
251316 11/29/08 12:25 adaq Target Field and Oven Temps snapshot
251315 11/29/08 12:25 adaq Target Cell Temps snapshot
251314 11/29/08 12:25 adaq Radiation and Ozone snapshot
251313 11/29/08 12:25 adaq Hall A Tools snapshot
251312 11/29/08 12:25 adaq BigBox Magnet snapshot
251311 11/29/08 12:25 adaq Beam Position and Current snapshot
251310 11/29/08 12:24 adaq Beam Energy in A and C snapshot
251307 11/29/08 12:05 Moffit HAPPEX DAQ: Feedback to use "bcm4" x10
251306 11/29/08 12:04 adaq Scaler rates (BigBite) snapshot
251305 11/29/08 11:21 Yi Qiang The ABU and BANU are not correct
251302 11/29/08 11:18 Yi Qiang Experiment Summary for Nov 28, Owl - Swing
251301 11/29/08 10:51 Lindgren Beam position monitor
251300 11/29/08 10:48 Lindgren Raster size
251299 11/29/08 10:45 Lindgren Hall A Tools and Beam current monitor
251296 11/29/08 09:30 adaq VDC efficiency screenshot
251293 11/29/08 08:31 yawei preshower and shower energy, run#4022
251292 11/29/08 08:25 Lindgren Day shift beginning
251289 11/29/08 08:06 adaq target: Spin Flip Report: spin flip system view
251287 11/29/08 08:04 adaq Scaler rates (BigBite) snapshot
251286 11/29/08 08:03 Yi Zhang Owl Shift Summary
251285 11/29/08 08:03 yawei laser absorption spectrum
251284 11/29/08 07:59 Yi Zhang HallA main tools
251283 11/29/08 07:57 adaq LHRS VDC efficiency
251282 11/29/08 07:41 Yi Zhang spot++ @ run4023
251275 11/29/08 04:24 adaq Target Field and Oven Temps snapshot
251274 11/29/08 04:24 adaq Target Cell Temps snapshot
251273 11/29/08 04:24 adaq Radiation and Ozone snapshot
251272 11/29/08 04:24 adaq Hall A Tools snapshot
251271 11/29/08 04:24 adaq BigBox Magnet snapshot
251270 11/29/08 04:24 adaq Beam Position and Current snapshot
251269 11/29/08 04:24 adaq Beam Energy in A and C snapshot
251266 11/29/08 04:04 adaq Scaler rates (BigBite) snapshot
251261 11/29/08 01:21 Katich Lumi HVs set during run 4016
251258 11/29/08 00:43 Katich Lumis 1,2,5 saturated
251255 11/29/08 00:06 Chiranjib Paged Joe
251254 11/29/08 00:04 adaq Scaler rates (BigBite) snapshot
251252 11/29/08 00:00 Gao swing shift summary
Friday
251251 11/28/08 23:58 Jin Lumi 1 & 2 abnormal
251250 11/28/08 23:45 M.Shabestari Laser absorption spectrum
251249 11/28/08 23:43 M.Shabestari Target stripcharts
251248 11/28/08 23:27 M.Shabestari Target ---> 3He
251247 11/28/08 23:18 Jin Spin flip is started for production
251245 11/28/08 23:12 adaq target: Spin Flip Report: Spin Flip for Production Started. Pol = ~66%
251244 11/28/08 22:59 Jin Target -> He3
251242 11/28/08 22:11 Chiranjib adaql10 rebooted
251241 11/28/08 22:06 M.Shabestari LASER aborption spectrum
251240 11/28/08 22:05 Jin adaql10 is down
251239 11/28/08 22:02 Chiranjib DISK problem /broken links
251234 11/28/08 20:24 adaq Target Field and Oven Temps snapshot
251233 11/28/08 20:24 adaq Target Cell Temps snapshot
251232 11/28/08 20:24 adaq Radiation and Ozone snapshot
251231 11/28/08 20:24 adaq Hall A Tools snapshot
251230 11/28/08 20:24 adaq BigBox Magnet snapshot
251229 11/28/08 20:24 adaq Beam Position and Current snapshot
251228 11/28/08 20:24 adaq Beam Energy in A and C snapshot
251227 11/28/08 20:12 M.Shabestari Target Stripcharts
251226 11/28/08 20:04 adaq Scaler rates (BigBite) snapshot
251222 11/28/08 18:41 M.Shabestari Target ---> H2
251218 11/28/08 18:00 xqian Hitting efficiency for BigBite Wire Chamber
251217 11/28/08 17:29 S. Sirca Spot++ check
251216 11/28/08 17:28 S. Sirca General Tools screenshot
251213 11/28/08 17:05 xqian BigBite Wire Chamber Standard Check
251212 11/28/08 17:05 M.Shabestari Target: N2 reference
251208 11/28/08 16:53 M. Shabestar Target: Empty reference
251204 11/28/08 16:32 xqian BiGBite Wire Chamber Hitting efficiency
251203 11/28/08 16:22 Lindgren spot++
251201 11/28/08 16:06 Lindgren day shift summary:
251200 11/28/08 16:04 adaq Scaler rates (BigBite) snapshot
251199 11/28/08 16:03 Kalyan Summary of reference cell runs with 5 pass upto run #3949
251198 11/28/08 15:50 Lindgren Beam polarization
251197 11/28/08 15:48 Lindgren Production running on reference cell
251195 11/28/08 15:47 Lindgren Rate as Raster varies
251194 11/28/08 15:46 Yi Qiang EPR Calibration for Transverse Pumping
251193 11/28/08 15:43 E.Chudakov Moller results
251192 11/28/08 15:40 Yi Qiang Fake spin signal enabled for reference cell runs
251191 11/28/08 15:36 bradshaw Move to Reference Cell
251190 11/28/08 15:29 bradshaw Moving to BeO
251189 11/28/08 15:02 Lindgren Polarization measurement
251188 11/28/08 14:57 bradshaw Moving Target to He3 Position
251187 11/28/08 14:45 bradshaw Move to BeO
251186 11/28/08 14:40 Lindgren Moller complete
251185 11/28/08 13:12 E.Chudakov Moller tune
251184 11/28/08 12:59 Moffit HAPPEX DAQ: investigation of dx3 signal to happex7 scaler.
251183 11/28/08 12:25 Lindgren Hall A Tools
251182 11/28/08 12:23 adaq Target Field and Oven Temps snapshot
251181 11/28/08 12:23 adaq Target Cell Temps snapshot
251180 11/28/08 12:23 adaq Radiation and Ozone snapshot
251179 11/28/08 12:23 adaq Hall A Tools snapshot
251178 11/28/08 12:23 adaq BigBox Magnet snapshot
251177 11/28/08 12:23 adaq Beam Position and Current snapshot
251176 11/28/08 12:23 adaq Beam Energy in A and C snapshot
251175 11/28/08 12:10 adaq target: Spin Flip Report: Auto spin flip are ready for Transverse Pumping
251174 11/28/08 12:04 adaq Scaler rates (BigBite) snapshot
251173 11/28/08 12:00 Chiranjib TRANSVERSE PUMPING STARTED
251172 11/28/08 11:18 A. Camsonne Compton rates stable
251171 11/28/08 11:12 Kalyan BB wirechamber HV turned off for Moller measurement
251170 11/28/08 11:09 Lindgren Moller Measurement
251169 11/28/08 11:06 adaq target: Spin Flip Report: Report for the End of the Vertical Pumping
251168 11/28/08 11:01 Yi/Chiranjib NMR Cross Calibration
251167 11/28/08 10:59 E.Chudakov Start Moller tune
251166 11/28/08 10:47 Lindgren Target spin change
251165 11/28/08 10:45 E.Chudakov Energy lock for low currents
251164 11/28/08 10:45 M. Iodice Rich Performances
251163 11/28/08 10:34 yawei L-arm VDC efficiency, run #4003
251162 11/28/08 10:31 Chiranjib/Yi EPR calibration
251161 11/28/08 10:31 yawei preshower and shower energy, run #4003
251160 11/28/08 10:30 Yi Qiang Target polarization change starts
251159 11/28/08 10:30 yawei preshower and shower energy, run #4003
251158 11/28/08 10:21 Yi Zhang TDC & ADC of BB Cherenkov after lowered the threshold.
251157 11/28/08 09:47 bradshaw Moved Target for Moeller Test
251153 11/28/08 08:54 E.Chudakov N2 cylinder changed
251152 11/28/08 08:30 Lindgen Day shift beginning
251146 11/28/08 08:04 Yi Zhang Owl Shift
251145 11/28/08 08:04 adaq Scaler rates (BigBite) snapshot
251144 11/28/08 07:53 Yi Zhang spot++ @ run#4001
251141 11/28/08 07:35 M. Iodice High Dead Time during run 4000
251140 11/28/08 07:35 Yi Zhang Hall A main tools
251139 11/28/08 06:06 M. Iodice Rich problem during HV scan
251134 11/28/08 04:29 Nelyubin Absorption Spectrum 11/28/08 4:25a.m.
251133 11/28/08 04:24 adaq target: Spin Flip Report: SpinFlipSystemStatus
251132 11/28/08 04:23 adaq Target Field and Oven Temps snapshot
251131 11/28/08 04:23 adaq Target Cell Temps snapshot
251130 11/28/08 04:23 adaq Radiation and Ozone snapshot
251129 11/28/08 04:23 adaq Hall A Tools snapshot
251128 11/28/08 04:23 adaq BigBox Magnet snapshot
251127 11/28/08 04:23 adaq Beam Position and Current snapshot
251126 11/28/08 04:23 adaq Beam Energy in A and C snapshot
251124 11/28/08 04:04 adaq Scaler rates (BigBite) snapshot
251114 11/28/08 02:10 M.Iodice Rich T/H delay scan completed
251113 11/28/08 01:53 ycwang Run # 3989 for RICH pedestal run with HV=2250V, Run 3990-3992 are junk.
251102 11/28/08 00:23 E.Chudakov Experiment summary for Nov 27(owl) - 27(swing)
251101 11/28/08 00:04 adaq Scaler rates (BigBite) snapshot
251100 11/28/08 00:01 S. Sirca Swing shift SUMMARY
Thursday
251096 11/27/08 23:07 adaq Target (3He) states
251093 11/27/08 22:26 ycwang L-arm VDC efficiency
251090 11/27/08 21:24 S. Sirca General Tools screen
251089 11/27/08 21:23 S. Sirca Spot++ check
251088 11/27/08 21:21 ycwang new pedestal for RICH after time delay to -152 ns
251087 11/27/08 21:09 ycwang pedestal run # 3983 for RICH
251082 11/27/08 20:33 adaq Target (3He) states
251081 11/27/08 20:23 adaq Target Field and Oven Temps snapshot
251080 11/27/08 20:23 adaq Target Cell Temps snapshot
251079 11/27/08 20:22 adaq Radiation and Ozone snapshot
251078 11/27/08 20:22 adaq Hall A Tools snapshot
251077 11/27/08 20:22 adaq BigBox Magnet snapshot
251076 11/27/08 20:22 adaq Beam Position and Current snapshot
251075 11/27/08 20:22 adaq Beam Energy in A and C snapshot
251074 11/27/08 20:03 adaq Scaler rates (BigBite) snapshot
251069 11/27/08 17:48 yawei deadtime turn to mormal now
251068 11/27/08 17:42 xqian Checked the BigBite drift time for each card
251066 11/27/08 17:41 brads BB Cerenkov beam-right PMTs turned off after run 3979
251065 11/27/08 17:39 xqian Wire Map looks good
251063 11/27/08 17:32 Kalyan Summary of reference cell runs taken after installing the new target cell
251062 11/27/08 17:14 brads deadtime calculation may be thrown by the non-zero T2, T7 rates
251061 11/27/08 16:52 E.Chudakov N2 cylinder replaced
251060 11/27/08 16:50 yawei high deadtime
251057 11/27/08 16:29 brads BB Cerenkov witness PMT rates with borated plastic shielding
251056 11/27/08 16:13 brads BB Cerenkov beam-right PMTs turned on during run 3978
251055 11/27/08 16:03 adaq Scaler rates (BigBite) snapshot
251054 11/27/08 16:01 Jin Huang Shift Summary, Nov 27 Day
251052 11/27/08 15:48 adaq target: Spin Flip Report: TO Report - Day Shift
251049 11/27/08 15:03 Kalyan VDC efficiency
251048 11/27/08 14:56 Jin Hall A General Tools Screen
251047 11/27/08 14:55 Jin Spot++ for Nov 27 Day shift
251044 11/27/08 13:20 Jin BB HV Error & Cleared
251041 11/27/08 12:22 adaq Target Field and Oven Temps snapshot
251040 11/27/08 12:22 adaq Target Cell Temps snapshot
251039 11/27/08 12:22 adaq Radiation and Ozone snapshot
251038 11/27/08 12:22 adaq Hall A Tools snapshot
251037 11/27/08 12:22 adaq BigBox Magnet snapshot
251036 11/27/08 12:22 adaq Beam Position and Current snapshot
251035 11/27/08 12:22 adaq Beam Energy in A and C snapshot
251032 11/27/08 12:03 adaq Scaler rates (BigBite) snapshot
251029 11/27/08 10:48 yawei preshower and shower energy check, run #3972
251028 11/27/08 10:24 Jin Golden Root File -> Run # 3969
251027 11/27/08 10:03 A. Camsonne Track and Hold delayed by 48 ns
251020 11/27/08 08:03 adaq Scaler rates (BigBite) snapshot
251018 11/27/08 07:56 Chiranjib OWL SHIFT SUMMARY
251015 11/27/08 06:49 M. Iodice Rich T/H delay scan
251014 11/27/08 06:12 A. Camsonne T/H current setting
251009 11/27/08 04:33 Nelyubin Absorption Spectrum
251008 11/27/08 04:28 adaq target: Spin Flip Report: SpinFlipSystemControl_112708_425am
251007 11/27/08 04:22 adaq Target Field and Oven Temps snapshot
251006 11/27/08 04:22 adaq Target Cell Temps snapshot
251005 11/27/08 04:22 adaq Radiation and Ozone snapshot
251004 11/27/08 04:22 adaq Hall A Tools snapshot
251003 11/27/08 04:22 adaq BigBox Magnet snapshot
251002 11/27/08 04:21 adaq Beam Position and Current snapshot
251001 11/27/08 04:21 adaq Beam Energy in A and C snapshot
251000 11/27/08 04:03 adaq Scaler rates (BigBite) snapshot
250999 11/27/08 03:59 M. Iodice Hall A General Tools Screen
250998 11/27/08 03:53 adaq empty channel on BB preshower
250993 11/27/08 02:22 Chiranjib spot++
250992 11/27/08 02:00 E.Chudakov Moller on Nov 17 - possible beam energy change
250989 11/27/08 01:47 M.Iodice Rich optimization T/H delay
250988 11/27/08 01:13 xqian Hitting efficiency is high
250987 11/27/08 01:10 M.Iodice End of Run 3963
250984 11/27/08 00:31 E.Chudakov Experiment summary for Nov 26(owl) - 26(swing)
250983 11/27/08 00:03 adaq Scaler rates (BigBite) snapshot
Wednesday
250981 11/26/08 23:54 J,C, Peng Swing shift summary
250978 11/26/08 23:06 Nelyubin Target System
250977 11/26/08 23:04 adaq target: Spin Flip Report: SpinFlipSystemControl
250976 11/26/08 22:48 adaq target: Spin Flip Report: Spin Flip System Status Report
250975 11/26/08 22:43 adaq Hall A tool screen
250972 11/26/08 21:48 Peng Raster spot
250971 11/26/08 21:39 Chiranjib DAQ reboot.
250969 11/26/08 20:40 L. Guo coda
250967 11/26/08 20:32 Chiranjib END of RUN 3958 problem
250965 11/26/08 20:21 adaq Target Field and Oven Temps snapshot
250964 11/26/08 20:21 adaq Target Cell Temps snapshot
250963 11/26/08 20:21 adaq Radiation and Ozone snapshot
250962 11/26/08 20:21 adaq Hall A Tools snapshot
250961 11/26/08 20:21 adaq BigBox Magnet snapshot
250960 11/26/08 20:21 adaq Beam Position and Current snapshot
250959 11/26/08 20:21 adaq Beam Energy in A and C snapshot
250956 11/26/08 20:10 ycwang RUN # 3958 already started at 19:16
250955 11/26/08 20:03 adaq Scaler rates (BigBite) snapshot
250953 11/26/08 18:39 Kalyan db_DL.dat database updated with BCM and trigger information for every event
250952 11/26/08 18:39 Kalyan db_DL.dat database updated with BCM and trigger information for every event
250951 11/26/08 18:38 Kalyan db_DL.dat database updated with BCM and trigger information for every event
250950 11/26/08 18:18 Kalyan Target polarization chart
250947 11/26/08 17:25 Moffit Re: dx3 in HAPPEX DAQ
250946 11/26/08 17:23 ycwang RUN # 3951 3952 3953 is not good data for RICH. From Run #3955 the data for RICH is good.
250945 11/26/08 17:10 Kalyan BCM d3 is back again in HAPPEX DAQ
250942 11/26/08 16:51 Jin Happex D3 Counting Dropping
250941 11/26/08 16:27 Peng Raster profile
250938 11/26/08 16:03 adaq Scaler rates (BigBite) snapshot
250933 11/26/08 15:54 gaskell Day shift summary
250932 11/26/08 15:52 adaq target: Spin Flip Report: TO - Day Shift
250929 11/26/08 15:06 gaskell Spot++ (checklist)
250927 11/26/08 14:35 Jin Spot++ for 1sr Production run
250926 11/26/08 14:32 Jin Spot++ for 2uA; raster 5x6.6
250924 11/26/08 14:31 gaskell Compton rates vs. slit position
250921 11/26/08 14:26 Jin Trigger Screen Shot; 12uA; 3x4 raster
250920 11/26/08 14:26 adaq target: Spin Flip Report: Real Spin flip started @ Pol.=~63%
250919 11/26/08 14:22 gaskell raster size scan
250917 11/26/08 14:20 Jin Trigger Screenshot; 5x6.6 raster
250915 11/26/08 14:16 Jin Trigger Screen follow up
250914 11/26/08 14:15 Jin Trigger Screenshot, 4x5 raster
250913 11/26/08 14:14 Jin Trigger, 3x4
250912 11/26/08 14:11 Jin Trigger 2uA, 2x2
250911 11/26/08 14:07 gaskell Finished with reference cell - moving to polarized 3He
250910 11/26/08 14:05 Kalyan RICH histograms doesn't look good
250908 11/26/08 14:02 S. Sirca Change in RICH spectra
250907 11/26/08 13:25 E.Chudakov Meeting today at 16:00. Current status.
250905 11/26/08 12:54 bradshaw Move to Reference Cell
250904 11/26/08 12:50 gaskell Compton ion chamber calibration
250903 11/26/08 12:21 adaq Target Field and Oven Temps snapshot
250902 11/26/08 12:21 adaq Target Cell Temps snapshot
250901 11/26/08 12:21 adaq Radiation and Ozone snapshot
250900 11/26/08 12:21 adaq Hall A Tools snapshot
250899 11/26/08 12:21 adaq BigBox Magnet snapshot
250898 11/26/08 12:21 adaq Beam Position and Current snapshot
250897 11/26/08 12:21 adaq Beam Energy in A and C snapshot
250896 11/26/08 12:03 adaq Scaler rates (BigBite) snapshot
250895 11/26/08 12:03 Jin HRS angle in database corrected
250894 11/26/08 11:58 bradshaw Move to BeO
250893 11/26/08 11:50 gaskell HRS ready - ask for beam
250891 11/26/08 11:36 gaskell Back to beam permit
250890 11/26/08 11:35 Kalyan Added T1, T3, T5 and T6 scaler counts in every event
250889 11/26/08 11:33 Jin Fake spin signal back ON
250888 11/26/08 11:28 Yi Qiang Heater is back
250887 11/26/08 11:21 ewing/fansle cryo trip
250886 11/26/08 11:05 Kalyan duplicate copy of T3 in L-HRS scaler
250885 11/26/08 11:05 Jin He3 Heater is Down
250884 11/26/08 10:49 Yi Qiang AFP Loss Test for Vertical Pumping
250883 11/26/08 10:46 Yi Qiang EPR Calibration for Maureen of Vertical Pumping
250882 11/26/08 10:43 E.Chudakov Faulty modules in the beam HV LeCroy
250881 11/26/08 10:40 Yi Qiang New Frequency Sweep Settings for Maureen Vertical pumping
250880 11/26/08 09:54 V. Sulkosky EPICS for Energy and BPMs
250879 11/26/08 09:44 A. Camsonne more RICH T/H
250878 11/26/08 09:41 Yi Qiang Fake spin signal stopped for calibration
250877 11/26/08 09:39 A. Camsonne Track and hold delay
250876 11/26/08 09:35 brads Energy and radiation EPICS data missing too
250872 11/26/08 09:28 brads BPM readbacks missing since 11pm last night?
250871 11/26/08 09:26 gaskell ESR down again - going to COntrolled Access
250870 11/26/08 09:00 gaskell IOC vs ESR takes down HRS
250868 11/26/08 08:22 A. Camsonne Owl shift summary
250866 11/26/08 08:05 A. Camsonne Beamline HV tripped
250865 11/26/08 08:03 adaq Scaler rates (BigBite) snapshot
250864 11/26/08 07:58 A. Camsonne Spot ++
250863 11/26/08 07:55 M.Iodice Rates with 12 uA on 110 psig He3
250862 11/26/08 07:51 A. Camsonne HV fine
250861 11/26/08 07:45 M.Iodice Rich Track/Hold scan results
250858 11/26/08 07:39 A. Camsonne HV alarm
250857 11/26/08 06:50 A. Camsonne End of run failed , restarted CODA and HAPPEX DAQ
250854 11/26/08 06:34 yawei Adjust BB chamber in 12uA when refrence cell filled into He3 to 110psig
250853 11/26/08 06:32 yawei fill He3 into reference cell to 110psig
250849 11/26/08 05:49 A. Camsonne RICH T/H to nominal value
250847 11/26/08 05:23 Chiranjib Laser absorption spectrum
250846 11/26/08 04:42 A. Camsonne RICH T/H to -64 ns from nominal value
250843 11/26/08 04:21 adaq Target Field and Oven Temps snapshot
250842 11/26/08 04:20 adaq Target Cell Temps snapshot
250841 11/26/08 04:20 adaq Radiation and Ozone snapshot
250840 11/26/08 04:20 adaq Hall A Tools snapshot
250839 11/26/08 04:20 adaq BigBox Magnet snapshot
250838 11/26/08 04:20 adaq Beam Position and Current snapshot
250837 11/26/08 04:20 adaq Beam Energy in A and C snapshot
250836 11/26/08 04:19 A. Camsonne Beam had trouble but now fine
250835 11/26/08 04:03 adaq Scaler rates (BigBite) snapshot
250834 11/26/08 03:43 A. Camsonne Compton Edet run
250833 11/26/08 03:42 Chiranjib rates with 12 uA on 110 psig H2
250832 11/26/08 03:39 yawei Adjust BB chamber in 12uA when reference cell filled into H2 to 110psig
250830 11/26/08 03:34 yawei fill H2 into refrence cell to 110psig
250829 11/26/08 03:30 A. Camsonne RICH T/H moved to +64 ns from nominal
250822 11/26/08 02:57 A. Camsonne ROC9 rebooted by itself during run 3935
250821 11/26/08 02:55 A. Camsonne RICH T/H moved +120 ns
250818 11/26/08 02:53 Yi Zhang RE: Set angle of HRS in Hall A tools
250817 11/26/08 02:41 A. Camsonne Set angle of HRS in Hall A tools
250816 11/26/08 02:32 A. Camsonne RICH T/H removed 120 ns
250815 11/26/08 02:31 A. Camsonne Run 3933 ROC9 crashed
250811 11/26/08 02:00 A. Camsonne Final beam position
250809 11/26/08 01:45 ywang Adjust BB chamber HV in 12 uA when reference cell filled into N2 to 26psig
250806 11/26/08 01:38 yawei fill N2 to 26 psig into reference cell
250803 11/26/08 01:33 A. Camsonne Empty reference cell 5 uA rates
250801 11/26/08 01:30 A. Camsonne Spot current setting run 3929
250798 11/26/08 01:17 ycwang RICH HV turned on
250796 11/26/08 01:12 A. Camsonne Spot Run 3927
250795 11/26/08 01:09 E.Chudakov iocha22 restored
250793 11/26/08 01:08 A. Camsonne Rates with Raster 5x6.5 MCC units 2 uA
250790 11/26/08 01:05 A. Camsonne Spot Run 3926
250789 11/26/08 01:01 A. Camsonne Rates with raster 4x5 mm ( MCC units ) 2uA
250787 11/26/08 01:00 E.Chudakov Experiment summary for Nov 24(swing) - 25(swing)
250785 11/26/08 00:56 A. Camsonne New beam position
250784 11/26/08 00:56 A. Camsonne Spot ++ Run 3925
250783 11/26/08 00:53 A. Camsonne Rates new position 2 uA raster 3x4 MCC units
250781 11/26/08 00:48 A. Camsonne Moved to Empty reference cell
250779 11/26/08 00:45 A. Camsonne Spot run 3924
250777 11/26/08 00:28 Jin Target -> BeO
250776 11/26/08 00:13 S. Sirca IOC problems
250775 11/26/08 00:11 Jin Snap shot of BeO with 1uA beam
250773 11/26/08 00:06 S. Sirca Swing shift SUMMARY
250772 11/26/08 00:03 adaq Scaler rates (BigBite) snapshot
Tuesday
250771 11/25/08 23:39 Jin Target -> Ref Cell
250770 11/25/08 23:15 Peng Move to BeO target
250768 11/25/08 22:58 Re: Issues remains & Full Scalar consistency check added to online plots
250764 11/25/08 22:50 Peng Empty reference target run
250761 11/25/08 22:45 S. Sirca Spot++ for raster 5x6.6
250760 11/25/08 22:42 Srish Compton Spectrum
250759 11/25/08 22:42 S. Sirca Rates study with increasing raster
250754 11/25/08 22:33 brads Firefox profiles regenerated again, also How to Fix a Broken Firefox
250750 11/25/08 22:22 adaq Move target positions
250746 11/25/08 22:11 S. Sirca Beam position on BeO
250745 11/25/08 22:09 S. Sirca Harp scans after Compton tune
250744 11/25/08 22:06 Sirish Compton Studies
250741 11/25/08 21:52 S. Sirca MCC sends CW
250740 11/25/08 21:43 adaq Target Field and Oven Temps snapshot
250739 11/25/08 21:43 adaq Target Cell Temps snapshot
250738 11/25/08 21:43 adaq Radiation and Ozone snapshot
250737 11/25/08 21:43 adaq Hall A Tools snapshot
250736 11/25/08 21:43 adaq BigBox Magnet snapshot
250735 11/25/08 21:43 adaq Beam Position and Current snapshot
250734 11/25/08 21:42 adaq Beam Energy in A and C snapshot
250730 11/25/08 20:03 adaq Scaler rates (BigBite) snapshot
250729 11/25/08 19:33 ycwang RICH HV turned off before Beam tuning up
250728 11/25/08 19:32 ycwang new rich pedestal run #3903
250726 11/25/08 19:30 Jin Fake spin signal is ON; Target is under pumping
250724 11/25/08 19:21 A. Camsonne Compton Electron detector current
250723 11/25/08 19:13 Sirish Ref Chicane Tune
250720 11/25/08 18:37 jpchen Target position survey
250719 11/25/08 18:35 jpchen Reference cell
250716 11/25/08 16:48 A. Camsonne Borrowed logic module from cabinet
250715 11/25/08 16:07 brads Added shielding to BB Cerenkov
250714 11/25/08 16:03 adaq Scaler rates (BigBite) snapshot
250712 11/25/08 16:01 Lingyan Zhu day shift Summary
250711 11/25/08 15:25 Lingyan Zhu controlled access
250700 11/25/08 12:33 Kalyan Re: Issues remains & Full Scalar consistency check added to online plots
250694 11/25/08 12:21 V. Sulkosky Run 3901 Pedestal Run
250687 11/25/08 12:03 adaq Scaler rates (BigBite) snapshot
250686 11/25/08 11:42 brads BB Cerenkov PMTs signal 15 and 16 swapped at ADC (fixed)
250685 11/25/08 11:36 brads BB Cerenkov TDC discrim. set to -100mV
250674 11/25/08 10:35 R. Michaels adaqfs filling at alarming rate
250670 11/25/08 10:12 Chiranjib EPR RF amplifier
250669 11/25/08 10:09 Kalyan Replaced 1877 TDCs on ROC10
250660 11/25/08 09:38 Yi Qiang Lasers are Locked for cell installation
250659 11/25/08 09:35 Yi Qiang NMR Cross Calibration
250656 11/25/08 09:02 xqian Hitting Efficiency
250655 11/25/08 09:01 Chiranjib EPR Calibration
250654 11/25/08 08:56 Lingyan Zhu restricted access
250653 11/25/08 08:40 adaq target: Spin Flip Report: Target report after beam stops and AFP loss test. Byebye, Astral ~~~~ you did a very nice job!!
250651 11/25/08 08:37 Yi Qiang AFP Loss Test for Vertical Pumping
250649 11/25/08 08:06 Chiranjib OWL SHIFT SUMMARY
250638 11/25/08 05:25 bradshaw Laser Absorption Spect- Parallel B and S
250637 11/25/08 05:16 bradshaw 400 Spin Flips
250636 11/25/08 05:16 adaq target: Spin Flip Report: TO - 400 Spin Flip Report
250632 11/25/08 04:02 adaq Scaler rates (BigBite) snapshot
250628 11/25/08 02:34 Chiranjib SPOT ++
250625 11/25/08 01:14 Chiranjib COMET LASER EPICS
250622 11/25/08 00:16 Chiranjib Hall A tools / Beam current.
250621 11/25/08 00:11 Chiranjib HAPPEX DAQ restarted
250618 11/25/08 00:03 S. Sirca Swing shift SUMMARY
250617 11/25/08 00:02 adaq Scaler rates (BigBite) snapshot
Monday
250615 11/24/08 23:57 E.Chudakov Experiment summary for Nov 23(swing) - 24(day)
250614 11/24/08 23:08 Chiranjib VDC TO calibration for the second period of Production
250611 11/24/08 22:52 adaq target state
250608 11/24/08 21:58 S. Sirca Spot size too small
250607 11/24/08 21:53 S. Sirca Spot++ snapshot (w/o overlapping windows)
250606 11/24/08 21:48 S. Sirca Spot++ snapshot
250603 11/24/08 20:20 adaq Target Field and Oven Temps snapshot
250602 11/24/08 20:20 adaq Target Cell Temps snapshot
250601 11/24/08 20:20 adaq Radiation and Ozone snapshot
250600 11/24/08 20:20 adaq Hall A Tools snapshot
250599 11/24/08 20:20 adaq BigBox Magnet snapshot
250598 11/24/08 20:20 adaq Beam Position and Current snapshot
250597 11/24/08 20:20 adaq Beam Energy in A and C snapshot
250596 11/24/08 20:02 adaq Scaler rates (BigBite) snapshot
250591 11/24/08 18:47 Katich Another quick Lumi update
250590 11/24/08 18:19 adaq target state
250589 11/24/08 18:13 adaq target state
250588 11/24/08 18:00 S. Sirca Logging of runs 3871, 3872
250587 11/24/08 17:56 S. Sirca hacsbc2 still in bad shape
250583 11/24/08 16:57 D. Parno Wire chamber -- bubbling again
250582 11/24/08 16:40 S. Sirca hacsbc2 behaves erratically
250581 11/24/08 16:36 S. Sirca Beam energy in A and C snapshot
250580 11/24/08 16:33 S. Sirca BigBite scalers snapshot
250579 11/24/08 16:30 S. Sirca Hall A Tools screen
250578 11/24/08 16:27 S. Sirca Spot++
250576 11/24/08 16:02 adaq Scaler rates (BigBite) snapshot
250574 11/24/08 15:57 gaskell Day Shift Summary
250573 11/24/08 15:51 gaskell Beam off: magnet problem
250571 11/24/08 15:30 brads Dump ion chamber readbacks reappeared
250570 11/24/08 15:11 brads BB Cerenkov ADC vs MWDC postion, vice versa
250567 11/24/08 14:37 Katich Some preliminary lumi anaylsis
250566 11/24/08 14:29 gaskell HALLA:p flatlined - restart BEM server
250565 11/24/08 14:16 V. Sulkosky Hall A Energy Readback Flatlined
250564 11/24/08 14:15 V. Sulkosky Hall A Energy Readback Flatlined
250559 11/24/08 12:33 D. Parno Left-arm VDC efficiency
250558 11/24/08 12:30 Jin Huang Issues remains & Full Scalar consistency check added to online plots
250557 11/24/08 12:20 adaq Target Field and Oven Temps snapshot
250556 11/24/08 12:20 adaq Target Cell Temps snapshot
250555 11/24/08 12:20 adaq Radiation and Ozone snapshot
250554 11/24/08 12:19 adaq Hall A Tools snapshot
250553 11/24/08 12:19 adaq BigBox Magnet snapshot
250552 11/24/08 12:19 adaq Beam Position and Current snapshot
250551 11/24/08 12:19 adaq Beam Energy in A and C snapshot
250550 11/24/08 12:06 xqian charge 2.76C
250549 11/24/08 12:02 adaq Scaler rates (BigBite) snapshot
250548 11/24/08 11:56 gaskell Hall A tools
250547 11/24/08 11:54 gaskell Spot++
250545 11/24/08 11:47 gaskell sprinkler system test
250544 11/24/08 11:46 lguo wire chamber
250541 11/24/08 11:17 gaskell beam was gone for RF problem
250539 11/24/08 10:28 D. Parno Shift in BPM x
250536 11/24/08 09:31 yawei preshower and shower energy check, run#3864
250535 11/24/08 09:01 V. Sulkosky Re: Increase in Accumulated Charge
250532 11/24/08 08:36 Moffit Bad BCM dx3 signal into RightHRS/happex7
250531 11/24/08 08:09 Markowitz Owl shift summary
250529 11/24/08 08:02 adaq Scaler rates (BigBite) snapshot
250528 11/24/08 07:59 bradshaw Laser Absorption Spect- Anti-Parallel B and S
250527 11/24/08 07:52 adaq target: Spin Flip Report: TO Report - Owl Shift
250522 11/24/08 06:24 adaq bcm & tools screenshots
250521 11/24/08 05:31 Jin Huang Coda Freeze @ 5:06AM
250517 11/24/08 04:19 adaq Target Field and Oven Temps snapshot
250516 11/24/08 04:19 adaq Target Cell Temps snapshot
250515 11/24/08 04:19 adaq Radiation and Ozone snapshot
250514 11/24/08 04:19 adaq Hall A Tools snapshot
250513 11/24/08 04:19 adaq BigBox Magnet snapshot
250512 11/24/08 04:19 adaq Beam Position and Current snapshot
250511 11/24/08 04:19 adaq Beam Energy in A and C snapshot
250510 11/24/08 04:02 adaq Scaler rates (BigBite) snapshot
250507 11/24/08 03:41 Markowitz bcm & tools screenshots
250500 11/24/08 00:04 JEn-Chieh Pe shift summary
250498 11/24/08 00:02 adaq Scaler rates (BigBite) snapshot
Sunday
250495 11/23/08 23:02 Jin Huang RICH is integrated into Main Replay
250494 11/23/08 21:58 Peng Beam is back
250492 11/23/08 21:37 Jin MCC called for 10min beam down
250490 11/23/08 21:13 Peng Dump IOC
250489 11/23/08 21:08 Peng Dump IOC
250488 11/23/08 20:54 brads Dump ion chamber readbacks are zero. IOC dead?
250485 11/23/08 20:19 adaq Target Field and Oven Temps snapshot
250484 11/23/08 20:19 adaq Target Cell Temps snapshot
250483 11/23/08 20:19 adaq Radiation and Ozone snapshot
250482 11/23/08 20:19 adaq Hall A Tools snapshot
250481 11/23/08 20:18 adaq BigBox Magnet snapshot
250480 11/23/08 20:18 adaq Beam Position and Current snapshot
250479 11/23/08 20:18 adaq Beam Energy in A and C snapshot
250478 11/23/08 20:02 adaq Scaler rates (BigBite) snapshot
250477 11/23/08 19:34 adaq target: Spin Flip Report: 300 Successful Flips Celebration!
250474 11/23/08 19:01 ycwang Mauro Iodice finds the kaon's ring for the rich at run 3799
250469 11/23/08 16:55 Peng Hall A General Tool
250468 11/23/08 16:21 E.Chudakov Experiment summary for Nov 22(swing) - 23(day)
250467 11/23/08 16:08 Bosted Day shift summary
250465 11/23/08 16:02 adaq Scaler rates (BigBite) snapshot
250464 11/23/08 15:57 xqian Two cards in BigBite Wire Chamber need to be replaced.
250461 11/23/08 15:36 V. Sulkosky Re: Ignore Trigger Download
250460 11/23/08 15:26 atrig Trigger-Download
250459 11/23/08 15:26 adaq target: Spin Flip Report: Target log day shift
250458 11/23/08 15:25 dalton Target stripchart screenshot
250457 11/23/08 14:52 D. Parno HAPPEX crashed. New HAPPEX run: 30450
250450 11/23/08 13:37 D. Parno 3845 is production
250448 11/23/08 13:35 Kalyan Run 3844 is production
250445 11/23/08 13:25 yawei preshower and shower energy check, run# 3841
250444 11/23/08 12:24 Kalyan Run 3843 : pedestals for BigBite
250443 11/23/08 12:23 D. Parno New HAPPEX run: 30449
250440 11/23/08 12:18 adaq Target Field and Oven Temps snapshot
250439 11/23/08 12:18 adaq Target Cell Temps snapshot
250438 11/23/08 12:18 adaq Radiation and Ozone snapshot
250437 11/23/08 12:18 adaq Hall A Tools snapshot
250436 11/23/08 12:18 adaq BigBox Magnet snapshot
250435 11/23/08 12:18 adaq Beam Position and Current snapshot
250434 11/23/08 12:18 adaq Beam Energy in A and C snapshot
250429 11/23/08 12:02 adaq Scaler rates (BigBite) snapshot
250428 11/23/08 11:52 D. Parno L-arm VDC efficiency
250423 11/23/08 09:45 D. Parno New HAPPEX run: 30448
250422 11/23/08 09:28 Jin run 3837 Replay is OK
250421 11/23/08 09:24 Bosted Tools screen snapshot
250420 11/23/08 09:17 Bosted Beam spot size
250419 11/23/08 09:09 Bosted CODA reboot
250413 11/23/08 08:09 Markowitz Owl shift summary
250410 11/23/08 08:02 adaq Scaler rates (BigBite) snapshot
250409 11/23/08 07:58 Yi Zhang Snapshot of target temperature
250406 11/23/08 07:43 Yi Zhang laser spectrum
250400 11/23/08 04:46 adaq target: Spin Flip Report: owl shift
250399 11/23/08 04:18 adaq Target Field and Oven Temps snapshot
250398 11/23/08 04:18 adaq Target Cell Temps snapshot
250397 11/23/08 04:18 adaq Radiation and Ozone snapshot
250396 11/23/08 04:18 adaq Hall A Tools snapshot
250395 11/23/08 04:18 adaq BigBox Magnet snapshot
250394 11/23/08 04:18 adaq Beam Position and Current snapshot
250393 11/23/08 04:17 adaq Beam Energy in A and C snapshot
250390 11/23/08 04:06 Markowitz poor beam quality (bcm & tools screenshots)
250389 11/23/08 04:02 adaq Scaler rates (BigBite) snapshot
250384 11/23/08 01:05 Markowitz spot++ screenshot
250383 11/23/08 00:47 Markowitz bcm screenshot
250381 11/23/08 00:20 E.Chudakov Experiment summary for Nov 21(swing) - 22(day)
250379 11/23/08 00:01 adaq Scaler rates (BigBite) snapshot
Saturday
250378 11/22/08 23:58 Jen-Chieh Pe shift summary
250376 11/22/08 23:33 Chiranjib Polarization / spin state status
250367 11/22/08 20:17 adaq Target Field and Oven Temps snapshot
250366 11/22/08 20:17 adaq Target Cell Temps snapshot
250365 11/22/08 20:17 adaq Radiation and Ozone snapshot
250364 11/22/08 20:17 adaq Hall A Tools snapshot
250363 11/22/08 20:17 adaq BigBox Magnet snapshot
250362 11/22/08 20:17 adaq Beam Position and Current snapshot
250361 11/22/08 20:17 adaq Beam Energy in A and C snapshot
250360 11/22/08 20:01 adaq Scaler rates (BigBite) snapshot
250357 11/22/08 19:24 adaq target: Spin Flip Report: spin flip system control panel
250356 11/22/08 19:15 yawei spin flip system control panel
250355 11/22/08 19:06 yawei screen shot laser absorption spectrum
250352 11/22/08 18:09 Kalyan fast clock (103.7kHz) in every event
250351 11/22/08 17:46 yawei preshower and shower energy check, run#3821
250348 11/22/08 16:44 Yawei Hall A general tools
250345 11/22/08 16:01 adaq Scaler rates (BigBite) snapshot
250343 11/22/08 15:55 Simonetti Day shift summary
250342 11/22/08 15:49 dalton Laser spectrum
250341 11/22/08 15:40 adaq target: Spin Flip Report: Day shift log
250340 11/22/08 15:20 D. Parno Run 3819 ended
250338 11/22/08 15:02 Kalyan BCM u3 and d3 scaler info for every event
250335 11/22/08 13:40 D. Parno New HAPPEX run: 30443
250332 11/22/08 13:30 A. Camsonne Compton Electron detector curent
250331 11/22/08 13:01 Simonetti Beam tuning ok
250330 11/22/08 12:41 Simonetti Beam position tuning
250328 11/22/08 12:30 Simonetti Vacuum problem solved
250327 11/22/08 12:21 Ewing Left spect. vacuum
250326 11/22/08 12:18 dalton Optics target in
250325 11/22/08 12:17 adaq Target Field and Oven Temps snapshot
250324 11/22/08 12:17 adaq Target Cell Temps snapshot
250323 11/22/08 12:17 adaq Radiation and Ozone snapshot
250322 11/22/08 12:17 adaq Hall A Tools snapshot
250321 11/22/08 12:17 adaq BigBox Magnet snapshot
250320 11/22/08 12:17 adaq Beam Position and Current snapshot
250319 11/22/08 12:17 adaq Beam Energy in A and C snapshot
250318 11/22/08 12:01 adaq Scaler rates (BigBite) snapshot
250317 11/22/08 11:49 Simonetti Controlled access request to check LHRS vacuum problem
250313 11/22/08 11:10 xqian BigBite Gas Cerenkov Study II ----- ADC
250312 11/22/08 10:52 xqian Hitting efficiency
250311 11/22/08 10:42 Simonetti Tech on call, Todd Ewing, will come at 12:00 to check the vacuum problem
250310 11/22/08 10:27 xqian BigBite Gas Cerenkov Study I ---- TDC
250309 11/22/08 10:24 Simonetti Call page number of "Tech on call" about vacuum
250306 11/22/08 10:09 Simonetti LHRS vacuum pressure
250305 11/22/08 09:29 dalton Target screenshot
250304 11/22/08 09:23 Jin 1% inconsistency between T3 scalar counts in bigbite and LHRS
250303 11/22/08 09:18 D. Parno Beam Position
250302 11/22/08 09:14 Simonetti LHRS vacuum pressure
250301 11/22/08 09:13 D. Parno New HAPPEX run: 30442
250298 11/22/08 08:58 D. Parno Left-arm VDC efficiency
250295 11/22/08 08:01 adaq Scaler rates (BigBite) snapshot
250294 11/22/08 08:01 Markowitz Owl shift summary
250292 11/22/08 07:48 Vilardi Laser Absorption Spectrum
250291 11/22/08 07:48 adaq target: Spin Flip Report: Halor Report owl shift 11/22/2008
250288 11/22/08 06:50 Jin LHRS vacuum followup
250287 11/22/08 06:39 Markowitz LHRS vacuum slowly worsens
250284 11/22/08 05:25 Jin re:re: HAPPEX run#
250281 11/22/08 05:12 Chiranjib HAPPEX DAQ restarted.
250276 11/22/08 04:16 adaq Target Field and Oven Temps snapshot
250275 11/22/08 04:16 adaq Target Cell Temps snapshot
250274 11/22/08 04:16 adaq Radiation and Ozone snapshot
250273 11/22/08 04:16 adaq Hall A Tools snapshot
250272 11/22/08 04:16 adaq BigBox Magnet snapshot
250271 11/22/08 04:16 adaq Beam Position and Current snapshot
250270 11/22/08 04:16 adaq Beam Energy in A and C snapshot
250269 11/22/08 04:07 R. Michaels re HAPPEX run#
250268 11/22/08 04:01 adaq Scaler rates (BigBite) snapshot
250263 11/22/08 02:31 Markowitz LHRS vacuum followup
250262 11/22/08 02:08 Jin CODA not working @ run 3804
250258 11/22/08 01:42 Markowitz RICH screenshots
250257 11/22/08 01:41 Markowitz hall A tools &bb screen
250256 11/22/08 01:39 Markowitz bcm screenshot
250253 11/22/08 00:05 Simonetti Beam problem solved
250252 11/22/08 00:02 Simonetti Swing shift summary
250249 11/22/08 00:01 adaq Scaler rates (BigBite) snapshot
Friday
250248 11/21/08 23:48 Simonetti LHSR vacuum pressure
250247 11/21/08 23:45 Simonetti Beam problem
250243 11/21/08 23:08 adaq target: Spin Flip Report: Swing shift log
250240 11/21/08 21:40 A. Camsonne Happex run number in regular DAQ start of run
250239 11/21/08 21:17 D. Parno New HAPPEX run: 30439
250234 11/21/08 20:28 Simonetti Beam problem
250233 11/21/08 20:16 adaq Target Field and Oven Temps snapshot
250232 11/21/08 20:16 adaq Target Cell Temps snapshot
250231 11/21/08 20:16 adaq Radiation and Ozone snapshot
250230 11/21/08 20:16 adaq Hall A Tools snapshot
250229 11/21/08 20:16 adaq BigBox Magnet snapshot
250228 11/21/08 20:16 adaq Beam Position and Current snapshot
250227 11/21/08 20:16 adaq Beam Energy in A and C snapshot
250224 11/21/08 20:01 adaq Scaler rates (BigBite) snapshot
250221 11/21/08 19:47 ycwang Pions' Ring for RICH at runnumber 3782'
250218 11/21/08 18:55 V. Sulkosky T0 Determined for Run 3789
250217 11/21/08 18:48 V. Sulkosky VDC T0 Calibration
250216 11/21/08 18:46 adaq target screenshot
250215 11/21/08 18:23 E.Chudakov Experiment summary for Nov 20(swing) - 21(day)
250214 11/21/08 18:12 Jin Golden Run -> 3787
250211 11/21/08 17:12 D. Parno Left-arm VDC efficiency
250210 11/21/08 17:07 D. Parno New HAPPEX run: 30438
250207 11/21/08 16:05 F. GaribaldI Day shift summary
250205 11/21/08 16:01 adaq Scaler rates (BigBite) snapshot
250202 11/21/08 15:00 brads BB Cerenkov rates with front added front shielding
250199 11/21/08 14:28 Holmstrom spot++
250198 11/21/08 13:47 xqian What I plotted is a preshower ADC
250197 11/21/08 13:37 xqian Another plot
250196 11/21/08 13:21 Yi Qiang A Typo .... :S
250194 11/21/08 13:15 Yi Qiang Some thoughts on the gas Cherenkov
250193 11/21/08 13:15 ellie HAPPEX Restarted, Run 30437
250191 11/21/08 12:32 Yi Zhang analysis on BB Cherenkov signal Fig3
250190 11/21/08 12:30 Yi Zhang analysis on BB Cherenkov signal
250189 11/21/08 12:16 adaq Target Field and Oven Temps snapshot
250188 11/21/08 12:16 adaq Target Cell Temps snapshot
250187 11/21/08 12:15 adaq Radiation and Ozone snapshot
250186 11/21/08 12:15 adaq Hall A Tools snapshot
250185 11/21/08 12:15 adaq BigBox Magnet snapshot
250184 11/21/08 12:15 adaq Beam Position and Current snapshot
250183 11/21/08 12:15 adaq Beam Energy in A and C snapshot
250180 11/21/08 12:01 adaq Scaler rates (BigBite) snapshot
250179 11/21/08 11:08 xqian The HV trip, and run starts before the chamber completely recover
250176 11/21/08 11:02 adaq target: Spin Flip Report: Stepper Motor #1423 warning
250175 11/21/08 11:01 V. Sulkosky Re: U10 BCM Double Pulsing (fixed)
250174 11/21/08 10:29 ycwang Good data for the RICH after turned on HV(2200V) from last night
250173 11/21/08 10:11 yawei preshower and shower energy check, run# 3785
250172 11/21/08 10:07 yawei preshower and shower energy check
250171 11/21/08 10:04 E.Chudakov Beam is back
250169 11/21/08 09:38 Correction of the time
250168 11/21/08 09:36 Correction
250167 11/21/08 09:12 E.Chudakov The compressor is turned off
250166 11/21/08 09:11 ellie HAPPEX Restarted, Run 30436
250165 11/21/08 09:09 E.Chudakov 1/2 wave plate is out
250163 11/21/08 08:45 A. Camsonne RICH data
250162 11/21/08 08:30 E.Chudakov Control access to address the smoke alert
250159 11/21/08 08:16 Moffit Re: U10 Double pulsing.
250158 11/21/08 08:15 xqian The Wire Chamber is not working at the start of this run
250156 11/21/08 08:01 adaq Scaler rates (BigBite) snapshot
250155 11/21/08 07:46 Markowitz Owl shift summary
250153 11/21/08 07:43 Vilardi La
250151 11/21/08 07:42 adaq target: Spin Flip Report: Halog Report Owl shift 11/21/2008
250150 11/21/08 07:32 adaq hitting efficiency for 3771
250149 11/21/08 07:26 R. Michaels re: CRL for RICH
250148 11/21/08 07:11 adaq tools screen
250147 11/21/08 06:37 adev BigBite WireChamber Measured Voltage change!
250146 11/21/08 06:36 Yi Zhang BB Chamber HV tripped
250144 11/21/08 06:27 R. Michaels how to back out (RICH DAQ change)
250143 11/21/08 06:26 adev BigBite WireChamber Measured Voltage change!
250137 11/21/08 05:10 xqian Negative reading in scalers
250136 11/21/08 04:50 Yi Zhang BB Cherenkov TDC Signal
250135 11/21/08 04:15 adaq Target Field and Oven Temps snapshot
250134 11/21/08 04:15 adaq Target Cell Temps snapshot
250133 11/21/08 04:15 adaq Radiation and Ozone snapshot
250132 11/21/08 04:15 adaq Hall A Tools snapshot
250131 11/21/08 04:15 adaq BigBox Magnet snapshot
250130 11/21/08 04:15 adaq Beam Position and Current snapshot
250129 11/21/08 04:15 adaq Beam Energy in A and C snapshot
250128 11/21/08 04:05 Chiranjib HAPPEX restarted
250126 11/21/08 04:01 adaq Scaler rates (BigBite) snapshot
250118 11/21/08 01:08 Chiranjib Internal Temperature test
250114 11/21/08 00:03 dhamilton Swing Shift Summary
250113 11/21/08 00:01 adaq Scaler rates (BigBite) snapshot
Thursday
250112 11/20/08 23:40 deng screen shot of Lumis check for run 30433
250111 11/20/08 23:38 dhamilton RE: Hall A VESDAs....
250110 11/20/08 23:25 Jin HAPPEX 30433 started
250109 11/20/08 23:18 Kalyan BCM U10 seems to be double pulsing again on LHRS
250106 11/20/08 22:54 dhamilton VESDA Warnings
250105 11/20/08 22:43 Chiranjib The RUN number for the "dead time issue" was 3774
250104 11/20/08 22:39 Chiranjib target spin flip screens / Laser absorption etc.
250103 11/20/08 22:37 Chiranjib Target screens
250102 11/20/08 22:36 dhamilton Beam position unstable around DAQ crash
250101 11/20/08 22:34 Chiranjib DEAD TIME PROBLEM OCCURED / ROCs REBOOTED
250099 11/20/08 21:59 LeRose Target: Laser Absorption Spectrum
250098 11/20/08 21:57 deng screen shot of lumi check for run 3774
250095 11/20/08 21:40 deng screen shot of run 3773
250094 11/20/08 21:32 ycwang RICH HV turned on, try the data after run # 3774
250093 11/20/08 21:22 ycwang HV tripped for RICH, probably that is the reason why data is not reasonable.
250092 11/20/08 21:18 Kalyan re: CRL for RICH
250091 11/20/08 21:13 dhamilton Hall A Tools
250090 11/20/08 21:13 ycwang One ADC has too much noise and there is something not right for new data taking from RICH ?
250089 11/20/08 21:03 R. Michaels re: CRL for RICH
250086 11/20/08 20:30 deng screen shot of run 3772
250085 11/20/08 20:15 adaq Target Field and Oven Temps snapshot
250084 11/20/08 20:15 adaq Target Cell Temps snapshot
250083 11/20/08 20:15 adaq Radiation and Ozone snapshot
250082 11/20/08 20:14 adaq Hall A Tools snapshot
250081 11/20/08 20:14 adaq BigBox Magnet snapshot
250080 11/20/08 20:14 adaq Beam Position and Current snapshot
250079 11/20/08 20:14 adaq Beam Energy in A and C snapshot
250078 11/20/08 20:00 adaq Scaler rates (BigBite) snapshot
250075 11/20/08 19:16 deng screen shot of coda and deadtime of run 3771
250074 11/20/08 18:52 Kalyan End run failed for 3770
250073 11/20/08 18:46 Kalyan Deadtime issue with DAQ
250071 11/20/08 18:07 adev BigBite WireChamber Measured Voltage change!
250070 11/20/08 17:55 adev BigBite WireChamber Measured Voltage change!
250067 11/20/08 17:51 ycwang A1 rates with 12 uA beam
250064 11/20/08 17:48 kalyan Trigger rate at 12uA on 3He
250063 11/20/08 17:47 dhamilton Begin Production Running
250062 11/20/08 17:42 E.Chudakov Experiment: summary for Nov 19 (swing) - 20 (day)
250061 11/20/08 17:33 LeRose 3He target in
250057 11/20/08 17:16 A. Camsonne moved CRL file to BOB's version
250056 11/20/08 17:05 adev BigBite WireChamber Measured Voltage change!
250035 11/20/08 16:21 LeRose Target BeO
250031 11/20/08 16:11 Jack (Techs) Left Q1 power supply
250029 11/20/08 16:04 Saha Shift Summary
250028 11/20/08 16:03 LeRose Left Q2/3 Cycled
250027 11/20/08 16:00 adaq Scaler rates (BigBite) snapshot
250026 11/20/08 15:58 Simonetti target moved to empty target
250025 11/20/08 15:31 fansler/Ewin lead flow
250024 11/20/08 15:26 A. Camsonne Edet event structure
250023 11/20/08 15:24 ycwang new pedestal for rich at run 3753
250011 11/20/08 14:46 Moffit Cleanup of some bcm cables in counting house. Mainly Dx3 and Dx10.
249984 11/20/08 13:48 brads Pb-glass test PMTs signal & HV channels
249978 11/20/08 13:31 simonetti Glitch during preparation of field
249977 11/20/08 13:30 Yi Qiang NMR Frequency/Field Sweep Cross Calibration for Vertical Pumping
249976 11/20/08 12:59 Bob/Ramesh R-HRS activity
249971 11/20/08 12:39 E.Chudakov Access status
249969 11/20/08 12:18 Michel Bross Electron detector dark current
249967 11/20/08 12:14 adaq Target Field and Oven Temps snapshot
249966 11/20/08 12:14 adaq Target Cell Temps snapshot
249965 11/20/08 12:14 adaq Radiation and Ozone snapshot
249964 11/20/08 12:14 adaq Hall A Tools snapshot
249963 11/20/08 12:14 adaq BigBox Magnet snapshot
249962 11/20/08 12:14 adaq Beam Position and Current snapshot
249961 11/20/08 12:14 adaq Beam Energy in A and C snapshot
249960 11/20/08 12:00 adaq Scaler rates (BigBite) snapshot
249959 11/20/08 11:58 Yi Qiang/YI EPR Calibration for vertical+ field, anti-parallel pumping
249957 11/20/08 11:10 brads BB Cerenkov witness PMT: shielding added in front
249956 11/20/08 10:56 V. Sulkosky Re: U10 Double Pulsing Checked
249954 11/20/08 10:41 Kalyan High dead time went away after I restarted coda
249950 11/20/08 10:27 A. Camsonne High dead time with T1 even without RICH
249943 11/20/08 10:15 A. Camsonne Serial adapter
249938 11/20/08 10:00 E.Chudakov Schedule for the access
249927 11/20/08 09:23 E.Chudakov No unwanted dispersive mode on Nov 17 - another followup
249924 11/20/08 08:53 Jin Huang Polarization History for Vertical Pumping Production Run
249923 11/20/08 08:39 E.Chudakov No unwanted dispersive mode on Nov 17
249922 11/20/08 08:18 Jin Huang AFP Loss Test @ Vert Pumping, 70% Pol & No Correction Coil
249921 11/20/08 08:15 Yi Zhang BB Cherenkov testing summary
249920 11/20/08 08:13 Yi Zhang BB Cherenkov testing summary
249918 11/20/08 08:00 adaq Scaler rates (BigBite) snapshot
249917 11/20/08 07:35 gilman owl shift summary
249916 11/20/08 07:31 gilman ask for controlled access
249913 11/20/08 07:09 gilman arc shunt problem
249912 11/20/08 06:59 Jin Huang Target Data is daily backed up to central Tape
249911 11/20/08 06:59 Jin Huang Coincident Cut applied on physics replay
249910 11/20/08 06:59 Jin Huang Protection on Golden Root File
249909 11/20/08 06:59 Jin Huang Online Plots divided to Shift Worker plots and Expert Plots
249907 11/20/08 06:52 gilman magnet off
249902 11/20/08 06:14 Moffit HAPPEX DAQ: Please keep runs less than 4 hours long.
249897 11/20/08 05:30 gilman spot++ for shift checklist
249896 11/20/08 05:26 gilman Hall A Gen Tools for shift checklist
249891 11/20/08 04:59 Jin Scalar & BCM for run 3712
249888 11/20/08 04:19 Jin Scalar & BCM for run 3711
249886 11/20/08 04:14 adaq Target Field and Oven Temps snapshot
249885 11/20/08 04:14 adaq Target Cell Temps snapshot
249884 11/20/08 04:14 adaq Radiation and Ozone snapshot
249883 11/20/08 04:14 adaq Hall A Tools snapshot
249882 11/20/08 04:13 adaq BigBox Magnet snapshot
249881 11/20/08 04:13 adaq Beam Position and Current snapshot
249880 11/20/08 04:13 adaq Beam Energy in A and C snapshot
249878 11/20/08 04:10 adev BigBite WireChamber Measured Voltage change!
249877 11/20/08 04:00 adaq Scaler rates (BigBite) snapshot
249876 11/20/08 03:59 adev BigBite WireChamber Measured Voltage change!
249875 11/20/08 03:48 gilman Moller done
249874 11/20/08 03:46 E.Chudakov Moller results
249859 11/20/08 01:27 A. Camsonne Trigger rates during moeller
249858 11/20/08 01:26 A. Camsonne Dead time with RICH
249857 11/20/08 01:21 A. Camsonne RICH readout change
249844 11/20/08 00:58 E.Chudakov Moller start
249841 11/20/08 00:42 A. Camsonne Compton electron detector current
249840 11/20/08 00:37 A. Camsonne Compton power meter channel in odd state
249838 11/20/08 00:27 adev BigBite WireChamber Measured Voltage change!
249834 11/20/08 00:20 A. Camsonne Compton power meter channel in odd state
249833 11/20/08 00:02 dhamilton Swing Shift Summary
249831 11/20/08 00:00 adaq Scaler rates (BigBite) snapshot
Wednesday
249818 11/19/08 23:38 brads BB Cerenkov tests for tonight
249814 11/19/08 22:56 E.Chudakov No Left Q1 for tonight
249811 11/19/08 22:24 adaq target: Spin Flip Report: Shift target report
249798 11/19/08 21:26 E.Chudakov Q1 problem
249793 11/19/08 20:32 E.Chudakov Experiment: summary for 11/18(day)-11/19(day)
249790 11/19/08 20:15 dhamilton Hall moved to controlled access
249788 11/19/08 20:13 adaq Target Field and Oven Temps snapshot
249787 11/19/08 20:13 adaq Target Cell Temps snapshot
249786 11/19/08 20:13 adaq Radiation and Ozone snapshot
249785 11/19/08 20:13 adaq Hall A Tools snapshot
249784 11/19/08 20:13 adaq BigBox Magnet snapshot
249783 11/19/08 20:13 adaq Beam Position and Current snapshot
249782 11/19/08 20:13 adaq Beam Energy in A and C snapshot
249781 11/19/08 20:00 adaq Scaler rates (BigBite) snapshot
249780 11/19/08 19:46 dhamilton Q1 again - contacted Jack
249779 11/19/08 19:36 dhamilton Hall moved to Beam Permit
249778 11/19/08 18:55 dhamilton Hall moved to controlled access
249777 11/19/08 18:48 E.Chudakov Problems with the accelerator and with Q1
249776 11/19/08 18:44 dhamilton Status update
249772 11/19/08 18:32 dhamilton Q1 Status Update
249770 11/19/08 18:00 dhamilton Another Q1 Quench
249769 11/19/08 17:56 xqian Coincidence events in BigBite
249767 11/19/08 17:38 Kalyan Scintillator rate on BigBite
249765 11/19/08 17:24 dhamilton Q1 back up to 2.35 GeV/c
249763 11/19/08 17:02 dhamilton Q1 Left Quenched again
249760 11/19/08 16:20 dhamilton Compton screenshot
249759 11/19/08 16:19 dhamilton BPM/BCM Screenshots
249758 11/19/08 16:17 dhamilton Q1 is back up
249755 11/19/08 16:00 adaq Scaler rates (BigBite) snapshot
249754 11/19/08 16:00 adaq target: Spin Flip Report: shift day 19-11
249753 11/19/08 15:59 gilman day shift summary - 19 Nov 2008
249752 11/19/08 15:52 brads BB detector stack's view of the target
249751 11/19/08 15:49 gilman Q1 trip / quench
249747 11/19/08 15:20 Kalyan Re: U10 Double-pulsing in Left HRS
249746 11/19/08 15:06 V. Sulkosky U10 Double-pulsing in Left HRS
249745 11/19/08 15:01 Katich BCM U10 is absent from HAPPEX DAQ
249742 11/19/08 14:07 gilman spot for shift checklist
249741 11/19/08 14:05 gilman Hall A Tools for shift checklist
249736 11/19/08 12:15 Katich Lumi HV Change
249735 11/19/08 12:13 adaq Target Field and Oven Temps snapshot
249734 11/19/08 12:13 adaq Target Cell Temps snapshot
249733 11/19/08 12:13 adaq Radiation and Ozone snapshot
249732 11/19/08 12:13 adaq Hall A Tools snapshot
249731 11/19/08 12:13 adaq BigBox Magnet snapshot
249730 11/19/08 12:13 adaq Beam Position and Current snapshot
249729 11/19/08 12:12 adaq Beam Energy in A and C snapshot
249728 11/19/08 12:00 adaq Scaler rates (BigBite) snapshot
249727 11/19/08 11:55 gilman beam returns ~11:43
249725 11/19/08 11:44 Moffit More Run Database (mysql) changes.
249724 11/19/08 11:29 xqian Beam Charge Counting now
249723 11/19/08 11:12 gilman 11 AM update
249721 11/19/08 10:33 Ole Hansen adaql10 diagnostic
249720 11/19/08 10:20 gilman beam coming back
249719 11/19/08 10:15 gilman logbook end of run error
249718 11/19/08 10:03 V. Sulkosky S2m Left # 4 ADC Appears Fine Now
249717 11/19/08 10:03 gilman ioc issue
249716 11/19/08 10:02 fansler LHRS vacuum
249714 11/19/08 09:47 V. Sulkosky Golden Root files Set to 3661
249713 11/19/08 09:46 adaq target: Spin Flip Report: after a short control access
249712 11/19/08 09:42 gilman controlled access done
249711 11/19/08 09:30 cisbani RICH rings cumulated
249709 11/19/08 09:10 gilman running nicely - sort off - controlled access
249708 11/19/08 08:58 X. Jiang MCC is planning a RF recovery tomorow, 8:30am, four hours
249705 11/19/08 08:15 Moffit MYSQL Run Database... added new Table.
249704 11/19/08 08:13 xqian Wire Chamber hitting efficiency per wire
249703 11/19/08 08:11 Yi Zhang /adaql10/work1/e06010/ROOTfiles/ problem sloved
249702 11/19/08 08:04 Simonetti Owl shift summary
249701 11/19/08 08:03 Simonetti LHSR vacuum pressure update
249699 11/19/08 08:00 adaq Scaler rates (BigBite) snapshot
249697 11/19/08 07:53 Vilardi Pumping laser Absorption spectrum
249695 11/19/08 07:50 adaq target: Spin Flip Report: Halog Report Owl shift 11/19/2008
249694 11/19/08 07:49 Yi Zhang /adaql10/work1/e06010/ROOTfiles/ access problem
249693 11/19/08 07:42 Moffit Lumi 7 and 8 Saturated
249692 11/19/08 07:32 xqian live time without RICH is 91%
249691 11/19/08 07:24 xqian BigBite Wire Chamber hitting efficiency
249688 11/19/08 06:22 Simonetti LHSR vacuum pressure update
249685 11/19/08 05:34 Simonetti LHRS vacuum pressure update
249682 11/19/08 05:01 Simonetti LHRS vacuum pressure read out
249679 11/19/08 04:12 adaq Target Field and Oven Temps snapshot
249678 11/19/08 04:12 adaq Target Cell Temps snapshot
249677 11/19/08 04:12 adaq Radiation and Ozone snapshot
249676 11/19/08 04:12 adaq Hall A Tools snapshot
249675 11/19/08 04:12 adaq BigBox Magnet snapshot
249674 11/19/08 04:12 adaq Beam Position and Current snapshot
249673 11/19/08 04:12 adaq Beam Energy in A and C snapshot
249672 11/19/08 04:06 Chiranjib LHRS vacuum pressure read out
249671 11/19/08 04:00 adaq Scaler rates (BigBite) snapshot
249668 11/19/08 03:17 Chiranjib We are running TWOARM FB
249663 11/19/08 01:43 Chiranjib LHRS vacuum pressure (or something like that) read out
249662 11/19/08 01:36 brads BB Cerenkov witness PMT rates with Pb surround shield
249661 11/19/08 01:26 Chiranjib RUN 3653 issue
249658 11/19/08 00:59 Chiranjib Trigger rates/ Prescales @12 uA
249657 11/19/08 00:48 Simonetti Helium alarm
249651 11/19/08 00:37 Jin Trigger @ 12uA, 3x4 raster
249650 11/19/08 00:35 Jin Spot++ @ 6x6 raster, go production beam
249647 11/19/08 00:31 Simonetti raster 5x5
249646 11/19/08 00:31 Jin Trigger @ 6x6 2uA
249643 11/19/08 00:28 Jin Trigger @ 5x5 2uA
249641 11/19/08 00:00 adaq Scaler rates (BigBite) snapshot
Tuesday
249640 11/18/08 23:57 A.Rakhman 11/18/2008 Swing Shift Summary
249639 11/18/08 23:38 Chiranjib Laser parameters
249638 11/18/08 23:26 Chiranjib Target screens II
249637 11/18/08 23:23 Chiranjib TARGET SCREENS
249636 11/18/08 23:19 A.Rakhman beam status
249635 11/18/08 22:57 Jin Trigger @ 3x4, 2uA, he3
249634 11/18/08 22:43 Camsonne / C RICH Activity during controlled access (Nov 18 2008)
249633 11/18/08 22:21 mcanan Target moved to He-3/ started spin flip
249632 11/18/08 22:15 A.Rakhman helium level alarm
249631 11/18/08 22:07 mcanan target moved to optics
249630 11/18/08 22:01 E.Chudakov Smoke problem in the hall
249629 11/18/08 22:01 A.Rakhman hall a tools snap shots
249628 11/18/08 21:47 brads BB Cerenkov ADC<->TDC mapping problem fixed
249627 11/18/08 21:44 brads BB Cerenkov discriminator calibration
249626 11/18/08 21:41 A.Rakhman hall a status update
249625 11/18/08 21:39 brads Modified shielding around BB Cerenkov witness PMT (iteration 2)
249624 11/18/08 21:37 brads Pump on left Q2 has been turned off, shift crew should watch gauge on monitor
249623 11/18/08 21:32 brads The smoke-puffing pump
249621 11/18/08 20:12 adaq Target Field and Oven Temps snapshot
249620 11/18/08 20:12 adaq Target Cell Temps snapshot
249619 11/18/08 20:12 adaq Radiation and Ozone snapshot
249618 11/18/08 20:12 adaq Hall A Tools snapshot
249617 11/18/08 20:12 adaq BigBox Magnet snapshot
249616 11/18/08 20:12 adaq Beam Position and Current snapshot
249615 11/18/08 20:12 adaq Beam Energy in A and C snapshot
249614 11/18/08 20:09 V. Sulkosky Changed S2m Left PMT # 4 ADC Delay Cable
249613 11/18/08 20:00 adaq Scaler rates (BigBite) snapshot
249612 11/18/08 19:41 A.Rakhman smoke detection alarm
249606 11/18/08 19:31 Jin Huang Target: Laser and RTD status at vertical pumping
249605 11/18/08 19:28 A.Rakhman left arm dipole
249604 11/18/08 19:23 Jack (Techs) Target IOC reboot and target motion - Make sure the beam is OFF
249598 11/18/08 19:12 Yi Qiang Target Activities for field rotation from transverse to vertical
249597 11/18/08 19:10 adaq target: Spin Flip Report: Vertical Pumping Auto Spin Flip Ready!
249596 11/18/08 19:09 A.Rakhman Compton
249593 11/18/08 19:06 Yi Qiang EPR calibration for Vertical Pumping
249592 11/18/08 19:05 A.Rakhman Alarm
249590 11/18/08 18:27 Jack (Techs) Right Q2 LHe level meter
249589 11/18/08 18:22 Jack (Techs) Smoke/haze in the hall - Watch the Left HRS vacuum gauge
249586 11/18/08 17:55 Yi Qiang EPR test w/ correction coil, transverse - field, - pumping
249585 11/18/08 17:53 Katich BCMup3 signal restored to HAPPEX DAQ
249584 11/18/08 17:51 brads BB Cerenkov PMT 5 plugged back into ADC
249583 11/18/08 17:50 brads T7 (BB Cerenkov-only) unplugged from Trigger Supervisor
249582 11/18/08 17:42 E.Chudakov Control access and smoke alert in the hall
249581 11/18/08 17:36 Yi Qiang EPR test w/ correction coil, transverse - field, + pumping
249577 11/18/08 17:22 Yi Qiang EPR Calibration for Transverse (anti_parallel) Pumping w/o correction coil
249576 11/18/08 17:06 Yi Qiang EPR Calibration for Transverse Pumping,
249575 11/18/08 17:05 A.Rakhman alarm
249574 11/18/08 17:02 Yi Qiang AFP Loss Test for Transverse Pumping w/o correction coil
249573 11/18/08 16:43 adev BigBite WireChamber Measured Voltage change!
249572 11/18/08 16:34 adaq target: Spin Flip Report: Target Report Before Rotation to Vertical Field
249571 11/18/08 16:31 adev BigBite WireChamber Measured Voltage change!
249570 11/18/08 16:23 A.Rakhman hall a status
249569 11/18/08 16:03 brads BB Cerenkov witness PMT rates with front shielding
249566 11/18/08 15:59 Saha Shift Summary
249565 11/18/08 15:47 Ole Hansen Shorter runs to avoid file splitting
249564 11/18/08 15:22 adaq Scaler rates (BigBite) snapshot
249561 11/18/08 14:24 Lingyan Zhu target: polarization graph
249560 11/18/08 14:24 Yi Qiang Spin Flip Error
249559 11/18/08 14:08 adaq target: Spin Flip Report: Auto spin flip continues
249552 11/18/08 12:31 D. Parno Left-arm VDC efficiency
249549 11/18/08 12:14 A. Camsonne Switched to Rich Holmes analyzer for Compton
249548 11/18/08 12:12 adaq Target Field and Oven Temps snapshot
249547 11/18/08 12:11 adaq Target Cell Temps snapshot
249546 11/18/08 12:11 adaq Radiation and Ozone snapshot
249545 11/18/08 12:11 adaq Hall A Tools snapshot
249544 11/18/08 12:11 adaq BigBox Magnet snapshot
249543 11/18/08 12:11 adaq Beam Position and Current snapshot
249542 11/18/08 12:11 adaq Beam Energy in A and C snapshot
249541 11/18/08 12:07 xqian live time now is about 81.5%, We just keep it now
249540 11/18/08 12:01 Moffit HAPPEX DAQ - Moved x10 BCM to adc15_1
249539 11/18/08 11:37 xqian The BigBite Chamber HVs are not adjusted for this run
249538 11/18/08 11:31 xqian Mapping problem for BB Gas Cerenkov
249537 11/18/08 11:27 xqian RUN 3631 has problem with the chamber HV.
249536 11/18/08 11:25 V. Sulkosky Re: Missing S2m ADC
249535 11/18/08 11:24 adev BigBite WireChamber Measured Voltage change!
249534 11/18/08 11:24 cisbani Run 3629 is a RICH pedestal RUN, with beam off
249532 11/18/08 11:13 adev BigBite WireChamber Measured Voltage change!
249530 11/18/08 11:10 Lingyan Zhu laster spectrum
249529 11/18/08 10:58 Jin Trigger @ 12uA, he3, 3x4 raster
249527 11/18/08 10:51 Jin Huang Spot++ @ 12uA, 3x4, he3 production
249525 11/18/08 10:48 V. Sulkosky VDC T0's Versus Time (resubmission)
249524 11/18/08 10:44 Ole Hansen Start Production
249521 11/18/08 10:36 adaq target: Spin Flip Report: Target Flip ON, Ready for production run, 72% Pol.
249519 11/18/08 10:32 Lingyan Zhu spot++ for raster 6*6
249517 11/18/08 10:29 Jin Trigger @ raster 6x6
249516 11/18/08 10:28 Lingyan Zhu spot++ for raster 5*5
249514 11/18/08 10:26 Ole Hansen Start/stop run scripts not appearing
249511 11/18/08 10:24 Lingyan Zhu spot++ for run 3626
249510 11/18/08 10:23 Ole Hansen CODA restart (again)
249509 11/18/08 10:22 Jin Trigger 2 Raster 5x5
249506 11/18/08 10:19 Jin Trgger @ Raster = 4x4
249505 11/18/08 10:17 Jin Trigger ScreenShot 3x3raster
249504 11/18/08 10:13 Jin Trigger 2uA 2x2
249503 11/18/08 10:13 Lingyan Zhu Target Camera Snapshot
249502 11/18/08 10:10 Jin Target -> He3
249501 11/18/08 10:01 Yi Zhang BB Cherenkov TDC peak
249500 11/18/08 09:39 D. Parno L S2m #4 went missing, came back
249495 11/18/08 08:27 Ole Hansen Run 3623 is junk
249492 11/18/08 08:06 Simonetti Owl shift summary
249491 11/18/08 08:05 Simonetti Beam tuning to BeO target
249486 11/18/08 07:51 ycwang Target motion problem has been solved
249484 11/18/08 07:48 Simonetti Target motion problem fixed
249483 11/18/08 07:37 adaq Laser Absorption Spectrum
249482 11/18/08 07:36 Moffit Re: HAPPEX DAQ - fixed... please restart
249481 11/18/08 07:33 Simonetti Pam works to solve target motion problem
249480 11/18/08 07:31 Simonetti Beam problem
249479 11/18/08 07:26 Moffit Re: HAPPEX DAQ - TS1
249478 11/18/08 07:24 adaq target: Spin Flip Report: target stay ar BeO
249476 11/18/08 06:36 Jin Trigger & BCM for run 3620
249473 11/18/08 06:11 Huaibo HAPPEX DAQ problem.
249472 11/18/08 06:08 Huaibo HAPPEX DAQ problem.
249469 11/18/08 05:51 Jin BCM & Trigger ScreenShot for run#3618
249466 11/18/08 05:28 Jin BCM & Trigger Rate for Run3617
249465 11/18/08 05:24 Huaibo triggers for beam current=0.5uA.
249464 11/18/08 05:21 Simonetti Beam tuning, new position for BPMA and BPMB
249462 11/18/08 05:16 Jin Beam Tuning for 0.5uA
249460 11/18/08 04:26 Jin Fake Spin Signal ON
249459 11/18/08 04:11 adaq Target Field and Oven Temps snapshot
249458 11/18/08 04:11 adaq Target Cell Temps snapshot
249457 11/18/08 04:11 adaq Radiation and Ozone snapshot
249456 11/18/08 04:11 adaq Hall A Tools snapshot
249455 11/18/08 04:11 adaq BigBox Magnet snapshot
249454 11/18/08 04:11 adaq Beam Position and Current snapshot
249453 11/18/08 04:11 adaq Beam Energy in A and C snapshot
249452 11/18/08 04:01 jpchen target motion problem
249451 11/18/08 03:52 Fansler/Sega Target motion controls
249450 11/18/08 02:16 Jin Huang hareboot24 is now accessable in counting house front room
249449 11/18/08 00:12 dhamilton Swing shift summary
Monday
249446 11/17/08 23:37 Jin Spin Flip Exit at Spin=-1 State
249444 11/17/08 23:00 Yi Qiang Upexpected target ladder movement
249443 11/17/08 22:56 dhamilton Move Hall to controlled access
249442 11/17/08 22:47 dhamilton Problem with target motion
249441 11/17/08 22:16 adev BigBite WireChamber Measured Voltage change!
249440 11/17/08 22:07 dhamilton BB HV on
249439 11/17/08 22:05 dhamilton Nominal beam position and harp scan complete
249438 11/17/08 21:45 Jin Huang The Scalar Counting Difference Climbing is due to missing on LHRS scalar
249437 11/17/08 21:42 dhamilton Tuning through Compton finished
249436 11/17/08 21:19 Sirish Compton Tune
249435 11/17/08 20:11 adaq Target Field and Oven Temps snapshot
249434 11/17/08 20:11 adaq Target Cell Temps snapshot
249433 11/17/08 20:11 adaq Radiation and Ozone snapshot
249432 11/17/08 20:11 adaq Hall A Tools snapshot
249431 11/17/08 20:10 adaq BigBox Magnet snapshot
249430 11/17/08 20:10 adaq Beam Position and Current snapshot
249429 11/17/08 20:10 adaq Beam Energy in A and C snapshot
249427 11/17/08 19:52 Saha Shift Summary
249426 11/17/08 19:47 E.Chudakov Moller results
249425 11/17/08 19:42 dalton Target: optics target in
249424 11/17/08 17:53 V. Sulkosky VDC T0 Comparison Run 3611
249423 11/17/08 17:36 xqian BB Wire Chamber
249422 11/17/08 17:31 xqian explaination
249421 11/17/08 17:10 dhamilton Dipole back up to 2.35 GeV/c
249420 11/17/08 15:50 gjin laser spectrum of empty target
249419 11/17/08 15:43 A. Camsonne Helium level low
249418 11/17/08 15:38 xqian power of retiming on BigBite
249417 11/17/08 15:24 adev BigBite WireChamber Measured Current change!
249415 11/17/08 15:20 V. Sulkosky Golden Root File Now 3611
249413 11/17/08 15:12 adev BigBite WireChamber Measured Voltage change!
249412 11/17/08 15:01 adev BigBite WireChamber Measured Voltage change!
249409 11/17/08 14:58 brads BB Cerenkov PMT 5 plugged into online scope
249408 11/17/08 14:57 V. Sulkosky S2m ADC # 4 Plot
249407 11/17/08 14:57 xqian explaination
249406 11/17/08 14:56 adaq target: Spin Flip Report: target status after stopping auto flip
249403 11/17/08 14:50 A. Camsonne RICH phase tweaked
249402 11/17/08 14:46 brads Added plate in front of BB Cerenkov witness PMT
249401 11/17/08 14:39 brads Cerenkov rate data taken on Sunday
249399 11/17/08 14:22 A. Camsonne Experiment status and plan, and meeting reminder
249398 11/17/08 14:18 Saha Spot++ for Run3611
249394 11/17/08 14:06 brads Results from the paddles set-up behind the BB Shower
249383 11/17/08 13:29 V. Sulkosky Checked Re-timing Signal with L1A
249382 11/17/08 13:28 brads Photos of added shielding on BB beamline side
249381 11/17/08 13:15 V. Sulkosky Moved S2m left PMT #4 Signal to a New ADC Channel
249380 11/17/08 12:54 V. Sulkosky Run 3604 Pedestal run
249376 11/17/08 12:47 Yi Qiang Comparison of spin up time with different pumping direction
249373 11/17/08 12:25 adaq Fixed the bad scintillator TDC chan
249368 11/17/08 12:10 adaq Target Field and Oven Temps snapshot
249367 11/17/08 12:10 adaq Target Cell Temps snapshot
249366 11/17/08 12:10 adaq Radiation and Ozone snapshot
249365 11/17/08 12:10 adaq Hall A Tools snapshot
249364 11/17/08 12:10 adaq BigBox Magnet snapshot
249363 11/17/08 12:10 adaq Beam Position and Current snapshot
249362 11/17/08 12:10 adaq Beam Energy in A and C snapshot
249355 11/17/08 11:39 Fansler RQ2/LQ1
249353 11/17/08 10:53 Yi Zhang BB Cherenkov test ADC plots
249352 11/17/08 10:52 adaq target: Spin Flip Report: target status before auto spin flip starts
249351 11/17/08 10:35 A. Camsonne Energy measurement complete
249350 11/17/08 10:32 Saha Arc Energy Measurement
249345 11/17/08 09:43 A. Camsonne Access for fixing cryo for spectrometer
249344 11/17/08 09:34 V. Sulkosky Left HRS Scaler Detector Rates Zero
249341 11/17/08 09:05 adev SCALER = -1
249338 11/17/08 08:58 adev SCALER = -1
249336 11/17/08 08:48 adaq target: Spin Flip Report: test
249332 11/17/08 08:01 Cisbani Owl Shift Summary
249331 11/17/08 08:01 Simonetti Laser spectrum and spin flip screen
249330 11/17/08 07:54 adaq target: Spin Flip Report: End Shift, Spin flip report
249329 11/17/08 07:43 cisbani BB Cherenkov tests
249326 11/17/08 07:09 Vilardi Missing channel, Left Arm S2 back again
249325 11/17/08 07:09 Vilardi Missing channel, Left Arm S2 back again
249320 11/17/08 05:44 vilardi I. Missing channel
249319 11/17/08 05:17 cisbani ROC5, 8, 9, 10 and 15 not responding
249315 11/17/08 04:42 adaq target: Spin Flip Report: Pol. Stablized to ~66%
249314 11/17/08 04:36 Jin Correction : Pol. History of Session # 241
249313 11/17/08 04:25 Jin Huang Target Polarization History during Production
249312 11/17/08 04:10 adaq Target Field and Oven Temps snapshot
249311 11/17/08 04:10 adaq Target Cell Temps snapshot
249310 11/17/08 04:10 adaq Radiation and Ozone snapshot
249309 11/17/08 04:10 adaq Hall A Tools snapshot
249308 11/17/08 04:10 adaq BigBox Magnet snapshot
249307 11/17/08 04:10 adaq Beam Position and Current snapshot
249306 11/17/08 04:10 adaq Beam Energy in A and C snapshot
249303 11/17/08 03:42 adaq Hall A Tools
249290 11/17/08 00:01 D. Dutta Shift Summary
249288 11/17/08 00:00 narayan Target Polarization
249287 11/17/08 00:00 adaq target: Spin Flip Report: Target General Update: the NMR looks bad
Sunday
249286 11/16/08 23:32 xqian Background study at 12 uA
249283 11/16/08 23:15 Yi Qiang Really Masing !!!
249282 11/16/08 23:04 A. Camsonne RICH data for ROC15
249281 11/16/08 23:01 Kalyan re: Dead time issue again,
249280 11/16/08 22:58 Jin Huang Update on scalar synchronization check
249279 11/16/08 22:46 Chiranjib Effect of correction coil
249276 11/16/08 22:43 D. Parno L-HRS S2m ADC 4 is back
249275 11/16/08 22:28 Chiranjib CORRECTION COIL TURNED OFF
249274 11/16/08 22:18 xqian Dead time issue again,
249273 11/16/08 22:12 D. Parno Data from S2m Left ADC 4
249272 11/16/08 22:08 narayan target test performed
249271 11/16/08 22:06 adaq target: Spin Flip Report: Correction Coil Off before Flip#28
249268 11/16/08 21:29 adaq Q1 Helium flow alarm
249265 11/16/08 20:57 adev BigBite WireChamber Measured Voltage change!
249263 11/16/08 20:44 adaq BCM/BPM strip scan
249261 11/16/08 20:42 adaq Hall A tools scan
249260 11/16/08 20:09 adaq Target Field and Oven Temps snapshot
249259 11/16/08 20:09 adaq Target Cell Temps snapshot
249258 11/16/08 20:09 adaq Radiation and Ozone snapshot
249257 11/16/08 20:09 adaq Hall A Tools snapshot
249256 11/16/08 20:09 adaq BigBox Magnet snapshot
249255 11/16/08 20:09 adaq Beam Position and Current snapshot
249254 11/16/08 20:09 adaq Beam Energy in A and C snapshot
249249 11/16/08 19:30 adaq target: Spin Flip Report: Regular Update: We initiated taking data a little, Regular update
249247 11/16/08 18:18 Kalyan BB Shower vs preshower for T1 and T6
249245 11/16/08 17:03 Kalyan Run 3566 is production
249244 11/16/08 16:51 Jin Golden Run -> 3565
249241 11/16/08 16:22 Chiranjib Polarization just before the production @12uA
249240 11/16/08 16:20 Chiranjib Target moved to He-3/ started spin flip
249239 11/16/08 16:18 Chiranjib BB wire ch dark current
249238 11/16/08 16:18 Chiranjib BB wire chamber dark currents
249237 11/16/08 16:16 Kalyan Trigger rates with 12uA current on 3He
249235 11/16/08 16:03 dhamilton Day Shift Summary
249233 11/16/08 15:52 Zhihong Ye Laser Spetrum Screen for Nov 16th Day shift
249232 11/16/08 15:20 Zhihong Ye Target moved to Optics Target Position
249231 11/16/08 14:41 Yi Qiang Target Strip Charts
249230 11/16/08 14:32 adev BigBite WireChamber Measured Voltage change!
249228 11/16/08 14:16 Kalyan BB scintillator right chan 10 TDC is missing
249227 11/16/08 14:04 Kalyan Run 3565 is production and not test
249226 11/16/08 13:54 Yi Qiang Target Spin Flip Code Restarted
249225 11/16/08 13:40 Kalyan Test for LHRS retiming
249224 11/16/08 13:34 D. Parno Paddle Timing
249223 11/16/08 13:34 adaq target: Spin Flip Report: Target Auto Spin Flip is back on line
249219 11/16/08 13:14 Kalyan Plot of HRS target z - BB target z for good e-pi coincidences
249217 11/16/08 12:51 D. Parno Timing Shifts -- Individual Paddles
249216 11/16/08 12:50 Kalyan Coincidence time for run 3562
249215 11/16/08 12:38 Kalyan S2m Left ADC 4 is missing
249212 11/16/08 12:13 D. Parno L-HRS TDC still shifted after trigger download
249211 11/16/08 12:09 adaq Target Field and Oven Temps snapshot
249210 11/16/08 12:09 adaq Target Cell Temps snapshot
249209 11/16/08 12:09 adaq Radiation and Ozone snapshot
249208 11/16/08 12:09 adaq Hall A Tools snapshot
249207 11/16/08 12:09 adaq BigBox Magnet snapshot
249206 11/16/08 12:09 adaq Beam Position and Current snapshot
249205 11/16/08 12:09 adaq Beam Energy in A and C snapshot
249203 11/16/08 11:56 A. Camsonne Redownload of the HRS trigger
249202 11/16/08 11:54 atrig Trigger-Download
249198 11/16/08 11:08 xqian Here is the plot
249197 11/16/08 11:03 Kalyan T3 delay was set to zero on 11th nov (log 248234 )
249194 11/16/08 10:24 D. Parno Shift in L-HRS TDC Spectrum
249193 11/16/08 10:22 xqian RUN 3559 and the run before that all habe timing problems in left HRS!!!
249192 11/16/08 10:21 xqian LEFT HRS out of window again!!!
249188 11/16/08 09:55 xqian live time again for RUN 3556
249186 11/16/08 09:52 xqian Update the drift time t0 for the new run , I created a new database of 20081113
249184 11/16/08 09:44 xqian Update in background rates
249183 11/16/08 09:21 dhamilton BPM/BCM strip charts
249182 11/16/08 09:20 dhamilton Hall A Tools
249179 11/16/08 08:28 xqian RUN 3553,3554,3555 need special treatment
249172 11/16/08 08:06 Simonetti Laser spectrum and Spin flip system control
249170 11/16/08 08:04 cisbani Owl Shift Summary
249166 11/16/08 07:56 xqian live time for RUN 3550
249165 11/16/08 07:54 adaq target: Spin Flip Report: Laser spectrum
249164 11/16/08 07:48 Simonetti Target activity
249162 11/16/08 07:22 xqian Effect of our new collimator
249161 11/16/08 07:17 Jin Huang Scalar Fast Clock Synchronization Check
249160 11/16/08 07:09 xqian Prescale study
249159 11/16/08 07:04 Jin Huang New Replay, New Online Plots
249157 11/16/08 06:53 cisbani On RICH deadtime during cosmics in Nov/15 swing shift
249147 11/16/08 06:23 xqian Background study after more shielding and collimator
249145 11/16/08 06:06 xqian BigBite Retming works, Great!!!
249144 11/16/08 06:02 xqian Live time for 2uA 3He, see plot
249142 11/16/08 05:31 cisbani spot++ 3He TPoc
249141 11/16/08 05:14 Jin Huang Real Spin Flip Start. Pol.=58%
249139 11/16/08 05:10 Simonetti Temperatures of He3 target during polarization
249138 11/16/08 04:58 Simonetti moving target to He3 cell
249136 11/16/08 04:44 xqian Live time in RUN3541
249135 11/16/08 04:32 xqian RUN 3541,3542,3543, have a lot of F1 TDC overflow. Why?
249133 11/16/08 04:23 cisbani BB chamber tripped again, run 3542
249131 11/16/08 04:12 cisbani BB Chamber HV Trip, run 3542
249130 11/16/08 04:10 adaq spot++ on BeO 3x4 MCC
249129 11/16/08 04:09 adaq Target Field and Oven Temps snapshot
249128 11/16/08 04:09 adaq Target Cell Temps snapshot
249127 11/16/08 04:08 adaq Radiation and Ozone snapshot
249126 11/16/08 04:08 adaq Hall A Tools snapshot
249125 11/16/08 04:08 adaq BigBox Magnet snapshot
249124 11/16/08 04:08 adaq Beam Position and Current snapshot
249123 11/16/08 04:08 adaq Beam Energy in A and C snapshot
249122 11/16/08 03:59 Simonetti Hall A Target Cell Temperature
249119 11/16/08 03:50 xqian Some thoughts about the dead time with our current trigger setup
249118 11/16/08 03:44 adaq target: Spin Flip Report: No Shift Worker Comment
249117 11/16/08 03:32 Simonetti Beam tuning and run with empty reference cell
249103 11/16/08 02:32 ycwang RICH HV tripped when tuning up beam
249093 11/16/08 01:54 adaq Harp scan problem solved, beam profile within 200 um
249091 11/16/08 01:10 A. Camsonne BCM strip chart
249090 11/16/08 00:42 cisbani HallA Tools
249089 11/16/08 00:24 Kalyan BB thresholds set to beam conditions
249086 11/16/08 00:10 adaq target: Spin Flip Report: test
Saturday
249083 11/15/08 23:49 R. Michaels notes after raster repair (TECH)
249069 11/15/08 23:17 adev SCALER = -1
249062 11/15/08 23:08 adev SCALER = -1
249057 11/15/08 22:59 adev SCALER = -1
249041 11/15/08 22:18 A. Camsonne Update about raster
249027 11/15/08 21:33 A. Camsonne Horizontal raster need to be replaced
249026 11/15/08 21:04 R. Michaels re: raster repair
249024 11/15/08 20:25 D. Dutta Raster problem
249023 11/15/08 20:08 adaq Target Field and Oven Temps snapshot
249022 11/15/08 20:08 adaq Target Cell Temps snapshot
249021 11/15/08 20:08 adaq Radiation and Ozone snapshot
249020 11/15/08 20:08 adaq Hall A Tools snapshot
249019 11/15/08 20:08 adaq BigBox Magnet snapshot
249018 11/15/08 20:08 adaq Beam Position and Current snapshot
249017 11/15/08 20:08 adaq Beam Energy in A and C snapshot
249016 11/15/08 20:01 Dutta Beam is back
249015 11/15/08 19:37 Jin Fake Spin Signal ON till Production
249014 11/15/08 19:06 adaq target: Spin Flip Report: Target Moved to Empty position. Auto Flip was already disabled, I fliped the target once, before moving it.
249013 11/15/08 19:02 Chiranjib EPR correction
249012 11/15/08 17:22 Chiranjib LHRS TDC spectra with T3 and T5
249009 11/15/08 16:55 A. Camsonne Left Q1 tripped while ramping down right arm
249007 11/15/08 16:04 Lindgren Summary
249006 11/15/08 15:03 Chiranjib EPR Calibration/ SPIN UP for TRANSVERSE + / POLARIZATION etc.
249005 11/15/08 13:42 ycwang A1 should be OK for now
248998 11/15/08 12:48 Lindgren Hall A Tools
248995 11/15/08 12:39 Lindgren Hall A Entry
248993 11/15/08 12:19 D. Parno Scaler problem fixed (new Compton photon DAQ)
248991 11/15/08 12:11 Lindgren Magnets
248990 11/15/08 12:08 adaq Target Field and Oven Temps snapshot
248989 11/15/08 12:08 adaq Target Cell Temps snapshot
248988 11/15/08 12:08 adaq Radiation and Ozone snapshot
248987 11/15/08 12:08 adaq Hall A Tools snapshot
248986 11/15/08 12:08 adaq BigBox Magnet snapshot
248985 11/15/08 12:08 adaq Beam Position and Current snapshot
248984 11/15/08 12:07 adaq Beam Energy in A and C snapshot
248976 11/15/08 11:29 Lindgren Hall A exit
248975 11/15/08 11:14 A. Camsonne Compton IOC on hareboot4 port 3
248974 11/15/08 11:01 A. Camsonne APC remote controlled plug
248972 11/15/08 11:00 A. Camsonne Compton webcams
248969 11/15/08 10:53 adev BigBite WireChamber Measured Voltage change!
248963 11/15/08 10:29 xqian I think that we should be able to run 12uA
248946 11/15/08 10:07 Lindgren Hall A Entry
248944 11/15/08 09:40 Lindgren Day shift begins
248941 11/15/08 08:54 A. Camsonne Remote controlled plug for compton electron detector
248940 11/15/08 08:52 A. Camsonne experiment status
248938 11/15/08 04:07 adaq Target Field and Oven Temps snapshot
248937 11/15/08 04:07 adaq Target Cell Temps snapshot
248936 11/15/08 04:07 adaq Radiation and Ozone snapshot
248935 11/15/08 04:07 adaq Hall A Tools snapshot
248934 11/15/08 04:07 adaq BigBox Magnet snapshot
248933 11/15/08 04:07 adaq Beam Position and Current snapshot
248932 11/15/08 04:07 adaq Beam Energy in A and C snapshot
248868 11/15/08 01:31 brads BB Cerenkov gain check results using LED (w/ snapshots)
248855 11/15/08 01:20 brads BB Cerenkov gain check results using LED
248829 11/15/08 00:39 brads Cerenkov work finished, DAQ now being used by RICH guys
Friday
248806 11/14/08 22:02 brads Run 3411 stopped for Cerenkov check out
248799 11/14/08 21:16 Jin Spin Up Test @ Spin + State
248798 11/14/08 20:56 brads Upper BigBite HV crate is beeping
248795 11/14/08 20:07 adaq Target Field and Oven Temps snapshot
248794 11/14/08 20:07 adaq Target Cell Temps snapshot
248793 11/14/08 20:07 adaq Radiation and Ozone snapshot
248792 11/14/08 20:07 adaq Hall A Tools snapshot
248791 11/14/08 20:07 adaq BigBox Magnet snapshot
248790 11/14/08 20:07 adaq Beam Position and Current snapshot
248789 11/14/08 20:07 adaq Beam Energy in A and C snapshot
248788 11/14/08 19:46 Chiranjib/Ji QUICK optics enclosure inspection.
248787 11/14/08 19:34 Yi Qiang AFP Loss Test
248786 11/14/08 18:57 V. Sulkosky Re: ADC Spectra with T3 and T5
248785 11/14/08 18:32 brads BigBite HV GUI now in VNC wrapper
248784 11/14/08 18:26 adev BigBite WireChamber Measured Voltage change!
248783 11/14/08 18:15 adev BigBite WireChamber Measured Voltage change!
248782 11/14/08 18:15 D. Parno BCM output wiring change
248781 11/14/08 17:53 D. Parno Compton VME crate moved to remote power control
248780 11/14/08 17:44 Chiranjib Temperature Test on ASTRAL
248779 11/14/08 17:22 Chiranjib/ Y EPR measurement
248778 11/14/08 17:14 brads BB Cerenkov bad signal cable on PMT 20 fixed
248764 11/14/08 15:23 A. Camsonne Webcam
248760 11/14/08 14:48 Kalyan Run 3402 ROC8 pedestals
248753 11/14/08 14:17 A. Camsonne Compton electron detector interlock checked out
248752 11/14/08 14:13 A. Camsonne Left Q1 in trouble
248751 11/14/08 14:07 Higinbotham Right HRS Q2 and Dipole Test
248748 11/14/08 13:03 R. Michaels HAPPEX DAQ to "safe" mode
248747 11/14/08 12:07 adaq Target Field and Oven Temps snapshot
248746 11/14/08 12:07 adaq Target Cell Temps snapshot
248745 11/14/08 12:06 adaq Radiation and Ozone snapshot
248744 11/14/08 12:06 adaq Hall A Tools snapshot
248743 11/14/08 12:06 adaq BigBox Magnet snapshot
248742 11/14/08 12:06 adaq Beam Position and Current snapshot
248741 11/14/08 12:06 adaq Beam Energy in A and C snapshot
248740 11/14/08 10:13 V. Sulkosky Coefficients for Left Gas Cerenkov Updated
248737 11/14/08 09:26 A. Camsonne HRS ramped down to 0 Amps
248736 11/14/08 09:20 A. Camsonne proxy for accelerator website
248735 11/14/08 08:57 ycwang The previous halog about A1 rates are without beam.
248734 11/14/08 08:52 ycwang A1 rate with 8uA beam
248733 11/14/08 08:30 A. Camsonne CHL still down
248732 11/14/08 07:57 Yi Zhang B Cherenkov ADC in low HV, 5 pass beam
248731 11/14/08 07:34 adev BigBite WireChamber Measured Voltage change!
248730 11/14/08 07:32 adev BigBite WireChamber Measured Voltage change!
248729 11/14/08 07:23 adev BigBite WireChamber Measured Current change!
248726 11/14/08 07:13 adev BigBite WireChamber Measured Current change!
248725 11/14/08 06:14 Yi Zhang BB Cherenkov ADC in low current, 5 pass beam
248722 11/14/08 04:06 adaq Target Field and Oven Temps snapshot
248721 11/14/08 04:06 adaq Target Cell Temps snapshot
248720 11/14/08 04:06 adaq Radiation and Ozone snapshot
248719 11/14/08 04:06 adaq Hall A Tools snapshot
248718 11/14/08 04:06 adaq BigBox Magnet snapshot
248717 11/14/08 04:06 adaq Beam Position and Current snapshot
248716 11/14/08 04:06 adaq Beam Energy in A and C snapshot
248715 11/14/08 04:02 Jin Target Activity of Nov-14 Owl
248712 11/14/08 01:24 adev BigBite WireChamber Measured Voltage change!
248709 11/14/08 01:14 adev BigBite WireChamber Measured Voltage change!
248708 11/14/08 00:11 adev BigBite WireChamber Measured Voltage change!
248707 11/14/08 00:11 brads BB Cerenkov work and todo list
248706 11/14/08 00:06 dhamilton Status Update - no owl shift
248705 11/14/08 00:01 Kalyan Change BB threshold to beam conditions
Thursday
248704 11/13/08 23:49 adev BigBite GUI Stopped!
248703 11/13/08 23:48 Jin Roc15 reboot, start run # 3393
248702 11/13/08 23:46 adev BigBite GUI Stopped!
248700 11/13/08 23:20 Jin Bigbite -> 710A
248698 11/13/08 22:25 xqian I may interprete wrong
248697 11/13/08 21:44 brads/kalyan BigBite retiming, potential gotchas with T6/T2
248696 11/13/08 21:43 xqian Pion+ with the gap on the top of our acceptance
248692 11/13/08 20:06 adaq Target Field and Oven Temps snapshot
248691 11/13/08 20:06 adaq Target Cell Temps snapshot
248690 11/13/08 20:06 adaq Radiation and Ozone snapshot
248689 11/13/08 20:06 adaq Hall A Tools snapshot
248688 11/13/08 20:06 adaq BigBox Magnet snapshot
248687 11/13/08 20:06 adaq Beam Position and Current snapshot
248686 11/13/08 20:05 adaq Beam Energy in A and C snapshot
248681 11/13/08 19:34 camsonne / c RICH activity during Nov/13/2008 access
248676 11/13/08 18:49 Chiranjib/yi Fiber -comet combo
248672 11/13/08 18:46 adev BigBite WireChamber Measured Voltage change!
248671 11/13/08 18:46 adev BigBite WireChamber Measured Voltage change!
248670 11/13/08 18:39 Yi/Jin/Chira Comet Laser Spectrum in Vertical Pumping
248669 11/13/08 18:35 adev BigBite WireChamber Measured Voltage change!
248662 11/13/08 17:46 ycwang A1 channel 03 come back
248661 11/13/08 17:43 V. Sulkosky Strobe was Missing in Scalers
248660 11/13/08 17:42 Yi Qiang Lasers are now UNLOCKED
248659 11/13/08 17:35 V. Sulkosky Re: Changes in Left HRS Trigger Signals in DecData Class
248657 11/13/08 17:17 Jack Left, not Right, Q1
248656 11/13/08 17:10 LeRose Wien Angle Change
248653 11/13/08 17:03 Jack (Techs) Right spectrometer status
248651 11/13/08 16:55 Jack (Techs) Right Q1 status
248634 11/13/08 16:32 Techs Check list
248601 11/13/08 15:42 V. Sulkosky Re: RF Time Signal
248566 11/13/08 14:10 A. Camsonne Beam dowm for CHL work expected return of beam around 8PM today
248561 11/13/08 13:21 Kalyan HV on two shower channels back to normal
248554 11/13/08 12:05 adaq Target Field and Oven Temps snapshot
248553 11/13/08 12:05 adaq Target Cell Temps snapshot
248552 11/13/08 12:05 adaq Radiation and Ozone snapshot
248551 11/13/08 12:05 adaq Hall A Tools snapshot
248550 11/13/08 12:05 adaq BigBox Magnet snapshot
248549 11/13/08 12:05 adaq Beam Position and Current snapshot
248548 11/13/08 12:05 adaq Beam Energy in A and C snapshot
248547 11/13/08 12:03 Yi Qiang BigBite Magnets is OFF
248545 11/13/08 12:00 Yi Qiang Lasers are locked for putting in new collimator
248543 11/13/08 11:57 Yi Qiang EPR Calibration
248540 11/13/08 10:58 adaq target: Spin Flip Report: Target Report before EPR measurement
248539 11/13/08 10:19 ycwang The base has problem for previous No.3 in A1
248533 11/13/08 09:18 Kalyan BB Shower display for background
248530 11/13/08 08:15 Z.-E. Mezian Shift Summary
248527 11/13/08 08:00 Yi Zhang Auto spin flip off
248519 11/13/08 06:20 Yi Zhang BB Cherenkov HV set
248517 11/13/08 06:15 Yi Zhang trig rates while beam on (8uA)
248511 11/13/08 05:34 Chiranjib Target Screens II
248510 11/13/08 05:32 Chiranjib Target Screens I
248507 11/13/08 04:54 Yi Zhang BB Cherenkov HV in channel L12.4
248504 11/13/08 04:45 Yi Zhang BB Cherenkov signal
248503 11/13/08 04:44 xqian Update the BigBite library
248502 11/13/08 04:20 Yi Zhang BB Cerenkov HV
248501 11/13/08 04:17 xqian plots
248499 11/13/08 04:13 xqian BigBite Wire Chamber Background study
248497 11/13/08 04:05 adaq Target Field and Oven Temps snapshot
248496 11/13/08 04:05 adaq Target Cell Temps snapshot
248495 11/13/08 04:05 adaq Radiation and Ozone snapshot
248494 11/13/08 04:05 adaq Hall A Tools snapshot
248493 11/13/08 04:05 adaq BigBox Magnet snapshot
248492 11/13/08 04:05 adaq Beam Position and Current snapshot
248491 11/13/08 04:05 adaq Beam Energy in A and C snapshot
248485 11/13/08 02:17 Yi Zhang BB Cerenkov HV tuning
248482 11/13/08 01:10 Chiranjib Target RTDs
248481 11/13/08 01:08 Chiranjib Spin flip screens and Laser spectrum
248478 11/13/08 00:38 Chiranjib S2 TDCs
248477 11/13/08 00:26 Chiranjb S1 TDCs
248474 11/13/08 00:04 cisbani Swing shift summary
Wednesday
248472 11/12/08 23:50 Chiranjb S1 ADCs look okay
248471 11/12/08 23:41 Yi Zhang golden run sets to 3306
248469 11/12/08 23:08 adaq target: Spin Flip Report: spin flip status
248467 11/12/08 22:56 nelyubin laser_line_spectrum
248464 11/12/08 22:23 Chiranjib RUN 3304 is a check for LHRS
248462 11/12/08 22:21 adev BigBite WireChamber Measured Voltage change!
248458 11/12/08 21:52 cisbani HallA Tools
248451 11/12/08 21:26 cisbani spot raster 6.6x6.6 mcc
248450 11/12/08 21:23 cisbani trigger rates, raster 6.6x6.6 mcc
248446 11/12/08 21:16 cisbani trigger rates raster 6x6 mcc
248444 11/12/08 21:14 cisbani spot, raster 5x5
248442 11/12/08 21:11 cisbani trigger rates, raster 5x5 mcc
248440 11/12/08 21:08 cisbani spot, raster 4x4 mcc
248438 11/12/08 21:04 cisbani trigger rates, raster 4x4 mcc
248436 11/12/08 21:01 cisbani spot, raster 3x3 mcc
248434 11/12/08 20:56 cisbani trigger rates, raster 3x3 mcc
248433 11/12/08 20:55 brads BB Cerenkov HVs turned on.
248431 11/12/08 20:54 cisbani spot, raster 2x2 mcc
248430 11/12/08 20:52 cisbani trigger rates raster 2x2 mcc
248428 11/12/08 20:46 adev BigBite WireChamber Measured Voltage change!
248427 11/12/08 20:41 E.Chudakov Moller results
248426 11/12/08 20:05 adaq Target Field and Oven Temps snapshot
248425 11/12/08 20:05 adaq Target Cell Temps snapshot
248424 11/12/08 20:04 adaq Radiation and Ozone snapshot
248423 11/12/08 20:04 adaq Hall A Tools snapshot
248422 11/12/08 20:04 adaq BigBox Magnet snapshot
248421 11/12/08 20:04 adaq Beam Position and Current snapshot
248420 11/12/08 20:04 adaq Beam Energy in A and C snapshot
248419 11/12/08 18:18 V. Sulkosky Left HRS Momentum 2.35 GeV/c
248418 11/12/08 17:13 Moffit HAPPEX DAQ Working... Bring it back up.
248417 11/12/08 16:31 Moffit Re: HAPPEX DAQ ROC23
248416 11/12/08 16:31 V. Sulkosky Left Q1 Tripped off
248414 11/12/08 16:04 averett Day Shift Summary
248411 11/12/08 15:57 bradshaw Spectra at Optics
248410 11/12/08 15:55 nelyubin absorption spectrum
248409 11/12/08 15:47 Sirish Compton Commissioning
248408 11/12/08 15:26 A. Camsonne HAPPEX DAQ down
248407 11/12/08 12:43 adev BigBite WireChamber Measured Voltage change!
248406 11/12/08 12:41 Kalyan wirechamber HV trip
248405 11/12/08 12:37 bradshaw Move Target to Optics
248404 11/12/08 12:33 adev BigBite WireChamber Measured Voltage change!
248403 11/12/08 12:04 adaq Target Field and Oven Temps snapshot
248402 11/12/08 12:04 adaq Target Cell Temps snapshot
248401 11/12/08 12:04 adaq Radiation and Ozone snapshot
248400 11/12/08 12:04 adaq Hall A Tools snapshot
248399 11/12/08 12:04 adaq BigBox Magnet snapshot
248398 11/12/08 12:04 adaq Beam Position and Current snapshot
248397 11/12/08 12:04 adaq Beam Energy in A and C snapshot
248396 11/12/08 11:04 averett target to empty
248395 11/12/08 11:04 adaq target: Spin Flip Report: Day Shift Target Report
248393 11/12/08 10:48 Yi Qiang Target Laser IOC is back online
248392 11/12/08 10:41 Kalyan coincidence time run 3291
248391 11/12/08 10:37 A. Camsonne RICH configuration in
248390 11/12/08 10:27 ycwang A1 rates with beam 8uA, previous one beam Trip
248387 11/12/08 10:15 ycwang Channel 03 has less events than the neighbors
248386 11/12/08 10:10 A. Camsonne HAPPEX DAQ crashing
248385 11/12/08 10:09 A. Camsonne HAPPEX DAQ crashing
248384 11/12/08 09:59 V. Sulkosky New Trigger Channels Working
248383 11/12/08 09:58 camsonne / c RICH activity during Nov/11/2008 Tue access
248378 11/12/08 08:46 ycwang A1 rate with beam 8uA
248377 11/12/08 08:32 Jin Target Temperature screen shot
248376 11/12/08 08:30 Jin Huang Target Activity Before Production Running
248374 11/12/08 08:19 Z.E. Meziani Shift Summary
248371 11/12/08 08:14 Lamiaa Spot++ for run 3290 w/. 5x6 mm (MCC unit)
248369 11/12/08 08:10 Lamiaa Spot++ for run 3289 w/. 4x5 mm raster (MCC unit)
248366 11/12/08 08:06 adaq target: Spin Flip Report: Spin FLip Subsystem Switched On and Take Over The Target.
248364 11/12/08 08:04 Lamiaa spot++ for run 3288 w/. 3x4 mm raster (MCC unit)
248361 11/12/08 07:58 Lamiaa Spot++ result for run 3287 w/. 2.5x3 mm raster (MCC unit)
248359 11/12/08 07:52 Yi Zhang maybe a timing problem?
248355 11/12/08 06:44 Yi Zhang Pion Rejector HV
248349 11/12/08 06:12 Jin Vac Ref Cell, Fill with H2 -> 100psig
248345 11/12/08 05:18 Jin Huang Beam Tuning for Reference cell run
248343 11/12/08 05:02 adev BigBite WireChamber Measured Voltage change!
248340 11/12/08 04:52 Yi Zhang BPM & raster size after 5 pass beam tuning.
248339 11/12/08 04:50 Jin Ref Cell -> 100 psig N2
248337 11/12/08 04:50 adaq target: Spin Flip Report: Fake Spin Signal ON, 20min/flip. True Target Pol. > 40%
248335 11/12/08 04:40 Jin Target -> Ref Cell
248334 11/12/08 04:40 Jin Target -> BeO
248330 11/12/08 04:25 Jin Target -> Carbon with Hole -> BeO
248322 11/12/08 04:04 adaq Target Field and Oven Temps snapshot
248321 11/12/08 04:04 adaq Target Cell Temps snapshot
248320 11/12/08 04:04 adaq Radiation and Ozone snapshot
248319 11/12/08 04:04 adaq Hall A Tools snapshot
248318 11/12/08 04:04 adaq BigBox Magnet snapshot
248317 11/12/08 04:03 adaq Beam Position and Current snapshot
248316 11/12/08 04:03 adaq Beam Energy in A and C snapshot
248312 11/12/08 03:53 Lamiaa Trigger screen shot. Empth ref, cell with raster 2.5x3 mm & BPMBX -1 mm
248305 11/12/08 03:28 adev BigBite WireChamber Measured Voltage change!
248304 11/12/08 03:28 Jin Trigger ScreenShot, Empty Ref, 3x3 MCC raster, 2uA
248301 11/12/08 03:24 Lamiaa Spot++ for run 3267
248299 11/12/08 03:20 Jin Trigger Screen Shot, Empty Ref Cell, 2.5x3 MCC Raster, 2uA
248297 11/12/08 03:17 adev BigBite WireChamber Measured Voltage change!
248296 11/12/08 03:12 Lamiaa Spot++ with Holy Carbon target
248294 11/12/08 03:10 xqian How does retiming work, I will explain this with the left HRS data.
248293 11/12/08 03:09 Jin For last screen shot, we asked for 3x3 to MCC
248292 11/12/08 03:06 Jin Trigger Screen Shot @ ref cell 2x2 2uA
248290 11/12/08 03:02 Jin Target -> Ref Cell
248289 11/12/08 03:02 xqian Trigger timing problem is still there. Solved half.
248287 11/12/08 02:36 Yi Zhang spot++ check
248286 11/12/08 02:35 Jin Target -> Carbon with Hole
248281 11/12/08 01:32 adev BigBite WireChamber Measured Voltage change!
248280 11/12/08 01:22 adev BigBite WireChamber Measured Voltage change!
248279 11/12/08 01:12 Jin Target -> BeO
248273 11/12/08 00:09 cisbani Swing Shift Summary
Tuesday
248259 11/11/08 23:08 V. Sulkosky Run 3253 Pedestal Run
248258 11/11/08 23:05 Yi Qiang Laser IOC is offline
248257 11/11/08 23:03 Jin/Chiranji Comet Laser Spectrum/Spin up start at ~23%
248256 11/11/08 22:52 Yi Qiang 2 COMET Lasers are ON for transverse pumping
248255 11/11/08 22:52 Yi Qiang 2 COMET Lasers are ON for transverse pumping
248251 11/11/08 22:34 V. Sulkosky New Trigger Signals in 1877 Empty
248250 11/11/08 22:31 V. Sulkosky Quick Access to fix RF Time Signals
248249 11/11/08 22:23 brads Plugged BB Cerenkov PMT 5 into networked scope
248239 11/11/08 21:53 adaq HallA Tools
248238 11/11/08 21:50 brads re: RHRS VDC discriminator problem
248237 11/11/08 21:07 V. Sulkosky Re: Checked RT on Left HRS
248236 11/11/08 21:06 Jin Target Holding field messed up by wired Bigbite setting, Pol -> 10%, target -> empty, test pumping up with new laser
248235 11/11/08 20:50 V. Sulkosky Added T1 and T3 to 1875
248234 11/11/08 20:47 Kalyan DAQ activity today
248233 11/11/08 20:41 V. Sulkosky Added TDC to ROC 4 in Slot 9 for Trigger Signals
248230 11/11/08 20:03 adaq Target Field and Oven Temps snapshot
248229 11/11/08 20:03 adaq Target Cell Temps snapshot
248228 11/11/08 20:03 adaq Radiation and Ozone snapshot
248227 11/11/08 20:03 adaq Hall A Tools snapshot
248226 11/11/08 20:03 adaq BigBox Magnet snapshot
248225 11/11/08 20:03 adaq Beam Position and Current snapshot
248224 11/11/08 20:03 adaq Beam Energy in A and C snapshot
248222 11/11/08 20:00 Jin Target -> He3 Cell
248221 11/11/08 19:56 adaq target: Spin Flip Report: Software Update Done
248220 11/11/08 19:39 A. Camsonne Webcam Compton
248214 11/11/08 19:15 brads BB Cerenkov PMTs 5, 6,7 shielded from beamline
248212 11/11/08 19:12 brads Two paddles set-up downstream of BB Shower detector
248209 11/11/08 18:45 Yi Qiang Lasers are now UNLOCKED
248202 11/11/08 18:28 ycwang The rate of A1 for different channnel
248201 11/11/08 18:21 ycwang A1 channel 03 has signal after we change base
248188 11/11/08 16:01 Saha Shift Summary
248187 11/11/08 15:44 Techs Checklist
248166 11/11/08 14:26 Jin Target Computer is rebooted
248151 11/11/08 13:07 ycwang Need a new base for A1
248148 11/11/08 12:49 ycwang After swap base 3 and base 4
248141 11/11/08 12:30 adaq target: Spin Flip Report: ScreenShot before quit
248140 11/11/08 12:03 adaq Target Field and Oven Temps snapshot
248139 11/11/08 12:03 adaq Target Cell Temps snapshot
248138 11/11/08 12:03 adaq Radiation and Ozone snapshot
248137 11/11/08 12:03 adaq Hall A Tools snapshot
248136 11/11/08 12:03 adaq BigBox Magnet snapshot
248135 11/11/08 12:03 adaq Beam Position and Current snapshot
248134 11/11/08 12:03 adaq Beam Energy in A and C snapshot
248122 11/11/08 10:25 xqian BigBite Shower Threshold to cosmics
248119 11/11/08 10:21 Yi Qiang Field rotated to vertical for correction coil check
248118 11/11/08 10:18 Yi Qiang EPR Correction
248117 11/11/08 10:11 Yi Qiang EPR Correction
248113 11/11/08 09:41 Chiranjb/Yi EPR measurement
248112 11/11/08 09:39 Yi Qiang Laser locked
248110 11/11/08 09:29 adev BigBite WireChamber Measured Voltage change!
248109 11/11/08 09:21 A. Camsonne Powering BigBite Magnet down
248108 11/11/08 09:01 Lingyan Zhu Laser spectrum
248106 11/11/08 08:45 R. Michaels re: coinc. timing
248105 11/11/08 08:40 xqian Just want to clarify
248104 11/11/08 08:34 R. Michaels re: trigger timing issues
248101 11/11/08 08:31 A. Camsonne Target work hazard analysis and plan
248100 11/11/08 08:20 Kalyan/Xin Trying to understand coincidence timing spectra
248099 11/11/08 08:18 xqian BigBite Wire Chamber T0 offset,
248096 11/11/08 08:02 Yi Zhang field direction
248095 11/11/08 08:02 Meziani Shift Summary
248093 11/11/08 07:56 Yi Zhang RTD & laser screen
248092 11/11/08 07:52 Yi Zhang Laser spectrum
248091 11/11/08 07:52 xqian BigBite Gas Cerenkov ADC signal
248090 11/11/08 07:52 Yi Zhang Laser spectrum
248089 11/11/08 07:50 adaq target: Spin Flip Report: owl shift on Nov.11
248083 11/11/08 06:32 Lamiaa Update: ROC synchronization distributions for run 3197.
248081 11/11/08 06:09 meziani Alarm handler acknowledment window not opening
248078 11/11/08 05:49 Lamiaa Empty ROC synchronization distributions for run 3196
248075 11/11/08 05:18 Moffit HAPPEX GUI: changed database
248070 11/11/08 04:03 adaq Target Field and Oven Temps snapshot
248069 11/11/08 04:03 adaq Target Cell Temps snapshot
248068 11/11/08 04:02 adaq Radiation and Ozone snapshot
248067 11/11/08 04:02 adaq Hall A Tools snapshot
248066 11/11/08 04:02 adaq BigBox Magnet snapshot
248065 11/11/08 04:02 adaq Beam Position and Current snapshot
248064 11/11/08 04:02 adaq Beam Energy in A and C snapshot
248061 11/11/08 03:24 xqian Possible timing problem; some suggestions
248051 11/11/08 01:13 Chiranjib Vertex Z II
248048 11/11/08 00:29 Jin Huang Update of Online Analysis Scripts. Golden run -> Run # 3171. More Events to replay
248039 11/11/08 00:04 E. Cisbani Swing Shift summary
248037 11/11/08 00:03 Kalyan Target z plot
Monday
248036 11/10/08 23:49 Chiranjib EPICS variable for the target ladder IN
248030 11/10/08 22:35 cisbani spot++ after long beam off
248027 11/10/08 22:29 Jin Target -> He3, Spin Flip recovers
248025 11/10/08 21:51 Kalyan/Xin Trying to understand coincidence timing spectra
248024 11/10/08 20:37 cisbani hallA tools
248023 11/10/08 20:02 adaq Target Field and Oven Temps snapshot
248022 11/10/08 20:02 adaq Target Cell Temps snapshot
248021 11/10/08 20:02 adaq Radiation and Ozone snapshot
248020 11/10/08 20:02 adaq Hall A Tools snapshot
248019 11/10/08 20:02 adaq BigBox Magnet snapshot
248018 11/10/08 20:02 adaq Beam Position and Current snapshot
248017 11/10/08 20:02 adaq Beam Energy in A and C snapshot
248016 11/10/08 19:47 Jin Target -> Optics
248015 11/10/08 19:46 Sirish Compton Status
248014 11/10/08 18:25 ycwang A1 all sum in 1 pass run # 2512 and 2 pass run #2777
248013 11/10/08 18:22 ycwang A1 all sum in 1 pass run # 2512 and 2 pass run #2777
248012 11/10/08 17:50 ycwang A1 rates Channel 10
248011 11/10/08 17:48 ycwang A1 rates when Channel 4 HV turn off and on
247999 11/10/08 17:19 ycwang A1 rates in cosmic state
247997 11/10/08 17:11 Yi Qiang Auto spin flip paused for beam down
247996 11/10/08 17:00 Jin MWDC hit drop highly depends on events range replayed
247993 11/10/08 16:00 Saha Day Shift Summary
247990 11/10/08 15:49 narayan Screen Shot Pumping Laser Absorption Spectrum
247987 11/10/08 15:18 Kalyan Re: LHRS aero A1 CH#3 dead
247986 11/10/08 15:08 Kalyan Al left HV screen
247985 11/10/08 14:47 yuan,Dimitri LHRS aero A1 CH#3 dead
247980 11/10/08 13:31 yuan SPOT++ for Run 3167
247977 11/10/08 13:16 yuan BPM and BCM charts
247976 11/10/08 13:13 Saha Hall A Tools Screen
247973 11/10/08 12:39 Kalyan shower channels
247971 11/10/08 12:23 narayan target moved back
247970 11/10/08 12:14 narayan target moved
247969 11/10/08 12:10 narayan narayan
247968 11/10/08 12:02 adaq Target Field and Oven Temps snapshot
247967 11/10/08 12:02 adaq Target Cell Temps snapshot
247966 11/10/08 12:02 adaq Radiation and Ozone snapshot
247965 11/10/08 12:02 adaq Hall A Tools snapshot
247964 11/10/08 12:02 adaq BigBox Magnet snapshot
247963 11/10/08 12:02 adaq Beam Position and Current snapshot
247962 11/10/08 12:01 adaq Beam Energy in A and C snapshot
247960 11/10/08 11:42 adaq target: Spin Flip Report: General Shift Update
247957 11/10/08 11:04 Kalyan increased HV on three shower blocks
247952 11/10/08 10:13 Kalyan BB wirechamber signals improved and back to normal again
247947 11/10/08 09:03 A. Camsonne Checked Argon/Ethane bottle
247946 11/10/08 08:34 adaq target: Spin Flip Report: for test
247942 11/10/08 08:07 Yi Zhang Owl shift summary
247939 11/10/08 07:43 ycwang something different with Bigbite wire Chamber RUN# 3152 3157
247938 11/10/08 07:30 ycwang pions and coincidence pions
247935 11/10/08 07:09 ycwang can gas Cherenkov make very clear electrons cut?
247934 11/10/08 07:06 Yi Zhang ROC9 rebooted
247931 11/10/08 06:51 Yi Zhang target-> 3He
247927 11/10/08 06:25 R. Michaels re: Qasy in BCM
247921 11/10/08 05:40 xqian About the BCM
247920 11/10/08 05:38 xqian It seems that there are a lot of "faked" trigger generated by the shower
247919 11/10/08 05:30 Yi Zhang Move target to BeO
247918 11/10/08 05:21 L. El Fassi Update : Hot Spot In BB Shower (run 3153 seems okay)
247917 11/10/08 05:14 Yi Zhang stop spin flip for beam trip
247915 11/10/08 05:00 L. El Fassi Update : Hot Spot In BB Shower (run 3152)
247910 11/10/08 04:01 adaq Target Field and Oven Temps snapshot
247909 11/10/08 04:01 adaq Target Cell Temps snapshot
247908 11/10/08 04:01 adaq Radiation and Ozone snapshot
247907 11/10/08 04:01 adaq Hall A Tools snapshot
247906 11/10/08 04:01 adaq BigBox Magnet snapshot
247905 11/10/08 04:01 adaq Beam Position and Current snapshot
247904 11/10/08 04:01 adaq Beam Energy in A and C snapshot
247901 11/10/08 03:33 Jin Laser Speacturm ScreenShot
247900 11/10/08 03:33 Jin Oven Temp and Laser ScreenShot
247898 11/10/08 03:32 adaq target: Spin Flip Report: Spin Flip Working OK
247896 11/10/08 03:16 Moffit Re: Feedback not working... Beam Charge asymmetry still small.
247893 11/10/08 02:57 Moffit HAPPEX Feedback: now using "bcm3"
247892 11/10/08 02:48 Jin Update : Hot Spot In BB Shower
247891 11/10/08 02:47 Moffit HAPPEX Feedback: not working
247888 11/10/08 02:21 adaq Hot spot in BB shower
247887 11/10/08 02:03 brads Target cell temperature changes a lot depending on ladder postion (w/o beam)
247884 11/10/08 01:51 brads BB ADC gates walking by 40ns+
Sunday
247876 11/09/08 23:53 xqian Swing Day Shift Summary
247875 11/09/08 23:50 adev BigBite WireChamber Measured Voltage change!
247873 11/09/08 23:42 adaq target: Spin Flip Report: Flip Paused due to beam off. Actived again before production
247869 11/09/08 23:34 Zhihong Ye Target moved to Polarized He3
247867 11/09/08 23:29 adev BigBite WireChamber Measured Voltage change!
247866 11/09/08 23:19 Zhihong Ye Target moved to Optics
247865 11/09/08 23:16 brads Target cell temperature fluctuations during beam off period
247864 11/09/08 23:10 brads Power cycled ROC8, ROC9, and ROC10
247859 11/09/08 22:56 brads Coda rate troubleshooting
247844 11/09/08 22:17 brads Coda rate slowed to a crawl
247843 11/09/08 22:11 Jin High Dead Time!
247841 11/09/08 22:00 Kalyan Re; ADC gate
247837 11/09/08 21:45 K Wang halla tools
247836 11/09/08 21:37 A. Camsonne ADC gate
247833 11/09/08 21:04 xqian BCM coefficient
247832 11/09/08 21:00 xqian It seems that only S1 ADC TDC got problems
247831 11/09/08 20:59 xqian ROC3 and 4 rebooted by themselves
247827 11/09/08 20:48 brads Please log the new BCM coeffs
247826 11/09/08 20:30 brads ADC gates should not be affected by coin. timing change(!?)
247823 11/09/08 20:11 brads FFB on standby, should it be?
247822 11/09/08 20:01 adaq Target Field and Oven Temps snapshot
247821 11/09/08 20:01 adaq Target Cell Temps snapshot
247820 11/09/08 20:01 adaq Radiation and Ozone snapshot
247819 11/09/08 20:01 adaq Hall A Tools snapshot
247818 11/09/08 20:01 adaq BigBox Magnet snapshot
247817 11/09/08 20:01 adaq Beam Position and Current snapshot
247816 11/09/08 20:01 adaq Beam Energy in A and C snapshot
247815 11/09/08 19:55 xqian Timing may screwed up in this run
247814 11/09/08 19:51 A. Camsonne T3 delay change
247812 11/09/08 19:44 Zhihong Ye Laser Pumping Spectrum Screen-shot
247810 11/09/08 19:41 Jin Target RTD & Laser ScreenShot @ Beam Comes
247809 11/09/08 19:23 xqian BCM Number Changed
247805 11/09/08 19:08 adev BigBite WireChamber Measured Voltage change!
247800 11/09/08 18:57 Zhihong Ye Target moved to He3, and Spin Flip
247799 11/09/08 18:48 Zhihong Ye Target moved to Optics
247798 11/09/08 18:36 adev BigBite WireChamber Measured Voltage change!
247795 11/09/08 18:31 Zhihong Ye Target Moved to Empty Target
247793 11/09/08 18:04 A. Camsonne Experiment status update : Hall A dump issue
247790 11/09/08 17:23 Kalyan Controlled access activity
247789 11/09/08 17:18 Yi Qiang Please following the special procedures to move target from/to He3 position
247788 11/09/08 16:50 A. Camsonne Requested controlled access while no beam
247785 11/09/08 16:08 Zhihong Ye Target Moved to Optics w/o hole
247781 11/09/08 16:02 Yi Day Shift Summary
247778 11/09/08 15:52 chench summary of effciency low peak position moving
247776 11/09/08 15:49 Chiranjib Hall A Tools
247775 11/09/08 15:32 adaq target: Spin Flip Report: Day shift report, transverse pumping
247774 11/09/08 15:19 Katich Hall A Tools Screen - Day shift
247773 11/09/08 15:17 V. Sulkosky Pion Cuts in Left HRS
247772 11/09/08 15:00 Chiranjib NO BEAM for almost more than 2 hours
247771 11/09/08 14:43 V. Sulkosky Pion Rejector Layer 2
247770 11/09/08 14:39 V. Sulkosky Pion Rejector Layer 1 Plots
247764 11/09/08 13:06 chench run3112 left arm VDC efficiency
247761 11/09/08 12:38 A. Camsonne Expect drop in efficiency because of gas shut down
247760 11/09/08 12:28 chench left arm efficiency for run3111
247759 11/09/08 12:11 Jack (Techs) Gas system shutdown
247756 11/09/08 12:03 Kalyan Plot of reaction point at target
247755 11/09/08 12:01 Yi Qiang A huge drop of left VDC effiencicy
247754 11/09/08 12:01 adaq Target Field and Oven Temps, snapshot
247753 11/09/08 12:01 adaq Target Cell Temps snapshot
247752 11/09/08 12:00 adaq Radiation and Ozone snapshot
247751 11/09/08 12:00 adaq Hall A Tools snapshot
247750 11/09/08 12:00 adaq BigBox Magnet snapshot
247749 11/09/08 12:00 adaq Beam Position and Current snapshot
247748 11/09/08 12:00 adaq Beam Energy in A and C snapshot
247747 11/09/08 11:55 chench run3110 left arm efficiency
247744 11/09/08 11:32 chench run3109 VDC efficiency
247743 11/09/08 11:18 Katich Correction to lumi 2 HV
247742 11/09/08 11:14 Katich Reset Lumi HVs
247741 11/09/08 11:12 chench run3108 left arm efficiency
247730 11/09/08 09:25 A. Camsonne Gas bottle swapped
247729 11/09/08 09:00 V. Sulkosky Start-of-Run tcl Modified
247724 11/09/08 08:15 cisbani Owl Shift Summary
247720 11/09/08 07:39 yawei dead time sometimes up to 44,usually less than 11
247715 11/09/08 06:46 cisbani Argon and CO2 gas pressure level low
247710 11/09/08 05:43 ycwang Coincidence peak in LHRS in Run3086
247707 11/09/08 05:22 Jin Huang RTD & Laser ScreenShot
247706 11/09/08 05:22 Jin Huang Laser absorption Screenshot
247705 11/09/08 05:20 adaq target: Spin Flip Report: Running OK.
247704 11/09/08 05:15 Jin Huang Physics Replay released, Run#3092 -> Golden Run
247701 11/09/08 04:33 cisbani spot++ run 3094
247698 11/09/08 04:12 cisbani HallA Tools
247697 11/09/08 04:10 cisbani bbps1:2001 HVOFF alarm, run 3094
247696 11/09/08 04:00 adaq Target Field and Oven Temps, snapshot
247695 11/09/08 04:00 adaq Target Cell Temps snapshot
247694 11/09/08 04:00 adaq Radiation and Ozone snapshot
247693 11/09/08 04:00 adaq Hall A Tools snapshot
247692 11/09/08 04:00 adaq BigBox Magnet snapshot
247691 11/09/08 04:00 adaq Beam Position and Current snapshot
247690 11/09/08 04:00 adaq Beam Energy in A and C snapshot
247671 11/09/08 00:01 Kebin Wang swing shift summary Nov 8
Saturday
247668 11/08/08 23:36 adaq target: Spin Flip Report: Running Smoothly
247667 11/08/08 23:34 Jin Laser Spectrum Screen Shot
247664 11/08/08 23:17 brads Added Target EPICs screens to auto_snap script
247661 11/08/08 22:31 xqian DEAD TIME on T5
247658 11/08/08 22:16 Yi Qiang EPICS strip chart screen shot
247648 11/08/08 20:22 K Wang Beamline HV tripped
247646 11/08/08 20:04 adaq Radiation and Ozone snapshot
247645 11/08/08 20:04 adaq Hall A Tools snapshot
247644 11/08/08 20:04 adaq BigBox Magnet snapshot
247643 11/08/08 20:04 adaq Beam Position and Current snapshot
247642 11/08/08 20:04 adaq Beam Energy in A and C snapshot
247639 11/08/08 19:36 Yi Qiang Suggestions on target spin flip while moving target
247634 11/08/08 18:54 Kalyan Coincidence time with 3He target
247631 11/08/08 18:31 Huaibo rate T3, T6
247626 11/08/08 17:50 V. Sulkosky BPM Calibration
247625 11/08/08 17:43 adaq Swiong shift tools
247620 11/08/08 16:52 Chiranjib EPR RF amplifier OK
247617 11/08/08 16:31 Chiranjib RTD readout screens
247616 11/08/08 16:29 Chiranjib Target spin flip screens/ laser absorption
247615 11/08/08 16:25 V. Sulkosky Harp Positions Versus EPICS (11072008 Bull's eye scan)
247613 11/08/08 16:20 Chiranjib timing offset for s2m
247609 11/08/08 15:55 E. Schulte Day Shift Summary, November 8, 2008
247607 11/08/08 15:55 E. Schulte Run Numbers for Day Shift starting at 11:40
247604 11/08/08 15:29 deng target reference cell RTD readbacks
247603 11/08/08 15:28 adaq Polarized He3 target RTD readbacks
247602 11/08/08 15:18 chench replay for run 3063
247593 11/08/08 13:48 adaq target: Spin Flip Report: 111
247592 11/08/08 13:39 E. Schulte Tools Screen
247589 11/08/08 13:11 V. Sulkosky Pion Rejector Layer 2 HV
247586 11/08/08 12:54 Kalyan Coincidence time with 3He target
247583 11/08/08 12:25 Kalyan BigBite scalers not counting T7
247580 11/08/08 12:12 debg Laser spectrum
247579 11/08/08 12:09 kalyan Trigger rates at 7uA on 3He target
247578 11/08/08 12:06 E. Schulte MCC having RF problems
247577 11/08/08 12:04 adaq Radiation and Ozone snapshot
247576 11/08/08 12:04 adaq Hall A Tools snapshot
247575 11/08/08 12:04 adaq BigBox Magnet snapshot
247574 11/08/08 12:04 adaq Beam Position and Current snapshot
247573 11/08/08 12:04 adaq Beam Energy in A and C snapshot
247572 11/08/08 11:48 E. Schulte Beam excersion on return of beam
247571 11/08/08 11:44 Kalyan T6 threshold changed to -120mV
247569 11/08/08 11:39 E. Schulte Spot++ After Extended Beam Down.
247566 11/08/08 11:29 deng Target moved to He3 cell
247565 11/08/08 11:11 Kalyan changed HV on few shower blocks
247562 11/08/08 10:46 E. Hall back in Beam Permit, but Accelerator has a Dead Magnet
247561 11/08/08 10:25 deng target moved to optics
247558 11/08/08 10:08 Kalyan Pedestals updated for BigBite ADC channels
247550 11/08/08 09:40 Chiranjib Target spin flip screen
247549 11/08/08 09:39 E. Schulte Hall in Controlled Access
247548 11/08/08 09:26 adaq target: Spin Flip Report: day shift
247547 11/08/08 09:17 E. Schulte MCC Needs to Work on Cooling Water Leak
247535 11/08/08 08:02 Cisbani Owl Shift Summary
247534 11/08/08 07:32 samuel NMR indication of transfer polarization is 42.7%
247533 11/08/08 07:01 Jin Huang Target Activity Summary @ Starting of Production
247532 11/08/08 06:55 ycwang Coincidence peak in run #3030
247525 11/08/08 05:53 xqian BigBite Wire Chamber background study
247524 11/08/08 05:16 online run3034,3035, 1700V chamber HV test,not 7000V
247523 11/08/08 05:14 adev BigBite WireChamber Measured Voltage change!
247519 11/08/08 05:03 adev BigBite WireChamber Measured Voltage change!
247514 11/08/08 04:54 online 1700V chamber HV test
247511 11/08/08 04:52 adev BigBite WireChamber Measured Voltage change!
247507 11/08/08 04:44 samuel Polarization reach to 40.5%
247505 11/08/08 04:42 adev BigBite WireChamber Measured Voltage change!
247498 11/08/08 04:31 adev BigBite WireChamber Measured Voltage change!
247497 11/08/08 04:31 adaq HallA Tools
247495 11/08/08 04:28 Cisbani spot++
247492 11/08/08 04:04 adaq Radiation and Ozone snapshot
247491 11/08/08 04:03 adaq Hall A Tools snapshot
247490 11/08/08 04:03 adaq BigBox Magnet snapshot
247489 11/08/08 04:03 adaq Beam Position and Current snapshot
247488 11/08/08 04:03 adaq Beam Energy in A and C snapshot
247481 11/08/08 03:12 Jin Temp. Golden Root File
247477 11/08/08 02:46 adev BigBite WireChamber Measured Voltage change!
247476 11/08/08 02:46 adaq target: Spin Flip Report: "Production run starts"; 20min Auto Spin Flip Actived
247472 11/08/08 02:35 adev BigBite WireChamber Measured Voltage change!
247470 11/08/08 02:32 adaq target: Spin Flip Report: Successful Spin Flip in 12 uA Beam. Pol.=36%
247467 11/08/08 02:25 adev BigBite WireChamber Measured Voltage change!
247463 11/08/08 02:14 adev BigBite WireChamber Measured Voltage change!
247462 11/08/08 02:13 Jin Huang Spin Flip software taking over the target
247457 11/08/08 02:03 adev BigBite WireChamber Measured Voltage change!
247455 11/08/08 02:02 adaq target: Spin Flip Report: Spin Flip System is taking over the target
247453 11/08/08 01:52 cisbani rate at raster 6.5x6 mcc
247450 11/08/08 01:43 xqian BB_PSh threshold is changed to 80mV from 20mV
247449 11/08/08 01:40 cisbani rate at raster 6x5 mcc
247444 11/08/08 01:33 cisbani Rate at Raster 4.5x4mcc
247441 11/08/08 01:24 adaq Rate at raster 3x3 mcc
247439 11/08/08 01:13 samuel move to He3 target position
247438 11/08/08 00:52 samuel move target to empty position
247437 11/08/08 00:52 Holmstrom / Bull-s Eyes Scan
247428 11/08/08 00:01 Holmstrom Swing Shift Summary
Friday
247426 11/07/08 23:50 adaq target: Spin Flip Report: New Code Test
247417 11/07/08 22:47 Kalyan Coincidence time peak with 5 pass beam
247414 11/07/08 21:59 yez Target Moved to Optics w/o hole
247413 11/07/08 21:55 A. Camsonne Reference cell well aligned good for setting ion chamber trip point
247411 11/07/08 21:50 A. Camsonne Rates reference cell N2 50 psi 2uA
247409 11/07/08 21:47 A. Camsonne Rates N2 50 psi 2 uA raster 4x4mm
247405 11/07/08 21:41 Yi Qiang Correction Tests
247404 11/07/08 21:40 Yi Qiang Correction Tests
247403 11/07/08 21:39 A. Camsonne Rates reference cell raster 3x3mm
247402 11/07/08 21:39 yez Target Moved to Refference Cell
247400 11/07/08 21:34 yez Target Moved to Refference Cell
247399 11/07/08 21:30 Yi Qiang Filled Reference Cell to 25 psig with N2
247394 11/07/08 21:02 brads (Tech) RHRS VDC discriminator problem?
247390 11/07/08 20:03 adaq Radiation and Ozone snapshot
247389 11/07/08 20:03 adaq Hall A Tools snapshot
247388 11/07/08 20:03 adaq BigBox Magnet snapshot
247387 11/07/08 20:03 adaq Beam Position and Current snapshot
247386 11/07/08 20:03 adaq Beam Energy in A and C snapshot
247384 11/07/08 20:02 A. Camsonne Update on HARP
247380 11/07/08 18:59 Kalyan Trigger rates at 10uA on carbon foil
247378 11/07/08 18:45 Holmstrom Harp Scan software problem
247377 11/07/08 18:22 Yi Zhang Target move to BeO target to do harpscan
247376 11/07/08 18:20 Yi Zhang Reference cell filled with N2(1 psi)
247375 11/07/08 18:15 yez Target Moved to Refference Cell
247374 11/07/08 17:39 Holmstrom Target Masking Bug?
247373 11/07/08 17:23 Yi Qiang Target move to empty
247369 11/07/08 15:55 E. Schulte Day Shift Summary, November 7, 2008
247366 11/07/08 15:37 A. Camsonne Preliminary results of target survey notes by Brad
247365 11/07/08 15:31 V. Sulkosky New Database Directory
247364 11/07/08 15:24 A. Camsonne Fan unit added above the helicity crate in counting house
247363 11/07/08 15:16 V. Sulkosky BPM Pedestal Run 2991
247362 11/07/08 15:05 V. Sulkosky BPM and Raster Work II
247361 11/07/08 14:34 Sulkosky/Sch BPM and Raster Work
247358 11/07/08 13:49 brads LHRS set to 2.35 GeV/c
247357 11/07/08 13:45 brads BigBite is back at 710A
247356 11/07/08 13:33 E. Cisbani / Switch network port
247354 11/07/08 13:30 Xiaodong Jia Need to fine tune left_HRS setting, to 2.35 GeV/c.
247353 11/07/08 13:30 Jack (Techs) Changed UPS
247352 11/07/08 13:10 Kalyan RE: BB chamber dark current check. Beam current:4uA
247351 11/07/08 12:55 A. Camsonne Transvsersity experiment status and reminder experiment meeting at 4PM in counting house 2nd floor meeting room
247345 11/07/08 12:26 Yi Qiang Left HRS Q1 is ON (set to 2.4GeV)
247340 11/07/08 12:07 Yi Qiang Left HRS Q1 and BigBite Magnets are OFF
247337 11/07/08 12:03 adaq Radiation and Ozone snapshot
247336 11/07/08 12:03 adaq Hall A Tools snapshot
247335 11/07/08 12:03 adaq BigBox Magnet snapshot
247334 11/07/08 12:03 adaq Beam Position and Current snapshot
247333 11/07/08 12:03 adaq Beam Energy in A and C snapshot
247320 11/07/08 10:44 brads Raster ADC change: heads up to analysis folks
247314 11/07/08 10:27 cusanno RICH dymanic library for positive polarity
247312 11/07/08 10:24 brads Must do Bull's Eye scan when we come back
247311 11/07/08 10:22 brads Hot spots found during
247309 11/07/08 10:11 Jin Target Ladder No Respond, Reboot IOC again
247308 11/07/08 10:05 Jin Target IOC Reboot
247307 11/07/08 09:45 V. Sulkosky Raster Overflowing and Underflowing
247306 11/07/08 09:25 Chiranjib/Ji SPIN UP @220 C
247305 11/07/08 09:24 E. Schulte Requested Controlled Access to Work on Compton
247304 11/07/08 09:22 adev BigBite WireChamber Measured Voltage change!
247303 11/07/08 09:12 Jin/Chiranji EPR Error, Pol is low now
247302 11/07/08 09:11 adev BigBite WireChamber Measured Voltage change!
247301 11/07/08 08:37 brads Software problem with target ladder control
247300 11/07/08 08:36 E. Schulte Tools screen for 3He
247299 11/07/08 08:34 Yi Zhang 3He laser off
247298 11/07/08 08:24 Kalyan Beam energy for runs 2936 to 2970 is 5.889GeV and not 2.39GeV
247297 11/07/08 08:16 ycwang Turn off HV of RICH for Target reason.
247296 11/07/08 08:12 E. Schulte Target Motion Control Problems
247294 11/07/08 08:06 E. Cisbani Owl Shift Summary
247291 11/07/08 07:45 yawei run2969, ROC9 no respond.
247288 11/07/08 07:38 yawei run2965,2966,2967,ROC9 no respond
247280 11/07/08 06:55 Huaibo no "optics with hole" selection
247279 11/07/08 06:50 Huaibo no "optics with hole" seclection
247268 11/07/08 05:46 ycwang A1 sum for check run#2512
247264 11/07/08 05:29 A. Camsonne Limited channel readout per ADC for RICH to 60
247261 11/07/08 05:09 E. Cisbani Left arm HV
247257 11/07/08 05:02 adaq Hall A tools
247255 11/07/08 04:59 adev BigBite WireChamber Measured Voltage change!
247254 11/07/08 04:51 Yi Zhang BB chamber test
247253 11/07/08 04:49 adev BigBite WireChamber Measured Voltage change!
247252 11/07/08 04:39 A. Camsonne recorded raster value at 6 GeV overflowing ADC ?
247250 11/07/08 04:35 ycwang Coincidence peak in run 2946
247249 11/07/08 04:34 A. Camsonne ROC9 rebooted by itself
247245 11/07/08 04:17 ycwang Target position for RUN 2946
247244 11/07/08 04:06 adev BigBite WireChamber Measured Voltage change!
247243 11/07/08 04:03 adaq Radiation and Ozone snapshot
247242 11/07/08 04:03 adaq Hall A Tools snapshot
247241 11/07/08 04:03 adaq BigBox Magnet snapshot
247240 11/07/08 04:03 adaq Beam Position and Current snapshot
247239 11/07/08 04:02 adaq Beam Energy in A and C snapshot
247236 11/07/08 03:56 adev BigBite WireChamber Measured Voltage change!
247235 11/07/08 03:55 ycwang lots of new warnings when replay
247234 11/07/08 03:55 Yi Zhang BB chamber dark current check. Beam current:4uA
247232 11/07/08 03:45 adev BigBite WireChamber Measured Voltage change!
247229 11/07/08 03:40 Yi Zhang BB chamber dark current under 3 uA
247226 11/07/08 03:34 Yi Zhang BB chamber dark current under 3 uA
247224 11/07/08 03:24 adev BigBite WireChamber Measured Voltage change!
247221 11/07/08 03:21 E. Cisbani Prescale all triggers but t3 to fix raster problem did not work
247219 11/07/08 03:18 Yi Zhang BB chamber dark current under 2 uA
247218 11/07/08 03:13 adev BigBite WireChamber Measured Voltage change!
247217 11/07/08 02:43 Huaibo move target to optics
247216 11/07/08 02:36 Yi Zhang THaRaster::"rb.Raster" problem
247215 11/07/08 02:18 E. Cisbani Raster Size on Holey Carbon
247212 11/07/08 02:13 online run2838,2839,target is optics with hole,sorry
247209 11/07/08 01:47 online target is Optics with hole,sorry
247208 11/07/08 01:43 online target is optics with hole, sorry
247201 11/07/08 01:13 E. Cisbani Harp Scan Ref
247200 11/07/08 01:08 Huaibo Move target to optics with hole
247199 11/07/08 00:58 brads Beam energy fluctuations
247198 11/07/08 00:51 brads 100x to 1000x the radiation dose during early 5 pass
247197 11/07/08 00:36 E. Cisbani spot++ BeO target 2uA
247192 11/07/08 00:27 brads Update on target vs. beam positions
247190 11/07/08 00:09 Ibrahim Swing Shift Summary
Thursday
247189 11/06/08 23:58 Huaibo Move target to optics
247188 11/06/08 23:15 cusanno RICH software updated
247187 11/06/08 23:07 cusanno RICH tripped, presently HV is OFF
247186 11/06/08 22:17 V. Sulkosky Left HRS HV GUI
247185 11/06/08 21:25 brads Wien angle setting: 90% source polarization to Hall
247184 11/06/08 20:59 brads BB MWDC ethanol bubbler engaged
247183 11/06/08 20:49 brads BB Shower threshold set to -80mV.
247182 11/06/08 20:44 adev BigBite WireChamber Measured Voltage change!
247180 11/06/08 20:06 doug Test of Web Entry with Photo
247179 11/06/08 20:02 adaq Radiation and Ozone snapshot
247178 11/06/08 20:02 adaq Hall A Tools snapshot
247177 11/06/08 20:02 adaq BigBox Magnet snapshot
247176 11/06/08 20:02 adaq Beam Position and Current snapshot
247175 11/06/08 20:02 adaq Beam Energy in A and C snapshot
247174 11/06/08 20:01 brads NOT steering through the Compton chicane
247171 11/06/08 19:53 Kalyan ROC8 pedestals updated
247170 11/06/08 19:51 brads testing weblog 3
247169 11/06/08 19:38 xiaohui Target moved to empty
247168 11/06/08 18:13 brads Web-based halog disabled for now
247167 11/06/08 17:24 V. Sulkosky Web based Halog Not Working
247166 11/06/08 17:22 V. Sulkosky Timing offsets for S2m updated
247165 11/06/08 17:18 V. Sulkosky Timing offsets for S1 updated
247164 11/06/08 17:08 Ibrahim Left Dipole Polarity Chaged to Negative
247162 11/06/08 16:01 adev BigBite WireChamber Measured Voltage change!
247158 11/06/08 14:25 Kalyan Run 2932 is pedestals for ROC8
247151 11/06/08 12:10 E. Schulte Tools Screen
247150 11/06/08 12:02 adaq Radiation and Ozone snapshot
247149 11/06/08 12:02 adaq Hall A Tools snapshot
247148 11/06/08 12:02 adaq BigBox Magnet snapshot
247147 11/06/08 12:02 adaq Beam Position and Current snapshot
247146 11/06/08 12:02 adaq Beam Energy in A and C snapshot
247136 11/06/08 11:04 brads Target -> He3 Cell
247134 11/06/08 11:00 brads Target -> Reference cell
247133 11/06/08 10:45 E. Schulte Gas Shed Air Flow Alarm
247132 11/06/08 10:41 Jin Huang Target Summary during Spin flip test @ Delta Resonance & further
247131 11/06/08 10:41 Yi Qiang Target to Optics (BeO)
247130 11/06/08 10:38 adaq target: Spin Flip Report: Some 1.6% holding filed drift begin from 9:55AM completely screwed up the polarization
247128 11/06/08 10:30 E. Schulte Target FP Loss Test Needs to be Performed Again
247127 11/06/08 10:30 brads LHRS moved to 16 degrees
247126 11/06/08 10:01 Yi Qiang Lasers Locked
247125 11/06/08 09:26 Jin Huang Spin Signal Locked to Spin Up = 1; Spin Down =0
247124 11/06/08 08:52 E. Schulte Hall to go to Controlled Access
247123 11/06/08 08:52 adaq target: Spin Flip Report: End of beam time
247121 11/06/08 08:40 gaskell HAPPEX run 30379 during run 2923
247118 11/06/08 08:30 Jin Huang he3 vertex and Inv Mass in 2 pass beam
247117 11/06/08 08:26 ycwang After RUN 2917, the ADCs of S1 and S2m have two peaks.
247116 11/06/08 08:22 ycwang After RUN 2917, the ADCs of S1 and S2m have two peaks.
247115 11/06/08 08:17 Jin Huang Real Spin Signal reads OK in DAQ
247113 11/06/08 08:04 Yi Zhang owl shift summay
247110 11/06/08 07:47 Yi Zhang hall a tools screen
247109 11/06/08 07:31 Jin Huang Laser Absorption Spectrum
247104 11/06/08 05:39 jpchen Raster size for beam on pol He3
247103 11/06/08 05:38 Yi Zhang BB chamber HV triped
247102 11/06/08 05:38 Yi Zhang spot++ of he3 run#2919,
247101 11/06/08 05:38 Jin Huang Automatic Spin Flip ON
247100 11/06/08 05:35 Yi Zhang BB chamber HV triped
247098 11/06/08 05:30 adev BigBite WireChamber Measured Voltage change!
247097 11/06/08 05:26 ycwang RICH Pedestal RUN with beam RUN # 2913
247096 11/06/08 05:23 A. Camsonne RICH runs summary and test
247095 11/06/08 05:19 adev BigBite WireChamber Measured Voltage change!
247094 11/06/08 05:18 adaq target: Spin Flip Report: Spin Flip System Fully Enabled with Transverse Pumping
247093 11/06/08 05:13 samuel 1 atm Nytrogemn gas in reference cell.
247092 11/06/08 05:09 samuel move target to He3 Position.
247086 11/06/08 04:54 A. Camsonne RICH run 2916
247079 11/06/08 04:38 A. Camsonne RICH good pedestal run 2913
247070 11/06/08 04:05 adaq LHRS momentum change to 600 MeV/C
247069 11/06/08 04:02 jpchen Use N2 for RICH study
247068 11/06/08 04:02 adaq Radiation and Ozone snapshot
247067 11/06/08 04:02 adaq Hall A Tools snapshot
247066 11/06/08 04:02 adaq BigBox Magnet snapshot
247065 11/06/08 04:02 adaq Beam Position and Current snapshot
247064 11/06/08 04:02 adaq Beam Energy in A and C snapshot
247062 11/06/08 03:59 ycwang No signal on BB Sum TDCs etc for Run #2902
247058 11/06/08 03:45 ycwang Run #2902 coincidence peak
247057 11/06/08 03:40 samuel fill reference cell with N2 to pressure 50PSI
247051 11/06/08 02:59 xqian The t0 offset for wire chamber for t5.
247049 11/06/08 02:41 adev BigBite WireChamber Measured Voltage change!
247043 11/06/08 01:45 xqian Delta in BigBite
247042 11/06/08 01:43 xqian BigBite Wire Chamber running condition study
247040 11/06/08 01:28 adev BigBite WireChamber Measured Voltage change!
247039 11/06/08 01:26 kalyan late entry : T3 delay changed
247036 11/06/08 00:51 Yi Qiang Typo on the error
247035 11/06/08 00:46 adev BigBite WireChamber Measured Voltage change!
247034 11/06/08 00:41 Yi Qiang Max Polarization for current Transverse and Vertical Pumping
247033 11/06/08 00:41 Yi Qiang Max Polarization for current Transverse and Vertical Pumping
247031 11/06/08 00:01 Ibrahim Swing Shift Summary
Wednesday
247027 11/05/08 23:27 Chiranjib/Xi Temperature Test
247026 11/05/08 23:27 Chiranjib/Xi Temperature Test
247025 11/05/08 23:04 xiaohui Temp. test done
247014 11/05/08 21:58 adev BigBite WireChamber Measured Voltage change!
247013 11/05/08 21:54 xiaohui 3He temp. test started
247004 11/05/08 21:37 adev BigBite WireChamber Measured Voltage change!
247002 11/05/08 21:30 xiaohui target moved to reference
246998 11/05/08 21:05 adev BigBite WireChamber Measured Voltage change!
246997 11/05/08 21:05 Ibrahim Harp Scan Results (20:35)
246995 11/05/08 20:55 adev BigBite WireChamber Measured Voltage change!
246992 11/05/08 20:53 chench 3x3 raster size check
246989 11/05/08 20:37 chench beam position confirm
246987 11/05/08 20:33 adev BigBite WireChamber Measured Voltage change!
246985 11/05/08 20:02 adaq Radiation and Ozone snapshot
246984 11/05/08 20:02 adaq Hall A Tools snapshot
246983 11/05/08 20:01 adaq BigBox Magnet snapshot
246982 11/05/08 20:01 adaq Beam Position and Current snapshot
246981 11/05/08 20:01 adaq Beam Energy in A and C snapshot
246976 11/05/08 19:47 adev SCALER = -1
246973 11/05/08 19:38 adaq target: Spin Flip Report: Tested for Trans Spin Flip
246970 11/05/08 19:34 Xiaohui Target moved to optics
246967 11/05/08 19:22 A. Camsonne new CODA configuration for pedestal
246964 11/05/08 19:09 adev BigBite WireChamber Measured Voltage change!
246963 11/05/08 19:06 brads BigBite Sieve in the OUT position
246962 11/05/08 19:05 xiaohui Target moved to empty
246953 11/05/08 18:46 Ibrahim Gas Shed Airflow Alarm
246951 11/05/08 18:44 chench ref. cell H2 cosmic run
246949 11/05/08 18:25 V. Sulkosky Harp A Scan Analysis from 16:58
246947 11/05/08 18:12 chench To modify the halog
246946 11/05/08 17:58 chench reference cell He
246945 11/05/08 17:55 chench Ref. Cell He
246944 11/05/08 17:47 Kalyan shift in coincidence peak by 16ns
246943 11/05/08 17:45 adev BigBite WireChamber Measured Voltage change!
246937 11/05/08 17:26 Katich BeO cross-hair position
246936 11/05/08 17:14 adev BigBite WireChamber Measured Voltage change!
246935 11/05/08 17:11 xiaohui target moved to reference cell
246934 11/05/08 17:00 brads Harp A Scan (needs analysis)
246933 11/05/08 16:59 brads Harp B scaA
246932 11/05/08 16:52 brads Harp B scan
246931 11/05/08 16:47 brads BPMs vs. BeO spot image
246929 11/05/08 16:05 V. Sulkosky Preliminary Left HRS Detector Position Measurements
246925 11/05/08 15:58 E. Schulte Day Shift Summary, November 5, 2008
246924 11/05/08 15:42 Chiranjib/yi EPR measurement
246923 11/05/08 15:32 gaskell BB tracking OK
246920 11/05/08 15:04 Kalyan Note on target-helicity gated scalers
246919 11/05/08 14:54 E. Schulte Gas Shed Ariflow Alarm
246918 11/05/08 14:49 gaskell Run 2863 is junk
246915 11/05/08 14:28 E. Schulte Spot++ for run 2862
246914 11/05/08 14:26 Yi Qiang Target to Reference (H2 135 psig)
246912 11/05/08 14:14 Yi Qiang Filled reference cell with H2 to 135 psig
246910 11/05/08 14:11 E. Schulte Run # 2861 No Raster Spot++
246908 11/05/08 14:02 Kalyan T3 delay changed
246906 11/05/08 13:57 Yi Qiang Target to Optics (BeO)
246905 11/05/08 13:55 A. Camsonne RICH crl change and logic change
246904 11/05/08 13:53 gaskell SPOT++ with holey target: raster=4x4
246899 11/05/08 13:39 gaskell Harp scan 1H03B
246898 11/05/08 13:39 R. Michaels xcefdmp updated
246897 11/05/08 13:37 gaskell Harp scan - 1H03A
246896 11/05/08 13:35 yez Target Moved to Optics w hole
246895 11/05/08 13:27 E. Schulte Tools screen with new positions.
246893 11/05/08 13:25 yez Target Moved to Empty Position
246892 11/05/08 13:24 gaskell spot++ screen grab
246890 11/05/08 13:20 E. Schulte Picture after position change
246889 11/05/08 13:10 E. Schulte Starting positions on BeO Target
246888 11/05/08 13:05 gaskell Runs 2854-2857 DAQ tests
246887 11/05/08 13:02 yez Target Moved to Optics w/o hold
246886 11/05/08 13:00 E. Schulte Wien Angle Changed
246885 11/05/08 12:56 Kalyan coincidence time peak
246883 11/05/08 12:54 V. Sulkosky Added Trigger T5 to 1875 TDC
246877 11/05/08 12:18 Yi Qiang Laser absorption spectrum for transverse pumping
246876 11/05/08 12:14 Yi Qiang Continue Transver pumping
246875 11/05/08 12:01 adaq Radiation and Ozone snapshot
246874 11/05/08 12:01 adaq Hall A Tools snapshot
246873 11/05/08 12:01 adaq BigBox Magnet snapshot
246872 11/05/08 12:01 adaq Beam Position and Current snapshot
246871 11/05/08 12:01 adaq Beam Energy in A and C snapshot
246868 11/05/08 11:54 E. Schulte Tools Screen
246867 11/05/08 11:49 E. Schulte Reference Cell Controller Back in Remote Mode
246866 11/05/08 11:43 brads Confirmed target alignment with laser level
246865 11/05/08 11:40 brads H2 pressure regulator swapped out
246863 11/05/08 11:23 E. Schulte Target Collimators are In
246862 11/05/08 11:04 V. Sulkosky Beam positions from run 2853
246861 11/05/08 11:00 V. Sulkosky BPM Calibration Completed
246860 11/05/08 10:47 E. Schulte Reference Cell Control Panel in Local
246859 11/05/08 10:34 Yi Qiang Laser locked for openning the enclosure
246858 11/05/08 10:32 E. Schulte Coincidence Time Location Tests
246857 11/05/08 10:11 brads Target alignment photos from Tuesday
246856 11/05/08 10:06 E. Schulte MCC to do Wein angle change
246855 11/05/08 10:04 E. Schulte Hall in Controlled Access
246851 11/05/08 09:38 E. Schulte 2uA RTD screens
246846 11/05/08 08:55 xqian Elastic proton in Left HRS, 2pass run 2847
246845 11/05/08 08:53 Yi Qiang Orientation of Spot++ plot and BeO photo
246844 11/05/08 08:23 xqian elastic electron at two pass
246840 11/05/08 08:01 Yi Zhang Owl shift summary
246839 11/05/08 07:55 xqian Sieve pattern with 2pass H2 data elastic
246838 11/05/08 07:53 xqian Unsuccesful extraction of the coincidence peak
246837 11/05/08 07:48 Yi Zhang Reference cell Raster size test
246836 11/05/08 07:36 adev BigBite WireChamber Measured Voltage change!
246835 11/05/08 07:33 Yi Zhang HallA main tools
246834 11/05/08 07:32 Yi Zhang BCM&BPM
246829 11/05/08 06:54 adev BigBite WireChamber Measured Voltage change!
246827 11/05/08 06:43 adev BigBite WireChamber Measured Voltage change!
246826 11/05/08 06:21 Jin Cannot fill in H2 to Reference Cell to Required Pressure
246825 11/05/08 06:17 Huaibo move target to reference cell(H2)
246824 11/05/08 06:12 Jin Huang Look into the second peak on BB momentum spectrum
246822 11/05/08 05:53 Yi Zhang raster size 1x1 mm
246821 11/05/08 05:51 Yi Zhang raster 3x3 mm(4x4 in MCC unit) run#2846
246819 11/05/08 05:44 Yi Zhang raster size
246817 11/05/08 05:36 Jin Might be able to further center spot
246816 11/05/08 05:29 adev BigBite WireChamber Measured Voltage change!
246812 11/05/08 05:19 adev BigBite WireChamber Measured Voltage change!
246811 11/05/08 05:10 Yi Zhang problem with spot++?
246810 11/05/08 05:01 Huaibo Move target to optics
246807 11/05/08 04:49 xqian Sieve at 2 pass with multi-carbon target
246800 11/05/08 04:19 xqian The charge asymmetry is amplified by factor of 10
246798 11/05/08 04:16 xqian False asymmetry study
246795 11/05/08 04:09 Jin Hole on Carbon
246794 11/05/08 04:05 brads spot++ BPMA_y is very different than 'Tools' BPMA_y
246793 11/05/08 04:05 adev BigBite WireChamber Measured Voltage change!
246792 11/05/08 04:01 adaq Radiation and Ozone snapshot
246791 11/05/08 04:01 adaq Hall A Tools snapshot
246790 11/05/08 04:01 adaq BigBox Magnet snapshot
246789 11/05/08 04:01 adaq Beam Position and Current snapshot
246788 11/05/08 04:01 adaq Beam Energy in A and C snapshot
246787 11/05/08 03:54 adev BigBite WireChamber Measured Voltage change!
246785 11/05/08 03:48 Huaibo Move target to optics with hole
246784 11/05/08 03:44 xqian vertex z for beam position
246783 11/05/08 03:43 adev BigBite WireChamber Measured Voltage change!
246781 11/05/08 03:36 adaq Vertex Z for spot++ run
246779 11/05/08 03:32 adev BigBite WireChamber Measured Voltage change!
246776 11/05/08 03:20 brads BeO snapshot with beam at -3.1, 3.2mm
246775 11/05/08 03:16 Huaibo Move target to optics
246774 11/05/08 03:10 adaq Carbon Hole Spot++ With Raster
246769 11/05/08 02:37 xqian Update on longitudinal asymmetry
246768 11/05/08 02:36 Huaibo Target moves to optics with hole
246767 11/05/08 02:33 xqian Transverse asymmetry in the delta region
246766 11/05/08 02:32 Huaibo Target moves to optics
246761 11/05/08 02:06 Yi Zhang spot positions @ 2 pass
246760 11/05/08 01:59 Huaibo Target moves to optics with hole
246759 11/05/08 01:59 Yi Zhang spot++ after beam postion turning
246756 11/05/08 01:35 Yi Zhang Harp Scan
246753 11/05/08 00:57 Huaibo Target move to optics
246752 11/05/08 00:38 Jin Huang Couple Tests On Target
246751 11/05/08 00:10 Jin Huang Transverse Pumping Status
246750 11/05/08 00:02 gilman swing shift summary, Nov 4, 2008
Tuesday
246748 11/04/08 23:55 brads BB Shower thresholds (T1, T6) set to -80mV
246747 11/04/08 23:08 adev BigBite WireChamber Measured Voltage change!
246746 11/04/08 23:01 gilman beam coming back, HV trip
246744 11/04/08 22:39 Yi Qiang Target wall collimators were taken out
246743 11/04/08 22:37 Yi Qiang Target activities during controlled access
246742 11/04/08 22:28 gilman controlled access over
246741 11/04/08 22:12 gilman moved spectrometer
246740 11/04/08 21:48 brads Closed valve on 3He bottle
246739 11/04/08 21:46 Jin Huang Vertical Pumping Laser Fiber Status
246738 11/04/08 21:09 gilman Ready to move spectrometer to 46.37 degr
246737 11/04/08 20:19 Techs Q1 slow ramp rate
246736 11/04/08 20:08 gilman Q1 problem
246735 11/04/08 20:01 adaq Radiation and Ozone snapshot
246734 11/04/08 20:01 adaq Hall A Tools snapshot
246733 11/04/08 20:01 adaq BigBox Magnet snapshot
246732 11/04/08 20:01 adaq Beam Position and Current snapshot
246731 11/04/08 20:00 adaq Beam Energy in A and C snapshot
246727 11/04/08 19:31 gilman setting spectrometer to + polarity, Q1problem
246726 11/04/08 18:36 V. Sulkosky Re: Restored to Previous ADC channels for S2m
246725 11/04/08 18:33 brads LHRS switched to POSITIVE polarity
246724 11/04/08 18:30 brads Center beam on BeO cross hairs, NOT holey carbon
246723 11/04/08 18:29 xiaohui target moved to empty
246722 11/04/08 18:28 brads New target ladder installed. NO upstream C foil
246721 11/04/08 18:25 xiaohui Target moved to reference cell
246712 11/04/08 17:41 adev BigBite WireChamber Chamber Turn on/off!
246710 11/04/08 17:40 xiaohui target IOC rebooted
246707 11/04/08 17:38 Techs Polarity change
246704 11/04/08 17:31 adev BigBite WireChamber Chamber Turn on/off!
246694 11/04/08 16:47 brads MWDC bubblers still bubbling
246693 11/04/08 16:46 brads BB Sieve IN
246689 11/04/08 16:28 adev BigBite GUI Stopped!
246685 11/04/08 16:18 adev BigBite GUI Stopped!
246680 11/04/08 16:01 E. Schulte Day Shift Summary, November 4, 2008: Election Day!
246678 11/04/08 15:59 brads Damage probably occured between 1am--3am, Monday Nov 3
246672 11/04/08 15:29 brads Damaged target ladder
246671 11/04/08 15:17 Kalyan database updated
246668 11/04/08 15:11 xqian BigBite Wire Chamber Update
246659 11/04/08 14:48 adev BigBite WireChamber Measured Voltage change!
246655 11/04/08 14:31 A. Camsonne RICH DAQ test
246650 11/04/08 13:33 adev BigBite WireChamber Threshold change!
246649 11/04/08 13:32 E. Schulte Tools Screen
246648 11/04/08 13:23 adev BigBite WireChamber Threshold change!
246647 11/04/08 13:12 adev BigBite WireChamber Measured Current change!
246640 11/04/08 12:51 adev BigBite WireChamber Threshold change!
246639 11/04/08 12:35 brads Gas shed ethane/argon mix line connected to BB MWDCs (ethanol bubbler bypassed)
246638 11/04/08 12:31 brads Hall in Controlled Access since 9am (late entry)
246637 11/04/08 12:28 R. Michaels R-HRS ROC1 2V busted
246636 11/04/08 12:13 R. Michaels R-HRS work
246635 11/04/08 12:03 V. Sulkosky Run 2794 was a Pedestal Run
246634 11/04/08 12:00 adaq Radiation and Ozone snapshot
246633 11/04/08 12:00 adaq Hall A Tools snapshot
246632 11/04/08 12:00 adaq BigBox Magnet snapshot
246631 11/04/08 12:00 adaq Beam Position and Current snapshot
246630 11/04/08 12:00 adaq Beam Energy in A and C snapshot
246621 11/04/08 11:22 V. Sulkosky New Analyzer Directory Created
246620 11/04/08 11:18 V. Sulkosky Moved S2m Right ADC Signals
246618 11/04/08 11:01 R. Michaels Re: TIR helicity (L-HRS)
246615 11/04/08 10:19 Chiranjib LaserS off
246611 11/04/08 10:10 A. Camsonne Compton computer crashed
246610 11/04/08 09:25 A. Camsonne Compton broadcast handler issue
246609 11/04/08 09:19 Lingyan Zhu target: move to empty
246608 11/04/08 09:15 A. Camsonne Compton centra crystal check
246607 11/04/08 09:07 Jin Huang NMR -> Pol. =52% at end of transverse run
246602 11/04/08 08:10 Jin Huang Screen Shot for Transverse Minus Spin State
246601 11/04/08 08:05 Yi Zhang Owl Shift summary
246600 11/04/08 08:05 Jin Huang Successful Manual Spin Flip (Pol. = 50%)
246599 11/04/08 08:03 zhangyi halla main tools
246597 11/04/08 07:35 samuel Turn on lasers #1,3,4,5,6
246596 11/04/08 07:26 samuel turn off lasers#1,3,4,5,6
246593 11/04/08 07:00 Huaibo G0 helicity change
246590 11/04/08 06:37 Jin Run#2783 is NOT junk run
246584 11/04/08 06:06 Jin NMR Measurement, 49%
246580 11/04/08 05:43 Moffit Beam Charge Asymmetry Feedback on.
246574 11/04/08 04:12 samuel move to He3 polarized target position
246573 11/04/08 04:06 Yi Zhang different VDC response in delta ragion
246572 11/04/08 04:00 adaq Radiation and Ozone snapshot
246571 11/04/08 04:00 adaq Hall A Tools snapshot
246570 11/04/08 04:00 adaq BigBox Magnet snapshot
246569 11/04/08 04:00 adaq Beam Position and Current snapshot
246568 11/04/08 04:00 adaq Beam Energy in A and C snapshot
246567 11/04/08 03:36 Jin Huang Target Long -> Trans
246566 11/04/08 03:28 samuel move target to optics position
246561 11/04/08 02:45 samuel move target to He3 cell position.
246560 11/04/08 02:43 samuel move to optics position and put lasers#1,3,4,5,6 on
246555 11/04/08 02:26 xqian With Helicity change sign
246554 11/04/08 02:25 samuel position for making NMR pol. measurement.
246549 11/04/08 01:07 samuel move to He3 cell
246548 11/04/08 01:02 samuel move to pickup coil position
246547 11/04/08 01:01 xqian 3He pressure curve makes sense
246545 11/04/08 00:47 xqian N2 pressure curve make sense.
246544 11/04/08 00:47 xqian N2 pressure curve make sense.
246543 11/04/08 00:45 Jin Temporary Golden Root File Run#2769 for the Night
246540 11/04/08 00:10 xqian more data from Nov. 3rd
246538 11/04/08 00:01 gilman swing shift summary Nov 3, 2008
Monday
246537 11/03/08 23:38 yez NMR Measurement: 42.1%
246535 11/03/08 23:33 V. Sulkosky BCM in End-of-Run File
246533 11/03/08 23:03 Kalyan Re: BCM gated scaler data are non-zero, but look bogus too
246531 11/03/08 22:47 xqian About the BCM
246529 11/03/08 22:44 Jin Re: BCM gated scaler data are non-zero, but look bogus too
246528 11/03/08 22:26 yez Target Moved to Polarized He3 before run 2769
246527 11/03/08 22:16 brads BCM gated scaler data are non-zero, but look bogus too
246525 11/03/08 21:57 gilman wave plate inserted
246524 11/03/08 21:50 yez Corrected: Target Moved to Pickup Coils for NMR Measurement
246523 11/03/08 21:46 yez Target Moved to He3, Longitudinal+, NMR Measurement=45.4%
246522 11/03/08 21:30 gilman pressure curves done
246516 11/03/08 20:53 Jin Target Signal
246513 11/03/08 20:43 xqian Polarized 3He run today will need special treatment
246507 11/03/08 20:08 yez Target Moved back to Reference Cell.
246506 11/03/08 20:04 yez Target Moved to He3 Cell for presure carve measurement.
246505 11/03/08 20:01 adaq Radiation and Ozone snapshot
246504 11/03/08 20:01 adaq Hall A Tools snapshot
246503 11/03/08 20:01 adaq BigBox Magnet snapshot
246502 11/03/08 20:01 adaq Beam Position and Current snapshot
246500 11/03/08 20:01 adaq Beam Energy in A and C snapshot
246499 11/03/08 20:01 yez NMR Mesearment, 43.3%
246498 11/03/08 19:55 gilman Hall A tools for shift checklist
246489 11/03/08 18:54 xqian BigBite Wire Chamber Noise Problem
246484 11/03/08 18:12 V. Sulkosky Fine Tune of S2m HVs
246480 11/03/08 17:51 yez Corretecd: Last NMR measurement result is 33.64%
246479 11/03/08 17:50 Jin Huang replay cratemap update
246478 11/03/08 17:42 yez Target at Reference Cells, Filed rotated to Transvers+, Laser on
246477 11/03/08 17:20 NMR Mesearment, target location pickup, RF vertical, pickup transverse
246472 11/03/08 16:03 E. Schulte 3He Longitudinal Polarization
246470 11/03/08 16:01 E. Schulte Day Shift Summary, November 3, 2008
246469 11/03/08 16:00 E. Schulte 3He Longitudinal Polarization
246468 11/03/08 15:55 Katich NMR measurement
246466 11/03/08 15:38 Jack Gas Shed air flow
246465 11/03/08 15:33 A. Camsonne Event size with RICH run 2721
246463 11/03/08 15:13 E. Schulte N. Okay asked about the "pickup coil" target position and controlled access procedures
246459 11/03/08 14:25 E. Schulte Bad Fan
246458 11/03/08 14:17 Katich EPR measurement
246457 11/03/08 14:17 V. Sulkosky Pedestal Shifts for S2m Right ADC Channels
246456 11/03/08 14:03 Katich NMR measurement
246453 11/03/08 13:45 E. Schulte Gas Shed Low Pressure Alarm
246452 11/03/08 13:43 E. Schulte Gas Shed Air Flow Alarm
246447 11/03/08 12:52 mercado Calibrated Lumi HVs
246444 11/03/08 12:20 E. Schulte Target & Reference Cell Readbacks
246443 11/03/08 12:11 E. Schulte Day Shift Tools Screen
246441 11/03/08 12:01 adaq Radiation and Ozone snapshot
246440 11/03/08 12:01 adaq Hall A Tools snapshot
246439 11/03/08 12:01 adaq BigBox Magnet snapshot
246438 11/03/08 12:01 adaq Beam Position and Current snapshot
246437 11/03/08 12:01 adaq Beam Energy in A and C snapshot
246436 11/03/08 12:00 Katich Target -> 3He Cell in beam
246435 11/03/08 11:47 Lindgren laser off
246434 11/03/08 11:39 jpchen Helium flow rate to beam line Be window
246433 11/03/08 11:33 brads Harp scan
246431 11/03/08 11:14 mercado Setting Lumi HVs
246430 11/03/08 11:14 V. Sulkosky Shortened Fast EPICS Logger Readout
246426 11/03/08 10:57 brads Added shielding around BB DAQ network switch
246425 11/03/08 10:54 brads 4x4 (MCC units) raster on BeO snapshot
246424 11/03/08 10:54 E. Schulte Back to Beam Permit, MCC to Deliver Beam
246423 11/03/08 10:47 brads/kalyan MWDC power supplies good, but cards are oscillating...
246421 11/03/08 10:39 brads Ethane/argon for MWDC is at 300psi
246419 11/03/08 10:36 R. Michaels Lumi saturated
246418 11/03/08 09:48 E. Schulte Controlled Access to Check Power Supplies
246416 11/03/08 09:24 brads Set MWDC 2, 3 threshold to 8V for now
246415 11/03/08 09:11 adaq Noise in BB wirechamber 2 and 3
246414 11/03/08 09:01 brads High deadtime issue connected to MWDCs
246412 11/03/08 08:39 E. Schulte HRS-L Q3 cryo event picture
246411 11/03/08 08:32 E. Schulte HRSL Q3 Helium Liquid Level fault
246409 11/03/08 08:14 ycwang re :A1 peak
246407 11/03/08 08:07 Ibrahim Owl Shift Summary
246401 11/03/08 07:37 Chiranjib target STATUS
246400 11/03/08 07:34 V. Sulkosky Re: A1 HV
246398 11/03/08 07:22 ycwang RICH main HV tripped during the beam tuning
246397 11/03/08 07:13 ycwang A1 peak corresponding to HV
246396 11/03/08 07:13 samuel on beam line is reference cell with 0 pressure N2
246395 11/03/08 07:12 Chiranjib NMR measurement
246394 11/03/08 07:05 samuel move to He3 cell position
246393 11/03/08 06:59 Ibrahim Deadtime is High for Production
246392 11/03/08 06:57 samuel move to pickup coil position
246390 11/03/08 06:38 Ibrahim Q3 is Up again.
246386 11/03/08 06:07 Ibrahim Q3 Tripped Again!
246385 11/03/08 06:04 xqian We See 3He polarization
246384 11/03/08 06:00 Chiranjib He-3 elastic peak
246383 11/03/08 05:59 xqian T7 rates goes extremely high suddenlly
246378 11/03/08 05:02 Chiranjib Raster check
246376 11/03/08 04:57 Chiranjib TARGET ACTIVITY
246372 11/03/08 04:42 samuel move to He3 Cell position
246371 11/03/08 04:35 Ibrahim Q3 Helium Level
246370 11/03/08 04:26 samuel move target to pickup coil position
246367 11/03/08 04:20 Ibrahim 1H03B Scanner
246364 11/03/08 04:01 adaq Radiation and Ozone snapshot
246363 11/03/08 04:01 adaq Hall A Tools snapshot
246362 11/03/08 04:01 adaq BigBox Magnet snapshot
246361 11/03/08 04:00 adaq Beam Position and Current snapshot
246360 11/03/08 04:00 adaq Beam Energy in A and C snapshot
246359 11/03/08 03:58 samuel move to he3 position
246358 11/03/08 03:42 samuel move to reference cell H2(100PSI)
246357 11/03/08 03:39 Ibrahim BeO Screen
246356 11/03/08 03:38 Ibrahim Spot++ for Run 2728
246355 11/03/08 03:36 Ibrahim Rebooted ROCs 8, 9 and 10.
246352 11/03/08 03:23 Ibrahim Tools Screen (Q3 is Up)
246351 11/03/08 03:17 Ibrahim MCC Harp Scan Results (1H03B Scanner is not working)
246350 11/03/08 03:06 Ibrahim Energy Monitor is Working
246349 11/03/08 02:56 Ibrahim Q3 Tripped
246348 11/03/08 02:50 Ibrahim MCC are still working on the Harp Scans
246347 11/03/08 02:16 jinhuang Some discuss on Polarization loss during last Rotation-NMR combination
246346 11/03/08 01:15 Ibrahim Q3 Lead Flows are Steady
246345 11/03/08 01:06 Ibrahim Beam Tuning
246344 11/03/08 00:35 adaq Radiation and Ozone snapshot
246343 11/03/08 00:35 adaq Hall A Tools snapshot
246342 11/03/08 00:35 adaq BigBox Magnet snapshot
246341 11/03/08 00:35 adaq Beam Position and Current snapshot
246340 11/03/08 00:35 adaq Beam Energy in A and C snapshot
246339 11/03/08 00:31 xqian BCM scaler calibration
246338 11/03/08 00:09 Jin Huang Target Expert Summary for Nov 02 Day-Swing Shift
Sunday
246336 11/02/08 23:56 gilman swing shift summary, 2 nov 2008
246335 11/02/08 23:42 gilman Q3 liquid levels
246334 11/02/08 23:35 gilman beam coming back
246333 11/02/08 23:27 Jin Target Back to Trans and Start pumping
246332 11/02/08 23:22 Jin Targeg Motion
246331 11/02/08 23:01 E.Chudakov Moller results
246330 11/02/08 22:54 Chiranjib TRANSVERSE PUMPING LASERS
246329 11/02/08 22:52 Jin Huang Stablized Target Trans Pumping Spec (Correct photo)
246328 11/02/08 22:49 Jin Huang Stablized Target Trans Pumping Spec
246325 11/02/08 22:15 Tareget Grou EPR Calibration @ Transverse Pumping
246324 11/02/08 22:04 gilman lead flows
246323 11/02/08 22:01 gilman Q3 helium
246322 11/02/08 21:55 gilman Q3 trip
246321 11/02/08 21:45 A. Camsonne Blaster test for RICH ROCs
246319 11/02/08 21:26 xqian BCM calibration BCM1 and BCM2 vs OLO2
246318 11/02/08 21:16 Target Group Target -> Trans Pumping
246312 11/02/08 20:23 brads Beam energy readbacks are frozen
246310 11/02/08 20:17 A. Camsonne RICH RUN during Moeller with timing scan
246305 11/02/08 20:01 adaq Radiation and Ozone snapshot
246304 11/02/08 20:01 adaq Hall A Tools snapshot
246303 11/02/08 20:00 adaq BigBox Magnet snapshot
246302 11/02/08 20:00 adaq Beam Position and Current snapshot
246301 11/02/08 20:00 adaq Beam Energy in A and C snapshot
246300 11/02/08 19:59 Jin Huang Laser Absorption
246298 11/02/08 19:39 Kalyan Beam RF time signal as seen by BigBite scintillator
246296 11/02/08 19:37 Chiranjib EPR measurement
246295 11/02/08 19:35 Kalyan BB wirechamber turned off
246294 11/02/08 19:34 gilman getting ready for Moller
246293 11/02/08 19:23 jinhuang Target -> Vertical+
246292 11/02/08 19:16 Chiranjib Vertical lasers turned ON
246291 11/02/08 19:05 Jin Huang Target Movement for NMR, Moller; Pol.= 28%
246290 11/02/08 19:01 gilman getting ready for Moller
246288 11/02/08 18:56 gilman beam positions
246287 11/02/08 18:50 Gilman shift plan followup 2
246284 11/02/08 18:31 gilman Moller to start at about 19:00
246282 11/02/08 18:11 jinhuang Target Movement for NMR; Pol.= 31%
246278 11/02/08 17:40 gilman shift plan followup
246275 11/02/08 17:09 gilman shift plan
246274 11/02/08 17:07 gilman cryo on call going home
246271 11/02/08 16:29 jinhuang Before taking beam@4PM, he3 Pol. is estimated to be 38%.
246268 11/02/08 15:59 Jin Target Move
246267 11/02/08 15:58 E. Schulte Day Shift Summary, November 2, 2008
246265 11/02/08 15:27 E. Schulte Request Cancelled
246263 11/02/08 15:12 E. Schulte Q3 Coming back online
246262 11/02/08 15:10 brads Auto-snapshots of Tools, BPMs, etc started
246261 11/02/08 15:08 adaq Radiation and Ozone snapshot
246260 11/02/08 15:08 adaq Hall A Tools snapshot
246259 11/02/08 15:08 adaq BigBox Magnet snapshot
246258 11/02/08 15:08 adaq Beam Position and Current snapshot
246257 11/02/08 15:08 adaq Beam Energy in A and C snapshot
246256 11/02/08 14:49 E. Schulte Controlled Access Requested
246255 11/02/08 14:39 E. Schulte Update on Cryo Situation
246254 11/02/08 14:21 zhangyi NMR calibration mistake & polarization after 1 run
246253 11/02/08 14:18 E. Schulte Just heard from Cryo On-Call
246252 11/02/08 14:13 E. Schulte Polarization really about 44%
246251 11/02/08 14:02 E. Schulte J. P. called, it's not so bad
246250 11/02/08 14:01 E. Schulte Beam is potentially back, still have cryo problems.
246249 11/02/08 14:01 brads (tech) Problem with LHe level readback on LQ3?
246248 11/02/08 13:52 E. Schulte Target polarization down to 14%
246247 11/02/08 13:43 eschulte Cryo-excersion tripped Q3
246246 11/02/08 13:31 E. Schulte HRS-L Q3 Interlock Fault
246245 11/02/08 13:14 E. Schulte Beam Down Again
246243 11/02/08 13:06 E. Schulte BTA not recording beam use properly
246242 11/02/08 13:02 E. Schulte Vince Responded
246239 11/02/08 12:41 E. Schulte Dead Time Estimate on DATAMON
246238 11/02/08 12:29 E. Schulte Tools Screen for Day Shift
246237 11/02/08 12:20 Kalyan Trigger rates vs N2 pressure in reference cell
246236 11/02/08 12:17 E. Schulte Production running under way!
246234 11/02/08 12:15 brads Pass change --> 2 pass has been deferred until Tuesday
246233 11/02/08 12:09 zhangyi target ->He3 cell
246232 11/02/08 12:02 zhangyi target-> pick-up coil
246229 11/02/08 11:58 E. Schulte Raster scan on 3He, 2uA Beam Current
246225 11/02/08 11:49 zhangyi Temperature screen print
246223 11/02/08 11:46 E. Schulte This was not the carbon optics with hole, it was the BeO one
246221 11/02/08 11:45 zhangyi NMR Calibration for Longitudinal Polarization
246219 11/02/08 11:24 E. Schulte Spot++ Quick Check after first NMR Measurement
246216 11/02/08 10:29 E. Schulte Carbon Hole Spot++ With Raster
246215 11/02/08 10:29 Yi Zhang turn off laser
246214 11/02/08 10:27 E. Schulte Harp Scan from MCC
246208 11/02/08 09:54 brads Modified names in runstart.tcl target list
246207 11/02/08 09:47 E. Schulte Spot++ Carbon Hole No Raster
246201 11/02/08 09:15 E. Schulte MCC Called: Sending CW beam
246200 11/02/08 08:24 Ibrahim Owl Shift Summary
246199 11/02/08 08:23 E. Schulte MCC Called, Checking Orbits
246198 11/02/08 07:56 Ibrahim No Beam
246197 11/02/08 07:53 ycwang A1 high voltage adjust (Right side)
246196 11/02/08 07:46 ycwang A1 high voltage adjust
246195 11/02/08 07:38 Chiranjib NMR measurement and VI
246193 11/02/08 07:26 Chiranjib NMR at the end of the shift
246191 11/02/08 07:11 adev BigBite WireChamber Measured Voltage change!
246190 11/02/08 07:02 Ibrahim MCC Harp Scan Results
246189 11/02/08 07:00 adev BigBite WireChamber Measured Voltage change!
246188 11/02/08 06:58 Chiranjib MCC working on the beam
246187 11/02/08 06:55 Ibrahim BeO Screen
246186 11/02/08 06:53 brads Beam wandering over last couple hours
246185 11/02/08 06:43 Chiranjib spot++ after harp scan
246183 11/02/08 06:30 Chiranjib ROC 9 issue could be radiation related??
246182 11/02/08 06:29 adev BigBite WireChamber Measured Voltage change!
246181 11/02/08 06:25 chiranjib Harp scan
246178 11/02/08 06:00 Chiranjib ROC 9 died again
246175 11/02/08 05:55 Ibrahim Tools Screen
246174 11/02/08 05:53 Chiranjib ROCs rebooted
246172 11/02/08 05:41 Chiranjib ROC 9 problem
246167 11/02/08 05:16 adev BigBite WireChamber Threshold change!
246166 11/02/08 05:13 ycwang Run # 2687 2688 2689 2690 are all junks.
246165 11/02/08 05:05 adev BigBite WireChamber Measured Voltage change!
246164 11/02/08 04:59 Ibrahim Beam is Very Trippy (RF Problems)
246163 11/02/08 04:56 chiranjib TARGET---> OPTICS
246162 11/02/08 04:55 adev BigBite WireChamber Measured Voltage change!
246161 11/02/08 04:44 adev BigBite WireChamber Measured Voltage change!
246160 11/02/08 04:37 Ibrahim BB Wire Chambers Trips
246159 11/02/08 04:33 adev BigBite WireChamber Measured Voltage change!
246158 11/02/08 04:30 Chiranjib ROC 9 died again
246157 11/02/08 04:23 adev BigBite WireChamber Measured Voltage change!
246155 11/02/08 04:19 Chiranjib ROC 9 rebooted
246153 11/02/08 04:12 adev BigBite WireChamber Measured Voltage change!
246150 11/02/08 04:01 adev BigBite WireChamber Measured Voltage change!
246144 11/02/08 03:51 adev BigBite WireChamber Threshold change!
246137 11/02/08 03:40 adev BigBite WireChamber Threshold change!
246132 11/02/08 03:30 adev BigBite WireChamber Threshold change!
246125 11/02/08 03:19 adev BigBite WireChamber Measured Voltage change!
246121 11/02/08 03:12 xqian residual
246117 11/02/08 03:08 adev BigBite WireChamber Threshold change!
246114 11/02/08 03:07 xqian BigBite Wire Chamber residual, need to recalibrate
246113 11/02/08 03:07 xqian BigBite Wire Chamber residual, need to recalibrate
246106 11/02/08 02:58 adev BigBite WireChamber Threshold change!
246100 11/02/08 02:52 Ibrahim ROC9 Rebooted Itself
246097 11/02/08 02:47 adev BigBite WireChamber Threshold change!
246092 11/02/08 02:35 Ibrahim Beam is unstable again
246091 11/02/08 02:20 xqian some thinkings
246090 11/02/08 02:16 adev BigBite WireChamber Measured Voltage change!
246089 11/02/08 02:14 xqian Background rates and the tracking efficiency, include multi track tracking efficiency
246087 11/02/08 02:07 Chiranjib Bb HV tripped ( #2664)
246085 11/02/08 02:05 adev BigBite WireChamber Measured Voltage change!
246084 11/02/08 01:54 Chiranjib BB HV tripped ( #2663)
246073 11/02/08 01:52 adev BigBite WireChamber Measured Voltage change!
246071 11/02/08 01:41 adev BigBite WireChamber Measured Voltage change!
246070 11/02/08 01:35 samuel move to H2 target with pressure 100PSI
246081 11/02/08 01:32 Ibrahim Clock Changes (End of Daylight Saving Time)
246069 11/02/08 01:32 adaq target: Spin Flip Report: It's confirmed that Q Wave Plate Direction is correct. Spin up in stepper motor control code -> Spin along holding field
246080 11/02/08 01:30 adaq target: Spin Flip Report: Spin Flip Subsystem is now calibrated and working for Vertical Pumping.
246067 11/02/08 01:24 Ibrahim Beam is stable at 2 uA.
246064 11/02/08 01:16 Ibrahim Results of the MCC Harp Scans
246063 11/02/08 01:10 adev BigBite WireChamber Measured Voltage change!
246059 11/02/08 01:04 Ibrahim BeO Screen Shot
246058 11/02/08 01:04 Ibrahim BeO Screen Shot
246056 11/02/08 00:59 Chiranjib Spot size with Be0
246051 11/02/08 00:22 Xiaofeng Swing shift summary for 11/01/2008
Saturday
246046 11/01/08 23:53 Xiaofeng snapshot of hall A tools screen
246045 11/01/08 23:51 cusanno RICH work at 2200V with 35 kHz T3
246044 11/01/08 23:35 Xiaofeng MCC harp scan
246042 11/01/08 23:07 Jin Spot ++ Screen shot
246040 11/01/08 22:59 xqian BigBite Wire Chambers trip a lot
246039 11/01/08 22:56 Jin Target -> Optics (BeO)
246038 11/01/08 22:55 adev BigBite WireChamber Measured Voltage change!
246037 11/01/08 22:48 Yi Qiang EPR calibration of vertical pumping
246036 11/01/08 22:44 adev BigBite WireChamber Measured Voltage change!
246035 11/01/08 22:40 xqian N2 pressure curve for RUN2513-RUN2524
246034 11/01/08 22:33 Jin Huang Target -> Reference Cell
246032 11/01/08 22:23 brads (techs) Shimmed BB sieve pole up by +4"
246026 11/01/08 22:00 brads MWDC gas supply at 400 psi
246025 11/01/08 21:59 brads ROC8, 9, 10 resets plugged back in
246024 11/01/08 21:58 brads BB Sieve OUT
246023 11/01/08 21:50 Techs/spiege Sieve movement
246022 11/01/08 21:34 Xiaofeng Hall switched to controlled access
246019 11/01/08 21:13 Jin Huang MWDC Position survey with Bigbite magnets off run # 2635
246018 11/01/08 21:08 A. Camsonne RICH HV to 2200V
246017 11/01/08 21:05 A. Camsonne RICH screens
246015 11/01/08 20:44 Xiaofeng beam current dropped from 2 uA to 0.5 uA from 20:30 to 20:40.
246014 11/01/08 20:39 Kalyan preshower vs shower energy
246012 11/01/08 20:19 Xiaofeng bigbite collimator was installed at ~11:00 am this morning.
246009 11/01/08 20:04 Xiaofeng bigbite magnet current at 710 A
246006 11/01/08 19:45 Xiaofeng Bigbite Magnet current: 600 A
246004 11/01/08 19:28 cusanno RICH stands at 2300V
246002 11/01/08 19:27 Xiaofeng bigbite magnet current ramp up to 400 A
245999 11/01/08 19:13 Xiaofeng Bigbite magnet current: 200A
245995 11/01/08 19:01 V. Sulkosky Analysis of Bulls eye Scan
245994 11/01/08 19:00 Kalyan DAQ is back again
245992 11/01/08 18:45 Kalyan ROC 8, 9 , 10 reboots
245989 11/01/08 18:28 brads More ROC9 investigation next CA
245985 11/01/08 18:15 Jin Target Temperature
245984 11/01/08 18:15 Yi Qiang Laser absorption for vertical line
245980 11/01/08 17:31 Xiaofeng bigbite magnet off
245978 11/01/08 17:23 Jin Target -> Optics
245976 11/01/08 17:20 adev BigBite WireChamber Measured Voltage change!
245974 11/01/08 17:09 adev BigBite WireChamber Measured Voltage change!
245972 11/01/08 16:55 Xiaofeng beam current increased to 2 uA
245969 11/01/08 16:50 ycwang rich pedestal with beam in run2623
245966 11/01/08 16:34 ycwang rich pedestal with beam in run 2623
245962 11/01/08 16:01 E. Schulte Day Shift Summary for November 1, 2008
245961 11/01/08 15:56 adev BigBite WireChamber Measured Voltage change!
245959 11/01/08 15:46 adev BigBite WireChamber Measured Voltage change!
245957 11/01/08 15:35 adev BigBite WireChamber Measured Voltage change!
245955 11/01/08 15:25 adev BigBite WireChamber Measured Voltage change!
245954 11/01/08 15:19 E. Schulte BigBite Wire chamber channels tripped again.
245953 11/01/08 15:15 E. Schulte MCC to send 1uA.
245952 11/01/08 15:07 E. Schulte MCC has taken beam to tune to Hall B
245950 11/01/08 15:04 adev BigBite WireChamber Measured Voltage change!
245946 11/01/08 14:59 E. Schulte BigBite tripped again, and other status related issues
245945 11/01/08 14:56 E. Schulte CODA end of run failed
245944 11/01/08 14:54 E. Schulte Another BigBite wire chamber channel trip.
245942 11/01/08 14:53 adev BigBite WireChamber Measured Voltage change!
245941 11/01/08 14:51 E. Schulte BigBite Wire Chamber channel L14.7 tripped again
245937 11/01/08 14:34 E. Schulte Beam's Back!
245934 11/01/08 14:05 E. Schulte RICH pedestal run for Alexandre
245932 11/01/08 13:53 E. Schulte MCC: Problem with an RF cavity
245931 11/01/08 13:44 adaq target: Spin Flip Report: No Shift Worker Comment
245929 11/01/08 13:11 Kalyan BB wire chamber HV
245928 11/01/08 13:11 E. Schulte Beam Back, BigBit On!
245926 11/01/08 12:58 adev BigBite WireChamber Measured Voltage change!
245925 11/01/08 12:53 Kalyan BB Chamber HV trips
245923 11/01/08 12:47 adev BigBite WireChamber Measured Voltage change!
245922 11/01/08 12:46 E. Schulte Hall A tools for Day Shift
245921 11/01/08 12:46 Kalyan wirechamber HV tripped
245919 11/01/08 12:32 E. Schulte BigBite channels tripped again, and the Beam is Back!
245918 11/01/08 12:26 adev BigBite WireChamber Measured Voltage change!
245917 11/01/08 12:21 E. Schulte Another BigBite wire chamber trip
245916 11/01/08 12:19 E. Schutle A few BigBite wire chamber channels tripped
245915 11/01/08 12:16 adev BigBite WireChamber Measured Voltage change!
245913 11/01/08 12:10 Katich H2 valve is currently open
245912 11/01/08 12:07 Katich reference cell pressure = 99psig
245910 11/01/08 12:02 Kalyan additions to db_DL.dat
245907 11/01/08 11:50 E. Schulte BigBite Wire Chamber Tripped
245906 11/01/08 11:47 E. Schulte Changed back to H2 Reference Cell
245905 11/01/08 11:27 Katich Target moved to Optics position
245904 11/01/08 11:25 E. Schulte Finished with Controlled Access & Starting with BeO target
245900 11/01/08 11:05 Kalyan Fixed the prescaled T1 channel in TDC
245897 11/01/08 10:47 V. Sulkosky BCM U10 Working
245896 11/01/08 10:30 V. Sulkosky Left HRS Gas Cerenkov HV
245894 11/01/08 10:20 cusanno RICH trips
245893 11/01/08 10:19 E. Schulte Compton Slot 15 Channel 10 was tripped off
245892 11/01/08 10:19 cusanno RICH trips
245889 11/01/08 09:52 E. Schulte H2 Reference Cell Valve Issue
245888 11/01/08 09:48 E. Schulte Controlled Access: Insert Target Collimator
245886 11/01/08 09:39 E. Schulte Awaiting the arrival of the RC for advice on run plan
245883 11/01/08 09:31 xqian ROC9 seems to be healed
245876 11/01/08 08:30 E. Schulte Beam Trippy
245873 11/01/08 08:14 eschulte BigBite Magnet Setting at Start of Shift
245872 11/01/08 08:04 Ibrahim Owl Shift Summary
245870 11/01/08 08:01 Jin Huang Transverse pumping laser pol. measurement
245867 11/01/08 07:52 Chiranjib Couple of issues
245866 11/01/08 07:47 riordan BigBite Momentum Reconstruction
245865 11/01/08 07:36 Jin & Chiran EPR Pol. = 31% (still pumping)
245864 11/01/08 07:35 adev BigBite WireChamber Measured Voltage change!
245861 11/01/08 07:31 xqian Wire Chamber Trip frequently when the beam is back
245860 11/01/08 07:24 adev BigBite WireChamber Measured Voltage change!
245858 11/01/08 07:16 xqian Vertex z and the sieve pattern for the LH2 run 2605
245856 11/01/08 07:13 adev BigBite WireChamber Measured Voltage change!
245855 11/01/08 07:08 riordan Frequent BigBite MWDC Trips
245854 11/01/08 07:03 adev BigBite WireChamber Measured Voltage change!
245853 11/01/08 06:53 xqian Scaler T1 trigger seems to not working for the prescaled one
245852 11/01/08 06:52 adev BigBite WireChamber Measured Voltage change!
245849 11/01/08 06:50 ycwang Run # 2581-2602 are all Junk
245848 11/01/08 06:42 xqian F1 TDC OUTPUT FIFO overflow IN run 2604 and 2605
245845 11/01/08 06:36 riordan H2 Replays F1 FIFO overflow
245844 11/01/08 06:16 xqian see the vertex too!
245842 11/01/08 06:15 xqian A plot with more stat.
245839 11/01/08 06:11 Ibrahim Beam is back
245836 11/01/08 06:02 adaq target: Spin Flip Report: Test
245833 11/01/08 06:00 riordan BigBite Sieve Plate
245832 11/01/08 05:59 samuel move to H2(100psi) reference cell
245830 11/01/08 05:52 Jin & Chiran EPR AFP working
245811 11/01/08 05:21 Ibrahim Controlled Access
245810 11/01/08 05:21 Ibrahim ROCs 8,9,10 cannot be rebooted remotely
245805 11/01/08 05:05 Jin & Chiran Got EPR signal during FM Sweep
245801 11/01/08 04:37 xqian BBNORM class is half working
245799 11/01/08 04:25 riordan ROC 9 problems
245794 11/01/08 04:05 adev BigBite WireChamber Measured Voltage change!
245793 11/01/08 04:05 Jin & Chiran NMR signal after vertical pumping
245791 11/01/08 04:02 Ibrahim BeO Spot Shot
245790 11/01/08 04:01 riordan ROC 9 Stopped Responding
245789 11/01/08 04:00 Ibrahim Tools Screen
245786 11/01/08 03:54 adev BigBite WireChamber Measured Voltage change!
245784 11/01/08 03:45 riordan Carbon Hole spot++
245780 11/01/08 03:30 jpchen He3 gas bottle valve is open, ready for reference cell run
245778 11/01/08 03:29 Chiranjib Laser absorption for vertical line
245777 11/01/08 03:26 jpchen target work
245776 11/01/08 03:25 samuel move to optics hole position.
245775 11/01/08 03:25 adaq spot++ on BeO
245773 11/01/08 03:21 Chiranjib Vertical Laser Pumping started
245772 11/01/08 03:21 Chiranjib Vertical Laser Pumping started
245768 11/01/08 02:53 samuel move to optics position
245766 11/01/08 00:36 adev BigBite WireChamber Measured Voltage change!
245764 11/01/08 00:05 Gao swing shift summary